A Complete Bibliography of Publications in the ICES Journal of Marine Science (1930–1939)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.04

Title word cross-reference

l_{1} [DH31]. l_{2} [DH31].

11.XI.1939 [Ano39d]. 12.IV.1878 [Ano39d]. 1929 [Büc30c, DM33]. 1930 [Tho30a]. 1931 [Gra31a]. 1935 [Kot36].

28 [DM33]. 28-VI-1929 [DM33]. 29th [Gra31a].

38 [Ber39].

A. [And31]. Ålands [Lis38]. Ålandsmeer [Lis38]. Aale [Ott36]. Absorption [UJ34]. abwärts [Ott31, Ott32, Ott36]. Acartia [Red34]. Account [Rai34, Rob36b]. Accuracy [EC34]. Acid [Buc33b, BG32]. Action [Hic39, Hod33]. adhesion [Kän32]. Aeration [Zen31]. af [Gra31a]. Age [Dix31b, Dix31a, Dix34, Moh30, Moh34, Büc30c]. Air [Atk32]. Albuminoid [RW34a, RW34b]. Alizarin [Rai35]. allgemeinen [Nie35]. along [Hes31]. Alters [Moh30, Moh34]. Alters- [Moh30, Moh34]. Alterszusammensetzung [Büc30c]. am [Rei31]. America [She31]. American [LBC30]. Ammonia [RW34a, RW34b]. Amoebae [SML35]. Analyses [CT37, Tul38, TC38]. Analysis [BW35b, BW36, Gib33, Par39]. Anglian [SH34]. Animals [Gar37, LBC30]. Annual [Sei35]. Anwendung [Nie38d]. Apparat [Sch38b]. Apparatus [Mou39, Web30, Sch38b]. Appearance [Rob33]. Application [BWF33, BW38, Gra35b, Wim38b]. Arctic [Sch37a]. area [AS39, RW34b]. Arkona [Ges33]. Arkonabecken [Ges33]. Arktis [Sch37a]. Arrangement [FB34]. ascribable [BW35a]. Aspects [Luc38c]. Asserson [S.37i]. Atlantic [Sei35, Sei37, Sei39, She31]. Atmospheric [RU39]. Atmungsmechanismus [Hen39]. Atoms [Tul38]. August [But34]. Ausscheiden [Lüb31].

bacteriological [Sch38b]. bakteriologische [Sch38b]. Balance [Wim38a]. Baltic [Dix31a, Dix34, Ges33, Gri37a, Hes31, Man37, SR39, ZO35, vB36]. Baltischen [ZO35]. Bann [Men36]. Barents [Awe35, Ide31, KV30, KV32, Ver32]. Barentsmeer [Awe34]. Barentsmeeres [Sch30b, Awe35]. basados [dB34a]. based [But34, dB34a]. Basin [Ges33, Sta36]. Basis [BW35b, BW36]. Bay [dB34b]. Becher [Lis30]. Bedeutung [Skl39]. Beeinflussung [Bus33]. Behaviour [CJ39, Gra31a, RB32]. Beitrag [Ges33]. Bemerkungen [Nie35]. benötigten [Awe33]. Beobachtungen [Awe35]. Bericht [Büc31b]. Berücksichtigung [Run33]. besonderer [Run33]. Bestimmung [Awe33, Kän32, Kre39, Moh30, Moh34, Nie35, Pal31c]. Between [Par36a, Lis38, Luc36a, Nie37a, Par36b, Rob36a]. Bibliographie [Moh30, Moh34]. Bibliography [Ano30a, Ano30b, Ano31a, Ano31b, Ano31c, Ano32b, Ano33a, Ano33b, Ano33c, Ano34a, Ano34b, Ano34c, Ano35a, Ano35b, Ano35c, Ano36a, Ano36b, Ano36c, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano39a, Ano39b, Lew34, RB32, TP32b, TP30a, TP32c, TP38a, TP38b, TP39, Moh30, Moh34]. bifilosa [Red34]. Bight [Büc31b, Kot36]. Biological [Hes31, Rus39b, Yon31b]. Biology [Spä35, Ges33]. Biscay [dB34b]. blue [VV34]. Board [Har35b, Har36]. Bodengreifer [Tha38a]. Bodengreifers [Tha38a]. body [Awe33]. Book [B.36b, B.36a, Jac36a, Sch36c]. Boric [Buc33b]. Borley [Mau39]. Bottle [Van37b]. Bottom [Spä35, Zen31]. Brackish [Red34, Red35]. Brandt [Rei31]. breathing [Hen39]. Breeding [Wim38a, Nol37]. brief [She31]. Bucht [Kot36, Büc31b]. Bucket [Wim37].

C [And31]. Calcium [Gri37a, Gri37b]. Calculating [Tul37]. California [Jan39]. Canada [She31]. Capt [Tån33]. Capturing [FB34]. Carbonic [Buc33b, BG32]. Caribbean [Par39, Sei38]. Carl [Gra31b]. Carmichael [M.31a]. Catch [Bor30]. Catches [Dix32]. Caught [B.S32, Dix34, Luc36b, Dix31a]. Celluloid [Nes34]. central [vB36]. centres [She31]. Centrifugal [GF37a]. Certain [Luc36a]. Chaetognatha [Fra37, Fra39]. Challenger [Gra31a]. Changes [KV30, SR32, Ver32]. Channel [Tai36]. characteristics [Run33]. Characters [Jen39a]. Charles [K.38]. Chemical [Lew34]. Chlorinity [CT37, TJW31, TW31, UTT34]. chlorophyll [Kre39]. Chlorophylls [Kre39]. Chrysaora [Kra34]. Circulation [Tho33, Zen31, Wüs30]. Civil [Lüb31]. Cl [CT37]. Classification [Mar32]. Classified [Ano32a]. Clean [Ort30]. Climat [Ber39]. Climate [Ber39]. Climatic [Par36b]. Closing [Van37a]. Closure [Red39]. Clupea [Awe35, Dah36, RD33, Tes38]. Clupeoid [Rob33]. coast [Has39, Hes31, Man37, SR39, ZO35, AS39, Dan33b]. Coasts [But34, RW34a, She31]. Cod [Dan32a, Dan32b, DS33, Gra34, Rol32, Tån33, Tån37, Tån38a]. Cold [Dan30a]. Collecting [GF37a]. Collection [EC34]. Collections [Gib33, Har35b]. Collectors [Lee30]. Colorimeter [Sun31a]. Colorimetric [TC38]. Colorimetry [Sun31a]. Commercial [Hod33]. Communities [Luc38c]. Community [Wim38a]. Comparative [Tha38a]. comparison [Hic31]. Component [BW35a]. Components [RU39, UM37]. Composition [Dix32, Büc30c, ZO35]. Computing [DH31]. Concentration [KV30, Sei31]. concerning [Har35a]. Condition [But34, Far36]. Conditions [Luc36a, RU39, Rus39b]. Conductivity [TTU34]. connection [Lis38]. conservation [Nol37]. consideration [Run33]. Considerations [Rus31]. constancy [Run33]. Constant [MT35]. Constructive [Ano33d]. Consumption [KV32]. Content [Gri37a, KV30, LBC30, Sei31, SR32, Ges33, vB36]. continuation [Moh30, Moh34]. contribution [Ges33]. Contributors [Ano34f, Ano34g]. Control [Fle39, Hav39a]. Convenient [Nes34]. Conversion [CT37]. Copepods [Red34]. Cornish [Hic39]. Corrections [Tul37]. Correspondence [Car30b]. Count [Tes38]. Crangon [Hav30a, Hav30b]. Cresol [MT35, MBR36]. critical [Bha31]. Criticism [Ano33d]. Criticisms [Nal32]. critique [Gil33]. Culture [Col38]. Cunningham [Bid35]. Current [Ano30a, Ano30b, Ano31a, Ano31b, Ano31c, Ano32b, Ano33a, Ano33b, Ano33c, Ano34a, Ano34b, Ano34c, Ano35a, Ano35b, Ano35c, Ano36a, Ano36b, Ano36c, Ano37a, Ano37b, Ano37c, Ano37d, Ano38a, Ano39a, Ano39b, Car35, Doo30, Ekm32, Par36b, Tho39a, TP32b, TP30a, TP32c, TP38a, TP38b, TP39, Wit32, Pal31c]. Current-Meter [Car35, Ekm32, Wit32]. Currents [Car30a]. Cutting [Tån38a].

Dab [Jen39a]. Dahl [Hic31]. Daily [Sei39]. Danemark [Knu39]. Danish [Kra34]. Danzig [Dix32]. Data [BW31b, Gra38b, Awe33]. Daten [Awe33]. David [M.31h]. Daylight [ACF^{+}38, Pet35, UJ34, UM37]. deadline [Pal31c]. Death [Dan30a]. Decline [TP36]. Deep [Tul37, Wüs30]. Deep-Sea [Tul37]. Degrees [TW31]. Dejecta [Web30]. Denigès’ [Iba33a, Iba33b]. depend [Bor30]. Depths [Par39, Sei38, Tho33, BG32, Sch38b]. Description [Sun31a]. Designed [VDJ34]. Determination [Gri37b, Iba33a, TB35a, vB38, Moh30, Moh34, Nie35, Kre39, Pal31c]. determine [Awe33]. determining [Kän32]. Deutschen [Büc31b, Kot36]. Developments [Dah36]. Device [Van37a]. Devices [Har35b]. Diatom [Fle39]. Diatoms [Bra37]. diesen [Skl39]. Different [DH31]. Diffusion [Par39, Sei38]. Digestive [Yon31a]. Direction [Par36b]. Dissociation [MT35, MBR36]. Distribution [BW35a, BW38, Fra37, Fra39]. Diurnal [Luc36b, Wim38a]. does [Bor30]. Dover [Car30a, vV38]. downwards [Ott31, Ott32, Ott36]. Dr [Ano33f]. Dr. [Nal32]. drei [Wüs30]. Drift [Far36, Hic39, Hod33, Pal31c]. Drift-Net [Hic39]. Drift-Nets [Far36]. due [Kän32]. Dunajec [Dix31b]. during [But34, Dan32b, Fra37, Ham33, RW34b]. Dutch [Hav30a, Hav30b]. Dynamic [Par36b].

Early [Dan32b]. East [SH34, SR39, TB35b]. eastern [Büc30c, Par39]. Eau [DM33]. Eaux [Ber39]. Eddy [Par39, Sei38]. Editor [Ble35a, BW34b, Gra35a, Hod35, Nie39, Ott34, Par37, vV37]. Editorial [Ano37d, Ano38b]. edulis [CJ39]. Eel [But34]. eels [Ott36]. Effect [Iba33a, MBR36, Red39]. Eggs [Mou39, Rol32]. Ehrenbaum [Lüb31, Lüb31]. Einfluss [BG32]. einige [Awe33, Awe35]. Ekmax [Pet30]. Electrical [TTU34, Van37a]. electricity [Lis38]. Elvers [Men36]. England [But34, Rob36b]. Entnahme [Sch38b]. Entrance [Sei31]. Equilibrium [Buc33b, BG32]. Ergebnisse [Awe35, vB36]. Errata [Gra38a]. Erratum [Ano32c, Ano34d, Ano37e]. Error [DH31, VV34]. Erwärmung [Sch37a]. Essay [Sto39]. Estimation [Iba33b, Wim38b]. Estudios [dB34a]. etc [DH31]. Étude [Ber39]. Evidence [Par39]. Evolution [Sta36]. exact [Nie35]. exakten [Nie35]. Examination [DM33]. Exceptional [Kra34, Car30a]. excessively [Bha31]. Expedition [Ham33, vE30, vR32]. Experimental [Bra37, Luc36a, Bus33]. Experimentelle [Bus33]. Experiments [Bow33, CJ39, DS33, GF37b, Gra38b, Hod33, Rae39, RD33, Büc31b]. Explanation [Sve38b]. Exploiting [Gra35b]. Expression [Tul38]. External [Par36b, Rol32]. Extrusion [Web30].

Fabr [Hav30a, Hav30b, Hav30a, Hav30b]. Factor [Far36]. Factors [DS33, Par36b]. Faroe [Rae39, Tai36]. Faroese [Tån38b]. Fat [Wim38b]. Fauna [Red35, Spä35]. Features [Dix31c, Tån37, Tai36]. fed [Bha31]. Feeding [Wim38a]. Fertilité [Ber39]. Fertility [Ber39]. Fertilized [Mou39]. Feuerschiff [Lis38]. Fifteen [Van37b]. Fifteen-Gallon [Van37b]. First [Bor30]. Fischbestandes [Awe33]. Fische [Skl39]. Fischen [Moh30, Moh34, Sch31c]. Fischerei [vB36]. Fischerei-Instituts [vB36]. Fish [Ano32d, Ano32f, Bor30, Dan30a, DH31, DS33, Far36, Ide31, Nes34, Red39, Rob36a, RB32, Tån38a, VDJ34, Awe33, Moh30, Moh34, Sch31c, Skl39]. Fisheries [Red39, TP36, vB36]. Fishery [BW38, EC34, Gra35b, Hic31, Hun38, Ily34, Rob36b, SH34]. Fishes [FB34, Men34, Rai35, Rou34, Yon31a]. Fishing [Gra38b, Rai36]. Fjords [SR32]. flatfish [Skl39]. flesus [Has39]. Flødevig [DS33]. floes [Büc31b]. floor [Tha38a]. flounder [Has39]. Fluctuations [Jen39a, RU39]. Fluoride [Iba33b]. Forellen [Ott32, Ott36, Ott31]. Form [Wim37]. formed [DD37]. Fortsetzung [Moh30, Moh34]. Free [RW34a, RW34b]. French [Ber39, dB34b]. Frequency [BW35a]. Freshly [Mou39]. Fridtjof [Tho30a]. Frühjahrshering [Run33]. Fulton [J.30]. Further [Dah36, Hod33].

Gadus [Dan33b]. Gallon [Van37b]. garnet [Hav30a, Hav30b]. Gascogne [dB34b]. Gear [Hic31]. Gebiet [AS39]. general [Nie35]. gentle [Büc31b]. Geographical [Sta36]. German [AS39, Awe33, Awe35, BG32, Büc30c, Büc31b, Büc34c, Bus33, Ges33, Has39, Hav30a, Hav30b, Hen39, Hen38, Kän32, Kot36, Kre39, Lis38, Lis30, Lüb31, Man37, Moh30, Moh34, Nie33, Nie35, Nie38d, Nol37, Ott31, Ott32, Ott33, Ott36, Pal31c, Run33, Sch38b, Sch37a, SR39, Sch31c, Skl39, Tha38a, VV34, Wüs30, ZO35, vB36]. gestorben [Rei31]. Gewässern [Hav30a, Hav30b]. Gewässers [Awe33]. Glass [Lee30]. Golfe [dB34b]. Graduation [BW31b]. Granat [Hav30a, Hav30b]. Graphical [CT37, Tul37]. grapple [Tha38a]. Grass [But34]. Gravity [TW31]. Grazing [Fle39]. greater [BG32]. Green [K.38]. Grösse [Awe33]. grösseren [BG32]. Groups [Luc36b]. Growth [Dix31b, Dix31a, Dix34, Moh30, Moh34]. Guide [Ano32d]. Gulf [Dix32, ZO35]. Gustaf [Pet30].

Haddock [Rai36, AS39]. Hake [Hic31]. Halibut [Tån38b]. halves [Wüs30]. Hansen [Gra31b, Tån33]. Hardy [GF37b]. harengus [Awe35]. Hatchery [DS33]. Hauptlinien [Büc34c]. Hensen [Gib39, Wim37]. Heringslaich [AS39]. Herring [Ano33d, BWF33, Dah36, Far36, For33, Gra31a, Luc36b, RD33, Run37a, SH34, Tes38, Wim38b, AS39, Run33]. Herrings [Dix32, Far38, Hes31, Lis34]. High [Dan32b]. Hole [Ano30d]. holländischen [Hav30a, Hav30b]. Horn [Doo30]. Horns [Tho39a]. Horns-Rev [Tho39a]. Hose [GF37a]. Hydrogen [KV30, Sei31]. Hydrogen-Ion [Sei31]. Hydrographic [BW38]. Hydrographical [Rus39b]. Hydrographie [Sch30b]. Hydrography [Tai36]. hysoscella [Kra34].

Iceland [Sei31]. Icelandic [Rae39]. II [Moh30, Skl39, Moh30, Skl39]. III [Moh34, Moh34]. Illumination [Utt33]. im [AS39, Awe34, Büc30c, Ges33, Kot36, Lis38, ZO35, vB36]. images [Bus33]. Importance [Leb33, Spä35, Skl39]. Impressions [Nes34]. Improved [Ekm32, Lis30]. inclusive [Bow33]. incretory [Skl39]. Index [Ano30c, Ano31d, Ano32e, Ano33e, Ano34e, Ano35d, Ano36d, Ano37f, Ano39c, TP38c]. Indian [Tho33]. Indicated [Rus39b]. Indicator [Fra37, GF37b]. Industry [She31]. Influence [Buc33b, Dan32a, DS33, Par36b, SH34, BG32]. influenced [Dan32b]. Influences [Rol32]. influencing [Bus33]. infolge [Kän32]. inkretorischen [Skl39]. Institute [vB36]. Institution [Ano30d]. Instituts [vB36]. Integration [Luc38c]. Intensity [Rai36, RU39]. Inter [Luc36a]. Inter-Relations [Luc36a]. international [Büc34c, Hun38]. internationalen [Büc34c]. Invertebrates [Yon31a]. inverted [Nie33]. Investigation [Spä35, Nie33]. Investigations [Run37a, Büc34c, Sch38b, vB36, Has39, Hun38, Skl39, ZO35]. Iodine [LBC30]. Ion [KV30, Sei31]. Ireland [Men36]. irideus [Bha31]. Irish [Far38, Lew34]. Iron [TB35a, TB35b]. Island [RW34a]. Isles [Yon31b]. IX [Ano34e].

J. [Pau32]. Jahre [Büc30c, Kot36, vB36, Büc34c]. James [Dan33a]. Januar [Rei31]. January [Gra31a]. Johannes [Ano33f, R.33a]. Johansen [And31]. John [Mau39]. Johnstone [Dan33a]. Jones [M.31h]. Joseph [Bid35]. jungen [Ott31, Ott32, Ott36]. Jungflundern [Has39]. Jungschollengründen [Büc31b].

Kalkgehalt [vB36]. Karl [Rei31]. Kilogram [CT37]. Kirstine [Ano39d]. Knipovich [Hjo39]. Knoll [Doo30]. Knud [N.38c]. Koenigsberg [vB36]. Kohlensäuregleichgewicht [BG32]. Königsberg [vB36]. Konservierung [Nol37]. Konstantenbestimmungen [BHM33]. Konstanz [Run33]. Küste [ZO35, AS39, SR39].

L [Has39, Has39]. L. [Dan33b, Mur32]. Laboratory [Mou39, Yon31b]. Lachse [Ott31, Ott32, Ott33, Ott36]. Larvae [CJ39, DS33, Nol37]. Larval [Leb33]. Lateral [Par36a]. Latvian [Man37, ZO35]. Layer [Sei37]. Lee [Rob36a]. Leistungsfähigkeit [Tha38a]. Lemon [Rae39]. Letter [BW34b, Gra35a, Nie39, Ott34, Par37, vV37]. Letters [Ble35a, Hod35]. lettländischen [Man37, ZO35]. letzten [Büc34c]. Light [Dan32a, Dan32b]. lightship [Lis38]. Lightvessels [Tho39a]. lime [vB36]. lines [Büc34c]. List [Ano32a, Ano33f]. literature [Beh34]. Litre [CT37, Tul38]. living [SML35]. Log [Car35]. London [Gra31a]. Low [Yon31b]. Lowestoft [Gra36, TP36]. Lunar [SH34].

M [Hjo39]. Machine [VDJ34]. Magen [AS39]. main [Büc34c]. Mallotus [Awe34]. Manipulation [Har35b]. Marine [Har35b, LBC30, Spä35, Yon31a, Nie33, Yon31b]. marinem [Nie33]. Marines [dB34b]. Mark [Men32, Mot31, Nal32]. Marked [Tån38b]. Marking [Ano32f, Bow33, Gra38b, Ide31, Rae39, Tån33]. Markings [Rae39]. Marks [Ano32d]. Matter [vB38]. Maxima [Sve38b]. May [Tho30a]. Mean [Tes38]. Measurement [ACF^{+}38, BW31b, Utt33]. Measuring [Har35a, Har36]. mechanism [Hen39]. Mediterranean [Ber39]. Méditerranéennes [Ber39]. Meerbusen [ZO35]. Meeres [ZO35]. Meerestiefen [BG32]. Meerwasser [Kre39]. Meerwasserproben [Sch38b]. Meerwassers [BHM33, ZO35]. Mer [DM33]. Meridional [Wüs30]. Meridionale [Wüs30]. Mesh [Far36]. Metamorphose [Skl39]. metamorphosis [Skl39]. Meteorological [PU35]. Meter [Car35, Ekm32, Wit32]. Method [Bra37, CT37, GF37a, Gri37b, Hav39a, Iba33a, Iba33b, Nes34, Tån33, Tån38a, Tul37, Wim38b, Nie35]. Methode [Nie35]. Methods [Ano33d, BWF33, Gra32, Hen38]. Microprojection [VDJ34]. microscope [Nie33]. Mid [Par39, Sei38]. Mid-Depths [Par39, Sei38]. Mielck [H.34b]. migrating [Ott31, Ott32, Ott36]. Migration [Men34, Rou34, Tån37]. Migrations [Tån38b]. Mikroskop [Nie33]. mil [CT37]. Milligram [Tul38]. Milligram-Atoms [Tul38]. Minima [Sve38b]. M’Intosh [M.31a]. mittleren [vB36]. Mixing [Par36a]. Mixture [Dix32]. Modern [BWF33, Gra35b]. Modified [Har36]. Mollusca [Leb33]. Monaco [Ber39, Ber39]. Montagnes [dB34b]. Morphometrical [Dix31c]. Mortality [Gra38b]. Mottram [Nal32]. Mountains [dB34b]. Movements [Car30a, Web30, vV38]. mug [Lis30].

N [Pau32, Yon31b]. N. [Hjo39]. Nachträge [Moh30, Moh34]. Nanoplancton [Ber39]. Nanoplankton [Ber39]. Nansen [Tho30a]. Natural [Gra38b]. needed [Awe33]. Needle [Tån33]. Net [Gib39, Har35a, Hic39, Van37a, Wim37]. Netherlands [Red34, Red35]. Nets [Far36, FB34, Hod33, WW36, Nie38d]. network [Lis30]. Netz [Lis30]. Netzfängen [Nie38d]. Neue [BHM33]. Nielsen [Pau32]. Nitrat [Ges33]. Nitrate [KV30, Ges33]. Nitrates [Ruu30]. Nitrite [SR32, Ver32]. Nitrogen [RW34a, RW34b, vB38]. nördlichen [Lis38]. Nordsee [Büc30c]. Nordwest [AS39]. normale [DM33]. North [Sei35, Sei37, Sei39, Tai36, vR36, Büc30c, Gra34, Gra35b, Mur32, Rai36, Rus39b, Sei31, She31, SR32, Sta36, TB35b, UJ34]. North-East [TB35b]. North-Norwegian [SR32]. Northern [Men36, Lis38, Tai36, vR36]. Northwest [AS39]. Norway [DD37]. Norwegian [AS39, Dan33b, Run33, Run37a, Sei31, SR32]. norwegischen [AS39, Run33]. Note [Ano38b, Ano38c, Har35a, Iba33a, Rae39, SML35, Tho39a]. Notes [BW37, Fra37, Rob33]. Notice [Ano32f, Ano34f, Ano34g, Ano37d]. Number [Dan32b, Dan33b, Far38, For33, Kän32, dB34a]. Numerical [Wim38a]. número [dB34a]. Nutrient [Coo33, KV32, Sei35].

Observations [CJ39, EC34, PU35, Sei31, Ver32, Awe35, Doo30, Pal31c]. obtained [Ham33]. Occurrence [Kra34, Red34, Rob36a, TB35b, AS39]. Ocean [TJW31, UTT34, BG32, RW34a, TB35b, Tho33, UJ34]. Oceanic [FB34]. Oceanographic [Ano30d, vB36]. Oceanography [Lew34]. Oceans [Sve38b, Wüs30]. October [Tho30a, But34]. octopus [Kra34]. off [RW34a, SR39]. Oliver [Mau39]. Optical [Rob33]. Organe [Skl39]. Organic [RW34a, RW34b, Sei35]. Organisms [Rus39b]. Organization [Ily34]. organs [Skl39]. Ostenfeld [Gra31b]. ostpreussischen [SR39]. Ostrea [CJ39]. Ostsee [SR39, vB36, Ges33]. Ostseeküste [Man37]. Other [Tån38a]. Otoliths [Tån38a]. Otter [Hic31]. Otto [H.34b]. Ova [Far38]. Overfishing [Rus31]. Oxygen [Sei37, Sve38b]. Oxygen-Poor [Sei37]. Oyster [Col38, Hav39a, Lee30, She31]. Oysters [Web30]. Ozeane [Wüs30]. ozeanographischen [vB36].

P [DM33]. Pacific [TB35b, UJ34, Dah36, RW34a, RD33, She31]. pallasii [Dah36, RD33, Tes38]. Paper [Lee30, Gra31a]. Papers [Ano33f]. Part [Ott32, Ott36]. Particulate [vB38]. Parts [DH31]. Passamaquoddy [Hun38]. Pelagic [Red34]. performance [Tha38a]. Petersen [Tha38a]. Petersen-Bodengreifers [Tha38a]. pH [BG32]. Phenol [MBR36]. phenomena [Kän32]. Phenomenon [Rob36a]. Philosophic [BW35b, BW36]. Phosphat [Ges33]. Phosphate [Gar37, Iba33a, Iba33b, KV30, Sei31, Ges33]. Phosphate-Content [Sei31]. Phosphates [Ruu30]. Phosphorus [Sei35]. Photometer [Gra36]. Physical [DS33, Lew34]. Phytoplankton [Gra32, Luc36a, Nie37a, Wim38a, Nie38d]. Phytoplanktonuntersuchungen [Nie38d]. Pilchard [Hic39]. pilchardus [dB34a]. Plaice [Ble33a, Bow33, Jen39a, TP36, Büc30c, Büc34c, Kot36, Man37]. Plankton [Bra37, Gib33, GF37a, GF37b, Har35a, Har35b, Leb33, Luc38c, Nie33, Rus39b, SR39, Van37a, Wim38b, WW36, Ges33, Hen38, Nie35, SR39]. Plankton-Net [Har35a, Van37a]. Planktonarbeiten [Nie35]. Planktonforschung [Hen38]. Planktongehalt [Ges33]. Planktonic [Gar37]. platessa [Man37]. Plattfische [Skl39]. Pleuronectes [Has39, Man37]. Plymouth [Rus34]. Polish [Bor30, Dix31a, Dix31c, Dix34]. Polluted [Dod30]. polychaete [Nol37]. Polychaetenlarven [Nol37]. ponds [Ott33]. Poor [Sei37]. population [Awe33, Büc30c, Kot36]. Populations [Fle39]. Pr [vB36]. Prediction [Hav39a]. Preliminary [Rai34, Rob36b]. Preparing [Nes34]. Present [But34, Ano33d]. Preservation [Iba33b]. pressure [BG32]. Prince [Knu39]. Probable [Par36a]. Problem [Dod30, Rus31, Sch31c]. Problems [Ano33d, BWF33, BW38, Gra31a, Gra32]. Procedure [TC38]. process [Skl39]. Processes [Par36a, Par39, Yon31a]. Production [Gar37, Ort30, Sei35, Ges33]. Produktionsbiologie [Ges33]. Professor [Ano33f, Lüb31]. Prussia [vB36]. Prussian [SR39]. Publications [Red35]. Puget [RW34b]. Pump [GF37a].

Quantitative [Bra37, Gib33, Spä35, vB38, Hen38, Nie33, Nie35, Nie38d]. quantitativen [Nie35, Nie38d]. Quantities [Nie37a]. Queensland [Yon31b].

R [Knu39]. Race [Ano33d, BWF33, Sch31c]. Races [Lis34]. Racial [Jen39a, Run33, dB34a]. raciales [dB34a]. Radiation [Atk32, RU39]. Radioactivity [DTU32]. Rassenmerkmale [Run33]. Rassenproblem [Sch31c]. Rassenverhältnisse [Run33]. Rate [Dix31b]. Rates [Gra38b]. Ratio [TJW31]. Rational [Coo33]. read [Gra31a]. Rearing [Mou39]. record [She31]. Recording [Web30]. Red [MT35, MBR36]. Reda [Dix31b]. Refractivity [UTT34]. Refractivity-Chlorinity-Temperature [UTT34]. regarding [Hes31]. Region [Sei38]. Regression [BW35a]. Reife [Ott33]. related [Wim38a]. relating [Par39]. Relation [Nie37a, Rob36a, TW31, Far36]. Relations [Luc36a, Run33]. Relationship [Par36a, Par36b]. Relationships [UTT34]. Remarks [BW31b, Nie35]. Replenishment [Rai36]. Reply [Nal32]. Report [Bor30, But34, RW34b, Büc31b]. Reporting [Coo33]. Requirement [Sei35]. Research [Bor30, Ily34, Hen38]. resignation [Lüb31]. Results [Ham33, Wit32, vB36, Awe35]. Rev [Tho39a, Doo30]. Reversing [Tul37]. Review [Run37a, Sto39]. Reviews [A.31a, A.31b, A.33, A35, Alm31, And30, Atk31a, Atk31b, Atk31c, BW30a, BW31a, BW34a, B.30a, B.30b, B.30c, B.31a, B.31b, B.31c, B.31d, B.32, B.33, B.34a, B.34b, B.34c, B.35, B.36b, B.36a, B.37b, B.37a, B.38a, B.38b, Bck33, Bck35, Beh34, Ber37, Bie30, Bie32a, Bie32b, BW30b, Ble30, Ble31a, Ble31b, Ble33b, Ble33c, Ble34, Ble35b, Bog37, Bra35a, Bra35b, Bra36, Bra39a, Bra39b, Bru32, Buc30a, Buc31a, Buc33a, Buc34a, Büc30b, Büc33c, Büc34b, Büc36, Büc38, Bul32, Bul38, C.30, C.32, C.34b, C.34c, C.34d, C.34a, C.35b, C.35a, C.36c, C.36a, C.36b, C.37a, C.37b, C.37c, C.38c, C.38d, C.38a, C.38b, C.38e, C.38f, C.39a, C.39b, C.39d, C.39c, Cal39, Car30c, Car31a, Car31b, Car32, Car33a, Car33b, Car34, Car39, Clo30, Cor34, D.30a, D.30b, D.35a, D.35b, D.37, D.38a, D.38b, Dah38, Dan30b, De 31, Dea37, Def30, Def38a, Def38b, Def38c, Def39a, Def39b, Die37b, Die37a, Die38d, Die38a]. Reviews [Die38b, Die38c, Die39a, Die39b, E.32, E.35, E.37a, E.37b, Eh30, Ehr36a, Ehr36b, Ehr37, Ehr38, Ehr39, Ekm30, Ekm33a, Ekm33b, Ekm34, Elm30, Ert34, F.30d, F.30e, F.30f, F.30b, F.30c, F.30a, F.31a, F.31b, F.31c, F.31d, F.32a, F.32b, F.33a, F.33b, F.34, F.35d, F.35a, F.35b, F.35c, F.36a, F.36b, F.36c, F.37a, F.37b, F.37c, F.38b, F.38c, F.38d, F.38a, F.38e, F.38f, F.39, Fag33, Fje34a, Fje34b, Fje34c, Fra34, G.30, G.31a, G.31d, G.31e, G.31b, G.31c, G.31f, G.31g, G.32a, G.32b, G.32c, G.33, G.34a, G.34c, G.34d, G.34e, G.34b, G.34f, G.35g, G.35k, G.35a, G.35b, G.35c, G.35d, G.35e, G.35h, G.35i, G.35l, G.35f, G.35j, G.36b, G.36c, G.36e, G.36d, G.37b, G.37a, G.38a, G.38b, G.39a, G.39e, G.39f, G.39c, G.39b, G.39d, Gar35, Gib37, Gil31, Gil34, Gra30, H.30, H.31d, H.31e, H.31a, H.31b, H.31c, H.31f, H.32d, H.32b, H.32c, H.32e, H.32f, H.32a, H.33c, H.33a, H.33b, H.34c, H.34d, H.34e, H.34a, H.34f, H35f, H.35a, H.35b, H.35c, H.35d, H.35e, H.36b]. Reviews [H.36c, H.36d, H.36a, H.37a, H.37b, H.37h, H.37g, H.37c, H.37d, H.37e, H.37f, H.38d, H.38a, H.38e, H.38f, H.38g, H.38h, H.38b, H.38c, H.39b, H39a, H.39c, H.39d, Hag31, Hag33b, Hag33c, Hag33a, Hag37a, Hag37b, Han38, Hav31, Hav33, Hav34, Hav37a, Hav37b, Hav39b, Hav39c, Hel31, Hel35, Hel39, Hen32, Hen31, Her30, I.34a, I.34b, I.35, I.37, I.38, I.39, Ise39, J.31, J.34, J.35, J.36, J.37, Jac30, Jac32a, Jac32b, Jac33a, Jac33b, Jac34, Jac35a, Jac35b, Jac36a, Jac36b, Jac37, Jef32, Jen34, Jen39b, Jes39, Jes30, Jes31a, Jes31b, Jes31c, Jes32a, Jes32b, Jes35, Jes37, Joh31, Joh32, Jou32, Jud31, K.39, KÜN32, Kän33a, Kän33b, Kän34, Kii32, Knu30, Kra39, KüU32, Kün31, L.30b, LH30, L.30c, L.30a, L.31b, L.31a, L.31c, L.32a, L.32c, L.32b, L.32d, L.32e, L.33b, L.33c, L.33d, L.33a, L.34b, L.34a, L.34c, L.35a, L.35b, L.36, L.37a, L.37b, L.37c, L.37d, L.37f, L.37g, L.37e]. Reviews [L.38a, L.38b, L.39, Lb.38, Luc38a, Luc38b, Luc39, M.30, M.31e, M.31f, M.31g, M.31b, M.31c, M.31d, MO31, M.32c, M.32a, M.32b, M.32d, M.33a, M.33b, M.34c, M.34a, M.34b, M.35b, M.35c, M.35a, M.37a, M.37b, M.37c, M.37d, M.38a, M.38b, M.38c, M.39a, M.39b, M.39c, M.39d, Mat30a, Mat30b, McE33, Mee30, Mie33, Mol35a, Mol35b, MI39, Mor31, Mos37a, Mos37b, Mos38, Mos39, N.30a, N.30b, N.31, N.34, N.38b, N.38a, Nes37, Nie37b, Nie38a, Nie38b, Nie38c, Nor36, Nor39a, Nor39b, Nor39c, O.30, O.31d, O.31a, O.31b, O.31c, O.38, O.39, Omm35, Ort34, Ort38, Ort39, Ost30, Ott37, Ott38, PT33, PT36, PT38, PT39, P.34a, P.34b, P.34c, P.35a, P.35b, P.36, P.37c, P.37a, P.37b, P.38g, P.38h, P.38i, P.38a, P.38b, P.38c, P.38d, P.38e, P.38f, P.39a, P.39b, P.39e, P.39c, P.39d, Pal31a, Pal31b, Pal32, Pen32, Pen33, Pen34a, Pen34b, Pen35, Pen37a, Pen37b, Pen38, Pou31, Pou32a, Pou32b, Pou38, Pra34, Pro31, Pro32]. Reviews [R.30, R.31a, R.31b, R.31c, R.32a, R.32b, R.32c, R.32d, R.33b, R.33c, R.34b, R.34a, R.35a, R.35b, R.35c, R.36a, R.36b, R.36c, R.36d, R.36f, R.36e, R.37a, R.37c, R.37d, R.37e, R.37f, R.37b, R.37g, R.37h, R.37i, R.38a, R.38b, R.38c, R39a, R.39b, R.39c, R.39d, R.39e, Ram30, Ram37a, Ram37b, Ram37c, RA39a, RA39b, RA39c, Rie32, Rit36, Rob39, Ros38, Run37b, Run37c, Run38, Run39, Rus39a, Ruu31a, Ruu31b, Ruu33, Ruu35a, Ruu35b, Ruu38, Ruu39, S.30g, S.30b, S.30d, S.30e, S.30f, S.30i, S.30c, S.30h, S.30a, S.31f, S.31a, S.31b, S.31d, S.31e, S.31c, S.32c, S.32k, S.32l, S.32m, S.32d, S.32h, S.32i, S.32a, S.32b, S.32n, S.32e, S.32f, S.32j, S.32g, S.33g, S.33b, S.33a, S.33d, S.33e, S.33f, S.33c, S.34c, S.34d, S.34b, S.34a, S.35c, S.35d, S.35e, S.35a, S.35b, S.36b, S.36c, S.36a, S.36d, S.36e, S.37e, S.37g, S.37h, S.37a, S.37d, S.37f, S.37j, S.37b, S.37c, S.38f, S.38e, S.38c, S.38a, S.38b, S.38d, S.39b, S.39a, Sch38a, Sch39a, Sch30a, Sch31a, Sch31b]. Reviews [Sch32a, Sch34a, Sch34b, Sch35a, Sch35b, Sch36a, Sch36b, Sch37b, Sch31d, Sch35c, Sch39b, Sch32c, Sch32b, Sch33a, Sch30c, Sch30d, Sch32d, Sch32e, Sch34c, Sch36c, Sch38c, Sch33b, Sch34d, Sch36d, She33a, She33b, She33c, Sie30, Sou30, Spä34, Spä36, Spe38, Ste30, Ste31, Ste37, Sun31c, Sun35, Sun36, Sun39, Sve38a, Sve31, TP31, TP32a, TP33a, TP33b, TP37, T.30a, T.30b, T.30c, T.30e, T.30f, T.30d, T.31c, T.31a, T.31d, T.31e, T.31b, T.31f, T.33b, T.33c, T.33a, T.33d, T.34, T.35a, T.35b, T.37, T.38b, T.38a, T.39a, T.39b, T.39c, T.39d, Tån31, Tån32, Tån39, Tai33, Tbo33, Tha38b, Tho30b, Tho30c, Tho38a, Tho39b, Tho30d, Tho32, Tho34a, Tho34b, Tho36, Tho37a, Tho37b, Tho37c, Tho38b, Tho38c, Tho39c, Tho39d, Tho39e, TP30b, Tre38a, Tre38b, Tut36a, Tut36b, U.38, U.39, W.30c, W.30a, W.30b, W.31, W.32a, W.32b, W.32c, W.32d, W.33a, W.33b, W.33c]. Reviews [W.33d, W.33h, W.33e, W.33f, W.33g, W.34a, W.34f, W.34b, W.34c, W.34d, W.34e, W.35a, W.35b, W.35d, W.35e, W.35c, W.35f, W.36, W.37b, W.37c, W.37a, W.38a, W.38b, W.38c, W.39a, W.39b, W.39c, W.39d, W.39e, W.39f, W.39g, Wat30a, Wat30b, Wat31, Wat32, Whi37, Whi38, Wil31, Wis38, Woo31, Wor33, Wri38, Wüs31, Wüs33, Wüs36, Yon30, Yon32, Yon34, Yon35, Yon37, dH36, dH37a, dH37b, dH39, uV37, vB37a, vB37b, vB37c, van38, van39a, van39b, vE36, vE37, vR35]. Revue [Gil33]. Rhizostoma [Kra34]. Riga [ZO35]. Rigaschen [ZO35].

Rings [Rob33]. Ripe [Ott33]. River [Men36]. Rivers [Dix31c, Dix31b]. Routine [Wim38b]. Rowland [G.36a]. Run [Men36]. Russian [Beh34].

S. [Knu39, S.37i]. Sagitta [Rus34]. salar [Ott33]. Salient [Tai36]. Salinity [MBR36, vR36]. Salmo [Bha31, Dix31b, Ott33]. Salmon [B.S32, Dix31a, Dix34, Mou39, Ott31, Ott32, Ott33, Ott36]. Salt [VV34]. Salts [Coo33, Iba33a, KV32]. Salz [VV34]. Salz- [VV34]. Samland [Has39]. Samlandküste [Has39]. Sample [SML35]. Sampler [Gib33, Sun31b]. Samples [GF37a, Iba33b, Kre39, Sch38b]. Sampling [WW36]. Sandeels [Rai34]. sardina [dB34a]. Sardine [Jan39, dB34a]. Sardinia [dB34a]. Saving [Bor30]. Saving-Trawls [Bor30]. Scale [Rob36a, Bus33]. Scales [Bha31, DH31, DD37, Men32, Mot31, Nal32, Nes34, Rob33, VDJ34]. Schellfischen [AS39]. Schichtung [Wüs30]. Schilddrüse [Skl39]. Schmidt [Ano33f, R.33a]. Scholle [Man37]. Schollenbestandes [Büc30c]. Schollenbevölkerung [Kot36]. Schollenuntersuchungen [Büc34c]. Schøning [N.38c]. Schontrawl [Büc31b]. Schöpfproben [Kre39]. Schuppenbildern [Bus33]. schwedischen [Büc31b]. Scientific [Ano33f]. scoop [Kre39]. Scottish [Fra37, Fra39, GF37b, Rai34]. screening [Hen38]. Sea [Buc33b, CT37, Coo33, DS33, DTU32, Dix31b, Dix31c, Gri37b, Iba33a, Iba33b, Men32, MT35, MBR36, Mot31, Nal32, Nie37a, Pet35, SML35, Sta36, Sun31a, TTU34, TW31, TB35a, Tul37, Tul38, TC38, vB38, Kre39, ZO35, Awe35, Büc30c, Ges33, Gra34, Gra35b, Ide31, KV30, KV32, Lew34, Lis38, Mur32, Rai36, Rus39b, SR39, Sei31, Sei38, Tai36, Ver32, ZO35, vR36, vB36]. Sea-Fish [DS33]. Sea-Trout [Dix31b, Dix31c]. Sea-Water [CT37, Tul38]. Seas [Ruu30]. Season [DD37]. Seasonal [KV30]. seawater [Sch38b]. Sections [Tån38a]. Secular [Tho39a]. Seihmethoden [Hen38]. Selected [RB32]. Selective [Hic39, Hod33]. Service [Lüb31]. Setting [CJ39, Hav39a]. Shellfish [Dod30, Ort30]. Shetland [Tai36, Bow33]. Ship [Har35b]. Simplified [Gri37b, TC38]. Size [Far38, Luc36b, Rob36a, Awe33]. Size-Groups [Luc36b]. Size-Relation [Rob36a]. Skagerak [Dan33b]. Skeletal [Rai35]. Small [FB34, Tån38b]. Smith [Ano39d, Doo30]. Snellius [Ham33, vE30, vR32]. Society [Gra31a]. Solar [Atk32, RU39]. Soles [Rae39]. Solution [BWF33, Dod30]. Some [BWF33, BW31b, CJ39, Gra31a, Har35b, Luc38c, PU35, Red35, Rus31, Tån37, Awe33, Rob33, She31, Wit32, Awe35]. Sound [RW34b]. Sources [DH31]. Sous [dB34b]. Sous-Marines [dB34b]. South [Sei31, Büc30c]. south-eastern [Büc30c]. Southern [Sei31, DD37, G.36a, Ruu30]. sowie [Ott36]. Space [BW38]. Spanish [dB34a]. Spat [Hav39a, Lee30]. spawn [AS39]. Spawning [Men32, Mot31, Nal32]. special [Run33]. Species [Fra37]. Specific [TW31]. Specimens [Bha31]. Speed [Sun31b]. Sprat [Dix32, Rob36a, Rob36b]. Sprats [Dix32]. spring [Run33]. Staatsdienst [Lüb31]. Stability [Par36a]. Stages [Dan32b]. Stain [Rai35]. standard [Lis30, Lis30]. Standard-Netz [Lis30]. starved [Bha31]. Statistical [Ano33d, BWF33, BW35b, BW36, BW38]. Statistics [EC34, Hes31]. Steam [Hic31]. Stock [Dix32, Rai36]. stomach [AS39]. Storbrotten [Lis38, Lis38]. Straits [Car30a, vV38]. stratification [Wüs30]. Strom [Lis38]. Structures [Rai35]. studies [dB34a]. Study [Bha31, Bra37, VDJ34, Ber39]. Submarine [ACF^{+}38, Pet35, Utt33, UM37]. substance [Gra31a]. Suction [GF37a]. südöstlichen [Büc30c]. Sulfate [TJW31]. Sulfate-Chlorinity [TJW31]. summer [She31]. Summers [RW34b]. Supplements [Moh30, Moh34]. Surface [Ham33, Sun31b]. Survey [But34]. surveys [Nie38d]. Susceptibility [Rol32]. Sweden [Hes31]. Swedish [Büc31b]. System [Col38, Coo33].

Tagging [Dah36, God36, Jan39, RD33]. Tags [God38]. taken [DH31]. taking [Sch38b]. Technical [BW37]. Teichen [Ott33]. Teil [Ott32, Ott36]. Teleosteer [Hen39]. Temperature [Dan32b, MBR36, Sei39, Tes38, UTT34, vR36, VV34]. Temperatures [Ham33]. Temperaturfehler [VV34]. Terminbeobachtungen [Pal31c]. tests [Tha38a]. their [BWF33]. theoretical [Rus31]. Theory [BW38, Gra35b]. Thermograph [Ham33]. Thermometers [Tul37]. Thomas [Bid35, M.31h, J.30]. three [Wüs30]. Thunnus [Mur32]. thymol [VV34]. Thymolblau [VV34]. thynnus [Mur32]. thyroid [Skl39]. Tides [Mar32]. Tiefe [Sch38b]. Tiefenzirkulation [Wüs30]. Time [Hav39a]. Tissues [Wim38b]. tonsa [Red34]. Topography [Par36b]. tour [She31]. Transmission [Atk32]. Transparency [Pet35]. Transplantations [Ble33a]. Trawl [Dix32, Hic31, Luc36b, Büc31b, Tha38a]. Trawl-Caught [Luc36b]. Trawling [Gra35b, Hic31]. Trawls [Bor30]. Triftstromes [Pal31c]. Tropical [Sei35]. Trout [Bha31, DD37, Dix31b, Dix31c, Men32, Mot31, Nal32, Ott31, Ott32, Ott36]. trutta [Dix31b]. Tuna [God36, God38]. Tunny [Mur32]. Turbinen [Ott31, Ott32, Ott36]. Turbines [Ott31, Ott32, Ott36]. Two [BW37, Jan39, Bha31, Red34]. Tyne [B.S32]. Type [Ekm32, Gib33].

U.S.S.R. [Ily34]. umgekehrtem [Nie33]. Uncertainties [Kän32]. undersized [Bor30]. Underwater [dB34b]. Units [Coo33]. Universität [vB36]. University [vB36]. Unsicherheiten [Kän32]. Untersuchung [Nie33]. Untersuchungen [vB36, Has39, Sch38b, Skl39, ZO35]. upon [MBR36]. Use [Gib33, Rai35, WW36, Wit32, Ano33d, Nie38d]. Utermöhls [Nie33]. Utermöl [Nie33].

V [Ano30c]. Valdemar [Knu39]. Vallées [dB34b]. Valleys [dB34b]. Valve [Web30]. Vancouver [RW34a]. Variation [For33, Luc36b, Tes38, Tho39a, Wim38a]. Variations [Ber39, Sei39, UM37, WW36, vR36]. Various [DS33]. Varne [Doo30]. Veen [Tha38a, Tha38a]. Veen- [Tha38a]. Veen-Bodengreifer [Tha38a]. verbesserter [Lis30]. Vergleichsversuche [Tha38a]. Versuche [Büc31b]. Vertebrae [Dan32b, Dan33b, For33, dB34a]. Vertebral [Tes38]. vértebras [dB34a]. Verteilung [Awe34]. Vertical [Car35, Par36a, Zen31]. Verwachsungserscheinungen [Kän32]. very [Dan30a]. VI [DM33, Ano31d]. Vicinity [B.S32]. Vigneron [Hic31]. VII [Ano32e]. VIII [Ano33e]. villosus [Awe34]. virens [Dan33b]. Volume [Ano30c, Ano31d, Ano32e, Ano33e, Ano34e, Ano35d, Ano36d, Ano37f, Ano39c, TP38c]. vor [SR39]. Vorgang [Skl39]. Vorkommen [AS39]. vortices [Kän32]. vulgaris [Hav30a, Hav30b]. Vyl [Tho39a].

Wachstums [Moh30, Moh34]. Wachstums-Bestimmung [Moh30, Moh34]. Walb [dB34a]. Walter [H.34b]. wandernden [Ott31, Ott32, Ott36]. Warming [Sch37a]. Washington [RW34a]. Wasserdruckes [BG32]. Water [Atk32, Car30a, CT37, Coo33, DTU32, Gri37a, Gri37b, Iba33a, Iba33b, MT35, MBR36, Pet35, SML35, Sei31, Tes38, TTU34, TW31, TB35a, Tul38, TC38, Van37b, vV38, Awe33, BG32, Kre39, ZO35]. Water-Bottle [Van37b]. Waters [Ber39, Bow33, Fra37, Fra39, GF37b, Kra34, Rai34, Red34, Red35, RW34a, RW34b, Tån38b, TJW31, TB35b, UTT34, Zen31, Hav30a, Hav30b]. Wemyss [J.30]. Western [Sei35, Wüs30, Sei37, Sei39]. Westhälften [Wüs30]. which [DD37]. Wilhelm [H.34b]. William [M.31a]. Wind [Lis38]. Winter [Car30a, DD37, Rob33]. Winters [Dan30a]. Wirbelzahl [Kän32]. Woods [Ano30d]. work [Nie35].

X [Ano35d]. XI [Ano36d]. XII [Ano37f]. XIII [TP38c]. XIV [Ano39c].

Year [SR32]. Years [Dix34, Jan39, Run37a, Büc34c, Doo30]. Young [Mou39, Büc31b, Has39, Ott31, Ott32, Ott36].

Zero [TW31]. Zones [DD37]. Zoo [Wim38a]. Zoo-Phytoplankton [Wim38a]. Zooplankton [Luc36a, Nie35, Nie37a, Wim38a]. Zostera [But34]. Zucht [Nol37]. Zuiderzee [Red39]. Zum [Sch31c, vB36]. zur [Awe33, Ges33, Nie35, Sch38b, Pal31c]. Zusammenhang [Lis38]. Zusammensetzung [ZO35]. zwischen [Lis38].

References

A:1931:Ra

[A.31a] W. R. G. A. Reviews. ICES Journal of Marine Science, 6(1):103, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

        (electronic). URL http://academic.oup.com/icesjms/article/6/
        1/103/641568.

A:1931:Rb

[A.31b] W. R. G. A. Reviews. ICES Journal of Marine Science, 6(1):103–104, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/103/641571.

A:1933:R

[A.33] W. R. G. A. Reviews. ICES Journal of Marine Science, 8(2):256–257, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/256/618372.

A:1935:R

[A35] W. R. G. A. Reviews. ICES Journal of Marine Science, 10(1):103, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/103/615386.

Atkins:1938:MSD

[ACF^{+}38] W. R. G. Atkins, G. L. Clarke, H. Fettersson, H. H. Poole, C. L. Utterback, and A. Angstrom. Measurement of submarine daylight. ICES Journal of Marine Science, 13(1):37–57, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/13/1/37/714612.

Alm:1931:R

[Alm31] Gunnar Alm. Reviews. ICES Journal of Marine Science, 6(1): 132–134, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/132/641678.

Andersson:1930:R

[And30] K. A. Andersson. Reviews. ICES Journal of Marine Science, 5(2): 248–249, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/248/627857.

Andersson:1931:CJ

[And31] K. A. Andersson. A. C. Johansen. (1867–1931). ICES Journal of Marine Science, 6(3):357–360, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/3/357/724399.

Anonymous:1930:CBa

[Ano30a] Anonymous. Current bibliography. ICES Journal of Marine Science, 5 (2):272–283, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/272/627910.

Anonymous:1930:CBb

[Ano30b] Anonymous. Current bibliography. ICES Journal of Marine Science, 5(3):438–448, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/3/438/648295.

Anonymous:1930:IVV

[Ano30c] Anonymous. Index to volume V. ICES Journal of Marine Science, 5(3):449–453, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/5/3/449/648296.

Anonymous:1930:WHO

[Ano30d] Anonymous. The Woods Hole Oceanographic Institution. ICES Journal of Marine Science, 5(2):226–228, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/2/226/627812.

Anonymous:1931:CBa

[Ano31a] Anonymous. Current bibliography. ICES Journal of Marine Science, 6 (1):142–152, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/142/641704.

Anonymous:1931:CBb

[Ano31b] Anonymous. Current bibliography. ICES Journal of Marine Science, 6(2):335–353, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/2/335/600677.

Anonymous:1931:CBc

[Ano31c] Anonymous. Current bibliography. ICES Journal of Marine Science, 6(3):506–518, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/3/506/724622.

Anonymous:1931:IVV

[Ano31d] Anonymous. Index to volume VI. ICES Journal of Marine Science, 6(3):519–524, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/6/3/519/724625.

Anonymous:1932:CL

[Ano32a] Anonymous. Classified list. ICES Journal of Marine Science, 7(2): 277–283, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/277/708212.

Anonymous:1932:CB

[Ano32b] Anonymous. Current bibliography. ICES Journal of Marine Science, 7(1):116–132, April 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/1/116/946971.

Anonymous:1932:E

[Ano32c] Anonymous. Erratum. ICES Journal of Marine Science, 7(2):321, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/321/708407.

Anonymous:1932:GFM

[Ano32d] Anonymous. A guide to fish marks. ICES Journal of Marine Science, 7(1):133–165, April 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/1/133/946976.

Anonymous:1932:IVV

[Ano32e] Anonymous. Index to volume VII. ICES Journal of Marine Science, 7(3):444–448, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/3/444/694327.

Anonymous:1932:NFM

[Ano32f] Anonymous. A notice of fish marking. ICES Journal of Marine Science, 7(3):407, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/3/407/694240.

Anonymous:1933:CBa

[Ano33a] Anonymous. Current bibliography. ICES Journal of Marine Science, 8(1):126–140, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/126/682442.

Anonymous:1933:CBb

[Ano33b] Anonymous. Current bibliography. ICES Journal of Marine Science, 8(2):287–305, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/2/287/618446.

Anonymous:1933:CBc

[Ano33c] Anonymous. Current bibliography. ICES Journal of Marine Science, 8(3):415–427, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/3/415/730436.

Anonymous:1933:HRP

[Ano33d] Anonymous. Herring race problems: a constructive criticism of statistical methods at present in use. ICES Journal of Marine Science, 8(1):9–47, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/9/682473.

Anonymous:1933:IVV

[Ano33e] Anonymous. Index to volume VIII. ICES Journal of Marine Science, 8(3):428–432, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/3/428/730441.

Anonymous:1933:LPD

[Ano33f] Anonymous. List of Professor Dr. Johannes Schmidt’s scientific papers 1899–1932. ICES Journal of Marine Science, 8(2):153–160, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/153/618340.

Anonymous:1934:CBa

[Ano34a] Anonymous. Current bibliography. ICES Journal of Marine Science, 9(1):140–156, April 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/1/140/945076.

Anonymous:1934:CBb

[Ano34b] Anonymous. Current bibliography. ICES Journal of Marine Science, 9(2):289–304, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/9/2/289/743496.

Anonymous:1934:CBc

[Ano34c] Anonymous. Current bibliography. ICES Journal of Marine Science, 9(3):432–449, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/9/3/432/615833.

Anonymous:1934:E

[Ano34d] Anonymous. Erratum. ICES Journal of Marine Science, 9(3):i, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/i/615842.

Anonymous:1934:IVI

[Ano34e] Anonymous. Index to volume IX. ICES Journal of Marine Science, 9(3):450–454, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/9/3/450/615837.

Anonymous:1934:NCa

[Ano34f] Anonymous. Notice to contributors. ICES Journal of Marine Science, 9(2):305, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/305/743508.

Anonymous:1934:NCb

[Ano34g] Anonymous. Notice to contributors. ICES Journal of Marine Science, 9(3):455, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/455/615840.

Anonymous:1935:CBa

[Ano35a] Anonymous. Current bibliography. ICES Journal of Marine Science, 10(1):127–140, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/127/615474.

Anonymous:1935:CBb

[Ano35b] Anonymous. Current bibliography. ICES Journal of Marine Science, 10(2):223–240, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/10/2/223/672942.

Anonymous:1935:CBc

[Ano35c] Anonymous. Current bibliography. ICES Journal of Marine Science, 10(3):345–361, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/10/3/345/659297.

Anonymous:1935:IVX

[Ano35d] Anonymous. Index to volume X. ICES Journal of Marine Science, 10(3):362–366, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/10/3/362/659301.

Anonymous:1936:CBa

[Ano36a] Anonymous. Current bibliography. ICES Journal of Marine Science, 11(1):117–130, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/117/640271.

Anonymous:1936:CBb

[Ano36b] Anonymous. Current bibliography. ICES Journal of Marine Science, 11(2):276–294, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/11/2/276/663251.

Anonymous:1936:CBc

[Ano36c] Anonymous. Current bibliography. ICES Journal of Marine Science, 11(3):386–400, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/11/3/386/621081.

Anonymous:1936:IVX

[Ano36d] Anonymous. Index to volume XI. ICES Journal of Marine Science, 11(3):401–404, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/11/3/401/621085.

Anonymous:1937:CBa

[Ano37a] Anonymous. Current bibliography. ICES Journal of Marine Science, 12(1):106–119, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/1/106/626344.

Anonymous:1937:CBb

[Ano37b] Anonymous. Current bibliography. ICES Journal of Marine Science, 12(2):237–253, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/2/237/753975.

Anonymous:1937:CBc

[Ano37c] Anonymous. Current bibliography. ICES Journal of Marine Science, 12(3):369–386, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/3/369/645491.

Anonymous:1937:ENC

[Ano37d] Anonymous. Editorial notice: the current bibliography. ICES Journal of Marine Science, 12(2):254–255, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/2/254/753976.

Anonymous:1937:E

[Ano37e] Anonymous. Erratum. ICES Journal of Marine Science, 12(3): 259, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/259/645322.

Anonymous:1937:IVX

[Ano37f] Anonymous. Index to volume XII. ICES Journal of Marine Science, 12(3):387–392, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/3/387/645499.

Anonymous:1938:CB

[Ano38a] Anonymous. Current bibliography. ICES Journal of Marine Science, 13(1):132–148, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/132/714604.

Anonymous:1938:EN

[Ano38b] Anonymous. Editorial note. ICES Journal of Marine Science, 13(1): 6, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/6/714624.

Anonymous:1938:N

[Ano38c] Anonymous. Note. ICES Journal of Marine Science, 13(2):236, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
236/783969.

Anonymous:1939:CBa

[Ano39a] Anonymous. Current bibliography. ICES Journal of Marine Science, 14(1):144–167, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/144/888495.

Anonymous:1939:CBb

[Ano39b] Anonymous. Current bibliography. ICES Journal of Marine Science, 14(3):432–448, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/3/432/721813.

Anonymous:1939:IVX

[Ano39c] Anonymous. Index to volume XIV. ICES Journal of Marine Science, 14(3):449–453, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/3/449/721815.

Anonymous:1939:KSI

[Ano39d] Anonymous. Kirstine Smith. 12.IV.1878–11.XI.1939. ICES Journal of Marine Science, 14(3):336, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/3/336/721625.

Aurich:1939:VHI

[AS39] Horst J. Aurich and Anneliese Stier. Über das Vorkommen von Heringslaich im Magen von Schellfischen aus dem Gebiet der norwegischen Nordwest–Küste. (German) [On the occurrence of herring spawn in the stomach of haddock from the Norwegian Northwest Coast area]. ICES Journal of Marine Science, 14(3):385–393, December 1939. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/14/3/385/721639.

Atkins:1931:Ra

[Atk31a] W. R. G. Atkins. Reviews. ICES Journal of Marine Science, 6(3): 476–477, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/476/724514.

Atkins:1931:Rb

[Atk31b] W. R. G. Atkins. Reviews. ICES Journal of Marine Science, 6(3):480, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/480/724534.

Atkins:1931:Rc

[Atk31c] W. R. G. Atkins. Reviews. ICES Journal of Marine Science, 6(3):505, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/505/724619.

Atkins:1932:SRT

[Atk32] W. R. G. Atkins. Solar radiation and its transmission through air and water. ICES Journal of Marine Science, 7(2):171–211, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/171/708158.

Awerinzew:1933:EBG

[Awe33] S. Awerinzew. Über einige zur Bestimmung der Grösse des Fischbestandes eines Gewässers benötigten Daten. (German) [On some of the data needed to determine the size of the fish population in a body of water]. ICES Journal of Marine Science, 8(2):230–234, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/230/618363.

Awerinzew:1934:VMV

[Awe34] S. Awerinzew. Über die verteilung von Mallotus villosus im barentsmeer. ICES Journal of Marine Science, 9(1):79–82, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/9/1/
79/945124.

Awerinzew:1935:EEB

[Awe35] S. Awerinzew. Einige Ergebnisse von Beobachtungen an Clupea harengus des Barentsmeeres. (German) [Some results of observations on Clupea harengus in the Barents Sea]. ICES Journal of Marine Science, 10(1):66–74, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/66/615494.

B:1930:Ra

[B.30a] G. B. Reviews. ICES Journal of Marine Science, 5(2):236–238, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/236/627827.

B:1930:Rb

[B.30b] G. B. Reviews. ICES Journal of Marine Science, 5(2):238–239, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/238/627833.

B:1930:Rc

[B.30c] G. B. Reviews. ICES Journal of Marine Science, 5(3):433–434, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/433/648283.

B:1931:Ra

[B.31a] J. R. B. Reviews. ICES Journal of Marine Science, 6(1):122–123, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/122/641637.

B:1931:Rb

[B.31b] J. R. B. Reviews. ICES Journal of Marine Science, 6(1):123, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/6/1/
123/641640.

B:1931:Rc

[B.31c] J. R. B. Reviews. ICES Journal of Marine Science, 6(1):123–124, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/123/641645.

B:1931:Rd

[B.31d] J. R. B. Reviews. ICES Journal of Marine Science, 6(2):330–331, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/330/600648.

B:1932:R

[B.32] C. E. P. B. Reviews. ICES Journal of Marine Science, 7(1):89–90, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/89/947082.

B:1933:R

[B.33] R. B. Reviews. ICES Journal of Marine Science, 8(1):121–123, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/121/682434.

B:1934:Ra

[B.34a] G. B. Reviews. ICES Journal of Marine Science, 9(3):397–398, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/397/615776.

B:1934:Rb

[B.34b] G. B. Reviews. ICES Journal of Marine Science, 9(3):398–399, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/398/615777.

B:1934:Rc

[B.34c] H. E. B. Reviews. ICES Journal of Marine Science, 9(3):412–413, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/412/615810.

B:1935:R

[B.35] H. B. B. Reviews. ICES Journal of Marine Science, 10(2):187–189, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/187/672841.

B:1936:BRb

[B.36a] K. B. Book reviews. ICES Journal of Marine Science, 11(3):369– 371, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/369/621010.

B:1936:BRa

[B.36b] P. H. B. Book reviews. ICES Journal of Marine Science, 11(3): 367–369, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/367/621006.

B:1937:Rb

[B.37a] A. B. Reviews. ICES Journal of Marine Science, 12(3):356–357, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/3/356/645424.

B:1937:Ra

[B.37b] T. B. Reviews. ICES Journal of Marine Science, 12(2):207–208, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
207/753842.

B:1938:Ra

[B.38a] T. B. Reviews. ICES Journal of Marine Science, 13(1):108–110, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
108/714472.

B:1938:Rb

[B.38b] T. B. Reviews. ICES Journal of Marine Science, 13(1):115, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
115/714497.

Bckm:1933:R

[Bck33] Bckm. Reviews. ICES Journal of Marine Science, 8(2):272–273, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/272/618418.

Bckm:1935:R

[Bck35] Bckm. Reviews. ICES Journal of Marine Science, 10(3):335–337, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/3/335/659274.

Behning:1934:RRR

[Beh34] A. L. Behning. Reviews of recent Russian literature. ICES Journal of Marine Science, 9(1):135–139, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/1/135/945072.

Berry:1937:R

[Ber37] John Berry. Reviews. ICES Journal of Marine Science, 12(1): 96–97, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/96/626453.

Bernard:1939:EVF

[Ber39] Francis Bernard. Étude sur les Variations de Fertilité des Eaux Méditerranéennes. Climat et Nanoplancton à Monaco en 1937–38. (French) [Study of the variations of fertility of Mediterranean waters. Climate and nanoplankton in Monaco in 1937–38]. ICES Journal of Marine Science, 14(2):228–241, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/2/228/737109.

Buch:1932:EWP

[BG32] Kurt Buch and Stina Gripenberg. Über den Einfluss des Wasserdruckes auf pH und das Kohlensäuregleichgewicht in grösseren Meerestiefen. (German) [On the influence of water pressure on ph and the carbonic acid equilibrium in greater ocean depths]. ICES Journal of Marine Science, 7(2):233–245, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/7/2/233/708198.

Bhatia:1931:CSS

[Bha31] D. Bhatia. A critical study of the scales of two specimens of starved and excessively fed trout — Salmo irideus. ICES Journal of Marine Science, 6(2):266–272, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/2/266/600444.

Bein:1933:NKM

[BHM33] Willy Bein, Heinz-Günther Hirsekorn, and Lotte Möller. Neue konstantenbestimmungen des meerwassers. ICES Journal of Marine Science, 8(1):48–58, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/8/1/48/682445.

Bidder:1935:JTC

[Bid35] Geo P. Bidder. Joseph Thomas Cunningham. (1859–1935). ICES Journal of Marine Science, 10(3):245–248, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/10/3/245/659211.

BierensDeHaan:1930:R

[Bie30] J. A. Bierens De Haan. Reviews. ICES Journal of Marine Science, 5(3):432–433, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/5/3/432/648282.

BierensDeHaan:1932:Ra

[Bie32a] J. A. Bierens De Haan. Reviews. ICES Journal of Marine Science, 7 (1):105–106, April 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/1/105/946914.

BierensDeHaan:1932:Rb

[Bie32b] J. A. Bierens De Haan. Reviews. ICES Journal of Marine Science, 7(1):106, April 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/1/106/946921.

Blegvad:1930:R

[Ble30] H. Blegvad. Reviews. ICES Journal of Marine Science, 5(2):247–248, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/247/627855.

Blegvad:1931:Ra

[Ble31a] H. Blegvad. Reviews. ICES Journal of Marine Science, 6(2):305–306, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/305/600495.

Blegvad:1931:Rb

[Ble31b] H. Blegvad. Reviews. ICES Journal of Marine Science, 6(2):306–308, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/306/600498.

Blegvad:1933:PT

[Ble33a] H. Blegvad. Plaice transplantations. ICES Journal of Marine Science, 8(2):161–180, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/2/161/618345.

Blegvad:1933:Ra

[Ble33b] H. Blegvad. Reviews. ICES Journal of Marine Science, 8(1):113–114, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/113/682408.

Blegvad:1933:Rb

[Ble33c] H. Blegvad. Reviews. ICES Journal of Marine Science, 8(1):120, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/120/682425.

Blegvad:1934:R

[Ble34] H. Blegvad. Reviews. ICES Journal of Marine Science, 9(3):414–416, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/414/615811.

Blegvad:1935:LE

[Ble35a] H. Blegvad. Letters to the Editor. ICES Journal of Marine Science, 10(1):99, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/99/615500.

Blegvad:1935:R

[Ble35b] H. Blegvad. Reviews. ICES Journal of Marine Science, 10(1): 113–115, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/113/615435.

Bogorov:1937:R

[Bog37] B. G. Bogorov. Reviews. ICES Journal of Marine Science, 12(3): 352–353, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/352/645415.

Borowik:1930:WDC

[Bor30] Joseph Borowik. On what does the catch of undersized fish depend? First report on the Polish research on saving-trawls. ICES Journal of Marine Science, 5(2):197–215, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/5/2/197/627803.

Bowman:1933:PME

[Bow33] Alexander Bowman. Plaice marking experiments: Shetland waters 1923–1931 (inclusive). ICES Journal of Marine Science, 8(2):223–229, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/223/618360.

Braarud:1935:Ra

[Bra35a] Trygve Braarud. Reviews. ICES Journal of Marine Science, 10(1): 104–105, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/104/615399.

Braarud:1935:Rb

[Bra35b] Trygve Braarud. Reviews. ICES Journal of Marine Science, 10(1): 105–106, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/105/615404.

Braarud:1936:R

[Bra36] Trygve Braarud. Reviews. ICES Journal of Marine Science, 11(1): 101–102, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/101/640222.

Braarud:1937:QME

[Bra37] Trygve Braarud. A quantitative method for the experimental study of plankton diatoms. ICES Journal of Marine Science, 12(3):321– 332, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/321/645339.

Braarud:1939:Ra

[Bra39a] T. Braarud. Reviews. ICES Journal of Marine Science, 14(2): 300–301, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/300/737318.

Braarud:1939:Rb

[Bra39b] T. Braarud. Reviews. ICES Journal of Marine Science, 14(3): 418, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/418/721766.

Bruun:1932:R

[Bru32] Anton Fr. Bruun. Reviews. ICES Journal of Marine Science, 7(1): 100–103, April 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/1/100/946898.

BStorrowMSc:1932:SCV

[B.S32] A. L. S. B.Storrow, M.Sc. Salmon caught in the vicinity of the Tyne, 1927–1931. ICES Journal of Marine Science, 7(3):396–406, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/396/694235.

Buch:1930:R

[Buc30a] Kurt Buch. Reviews. ICES Journal of Marine Science, 5(3):403–404, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/403/648204.

Buckmann:1930:R

[Büc30b] Bückmann. Reviews. ICES Journal of Marine Science, 5(3):408–410, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/408/648216.

Buckmann:1930:ASS

[Büc30c] A. Bückmann. Die Alterszusammensetzung des Schollenbestandes der südöstlichen Nordsee im Jahre 1929. (German) [Age composition of the plaice population in the south-eastern North Sea in 1929]. ICES Journal of Marine Science, 5(1):33–54, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/1/33/676369.

Buch:1931:R

[Buc31a] Kurt Buch. Reviews. ICES Journal of Marine Science, 6(2):293–294, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/293/600466.

Buckmann:1931:BVS

[Büc31b] A. Bückmann. Bericht über Versuche mit dem schwedischen Schontrawl auf den Jungschollengründen der Deutschen Bucht. (German) [Report on experiments with the Swedish gentle trawl on the young floes in the German Bight]. ICES Journal of Marine Science, 6(1):41– 46, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/41/641718.

Buch:1933:R

[Buc33a] K. Buch. Reviews. ICES Journal of Marine Science, 8(2):257–258, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/257/618375.

Buch:1933:BAS

[Buc33b] Kurt Buch. On boric acid in the sea and its influence on the carbonic acid equilibrium. ICES Journal of Marine Science, 8(3):309–325, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/309/730284.

Buckmann:1933:R

[Büc33c] A. Bückmann. Reviews. ICES Journal of Marine Science, 8(3):408– 409, December 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/3/408/730408.

Buch:1934:R

[Buc34a] Kurt Buch. Reviews. ICES Journal of Marine Science, 9(1):99, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/99/945140.

Buckmann:1934:R

[Büc34b] Bückmann. Reviews. ICES Journal of Marine Science, 9(2):280–283, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/280/743443.

Buckmann:1934:HIS

[Büc34c] A. Bückmann. Die Hauptlinien der internationalen Schollenuntersuchungen der letzten Jahre. (German) [The main lines of the international plaice investigations in recent years]. ICES Journal of Marine Science, 9(3):309–338, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/3/309/615748.

Buckmann:1936:R

[Büc36] Bückmann. Reviews. ICES Journal of Marine Science, 11(3):377– 378, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/377/621037.

Buckmann:1938:R

[Büc38] Bückmann. Reviews. ICES Journal of Marine Science, 13(2):256– 257, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/256/784068.

Bull:1932:R

[Bul32] H. O. Bull. Reviews. ICES Journal of Marine Science, 7(1):113, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/113/946956.

Bull:1938:R

[Bul38] H. O. Bull. Reviews. ICES Journal of Marine Science, 13(2): 260–261, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/260/784089.

Buschkiel:1933:EBS

[Bus33] A. Buschkiel. Experimentelle Beeinflussung von Schuppenbildern. (German) [Experimental influencing of scale images]. ICES Journal of Marine Science, 8(1):80–82, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/8/1/80/682465.

Butcher:1934:ZRP

[But34] R. W. Butcher. Zostera. report on the present condition of eel grass on the coasts of England, based on a survey during August to October, 1933. ICES Journal of Marine Science, 9(1):49–65, April 1934. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/9/1/49/945120.

B-W:1930:R

[BW30a] H. J. B.-W. Reviews. ICES Journal of Marine Science, 5(3):434–436, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/434/648286.

Blake:1930:R

[BW30b] Isobel Blake and I. J. F. Williamson. Reviews. ICES Journal of Marine Science, 5(3):424–426, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/5/3/424/648261.

B-W:1931:R

[BW31a] H. J. B.-W. Reviews. ICES Journal of Marine Science, 6(3):488–490, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/488/724584.

Buchanan-Wollaston:1931:SRG

[BW31b] H. J. Buchanan-Wollaston. Some remarks on the graduation of measurement data. ICES Journal of Marine Science, 6(1):47–63, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/6/1/
47/641720.

B-W:1934:R

[BW34a] H. J. B.-W. Reviews. ICES Journal of Marine Science, 9(1):119–124, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/119/945013.

Buchanan-Wollaston:1934:LE

[BW34b] H. J. Buchanan-Wollaston. Letter to the Editor. ICES Journal of Marine Science, 9(3):394–396, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/3/394/615772.

Buchanan-Wollaston:1935:CFD

[BW35a] H. J. Buchanan-Wollaston. On the component of a frequency distribution ascribable to regression. ICES Journal of Marine Science, 10 (1):81–98, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/81/615499.

Buchanan-Wollaston:1935:PBS

[BW35b] H. J. Buchanan-Wollaston. The philosophic basis of statistical analysis. ICES Journal of Marine Science, 10(3):249–263, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
249/659213.

Buchanan-Wollaston:1936:PBS

[BW36] H. J. Buchanan-Wollaston. The philosophic basis of statistical analysis. ICES Journal of Marine Science, 11(1):7–26, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/11/1/7/640298.

Buchanan-Wollaston:1937:TTN

[BW37] H. J. Buchanan-Wollaston. Two technical notes. ICES Journal of Marine Science, 12(3):333–334, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/12/3/333/645341.

Buchanan-Wollaston:1938:AST

[BW38] H. J. Buchanan-Wollaston. On the application of the statistical theory of space distribution to hydrographic and fishery problems. ICES Journal of Marine Science, 13(2):173–186, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/13/2/173/783925.

Buchanan-Wol:1933:SMS

[BWF33] H. J. Buchanan-Wol and R. A. Fisher. Some modern statistical methods: their application to the solution of herring race problems. ICES Journal of Marine Science, 8(1):7–8, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/8/1/7/682457.

C:1930:R

[C.30] R. S. C. Reviews. ICES Journal of Marine Science, 5(2):256–258, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/256/627869.

C:1932:R

[C.32] A. D. C. Reviews. ICES Journal of Marine Science, 7(3):421–422, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/421/694291.

C:1934:Rd

[C.34a] A. C. Reviews. ICES Journal of Marine Science, 9(3):419–423, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/419/615821.

C:1934:Ra

[C.34b] L. H. N. C. Reviews. ICES Journal of Marine Science, 9(1):99–100, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/99/945146.

C:1934:Rb

[C.34c] L. H. N. C. Reviews. ICES Journal of Marine Science, 9(2):263–265, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/263/743277.

C:1934:Rc

[C.34d] L. H. N. C. Reviews. ICES Journal of Marine Science, 9(3):404–405, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/404/615792.

C:1935:Rb

[C.35a] H. A. C. Reviews. ICES Journal of Marine Science, 10(2):213–215, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/213/672921.

C:1935:Ra

[C.35b] L. H. N. C. Reviews. ICES Journal of Marine Science, 10(2): 197–198, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/197/672873.

C:1936:Rb

[C.36a] H. A. C. Reviews. ICES Journal of Marine Science, 11(1):113–114, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/113/640259.

C:1936:Rc

[C.36b] H. A. C. Reviews. ICES Journal of Marine Science, 11(1):114–115, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/114/640261.

C:1936:Ra

[C.36c] L. H. N. C. Reviews. ICES Journal of Marine Science, 11(1):100, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/100/640219.

C:1937:Ra

[C.37a] H. A. C. Reviews. ICES Journal of Marine Science, 12(1):100, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
100/626322.

C:1937:Rb

[C.37b] L. H. N. C. Reviews. ICES Journal of Marine Science, 12(2): 206–207, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/206/753838.

C:1937:Rc

[C.37c] L. H. N. C. Reviews. ICES Journal of Marine Science, 12(3):344– 345, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/344/645385.

C:1938:Rc

[C.38a] H. A. C. Reviews. ICES Journal of Marine Science, 13(1):127–128, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/127/714586.

C:1938:Rd

[C.38b] H. A. C. Reviews. ICES Journal of Marine Science, 13(1):128–130, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/128/714592.

C:1938:Ra

[C.38c] L. H. N. C. Reviews. ICES Journal of Marine Science, 13(1):110, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/110/714474.

C:1938:Rb

[C.38d] L. H. N. C. Reviews. ICES Journal of Marine Science, 13(1):110, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/110/714477.

C:1938:Re

[C.38e] L. H. N. C. Reviews. ICES Journal of Marine Science, 13(3): 386, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/386/635464.

C:1938:Rf

[C.38f] W. T. C. Reviews. ICES Journal of Marine Science, 13(3):396, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/3/396/635494.

C:1939:Ra

[C.39a] H. A. C. Reviews. ICES Journal of Marine Science, 14(1):140–141, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/140/888475.

C:1939:Rb

[C.39b] H. A. C. Reviews. ICES Journal of Marine Science, 14(1):141–142, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/141/888480.

C:1939:Rd

[C.39c] L. H. N. C. Reviews. ICES Journal of Marine Science, 14(3):405– 406, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/405/721651.

C:1939:Rc

[C.39d] W. T. C. Reviews. ICES Journal of Marine Science, 14(2):313– 315, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/313/737374.

Calman:1939:R

[Cal39] W. T. Calman. Reviews. ICES Journal of Marine Science, 14(2): 290–292, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/290/737238.

Carruthers:1930:WMS

[Car30a] D. Sc. J. N. Carruthers. The water movements in the Straits of Dover. Exceptional currents in winter 1929–30. ICES Journal of Marine Science, 5(2):167–191, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/2/167/627792.

Carruthers:1930:C

[Car30b] J. N. Carruthers. Correspondence. ICES Journal of Marine Science, 5(3):437, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/437/648292.

Carruthers:1930:R

[Car30c] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 5(1): 103–105, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/103/676246.

Carruthers:1931:Ra

[Car31a] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 6(2): 291–293, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/291/600462.

Carruthers:1931:Rb

[Car31b] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 6(3): 472–476, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/472/724506.

Carruthers:1932:R

[Car32] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 7(3): 408–409, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/408/694243.

Carruthers:1933:Ra

[Car33a] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 8(1): 112–113, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/112/682405.

Carruthers:1933:Rb

[Car33b] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 8(2): 258–260, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/258/618379.

Carruthers:1934:R

[Car34] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 9(3): 405–407, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/405/615794.

Carruthers:1935:VLC

[Car35] J. N. Carruthers. A “vertical log” current-meter. ICES Journal of Marine Science, 10(2):151–168, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/10/2/151/672826.

Carruthers:1939:R

[Car39] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 14(3): 406–408, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/406/721656.

Cole:1939:SOE

[CJ39] H. A. Cole and E. W. Knight Jones. Some observations and experiments on the setting behaviour of larvae of Ostrea edulis. ICES Journal of Marine Science, 14(1):86–105, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/1/86/888521.

Closs:1930:R

[Clo30] Karl Closs. Reviews. ICES Journal of Marine Science, 5(3):402–403, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/402/648201.

Cole:1938:SOC

[Col38] H. A. Cole. A system of oyster culture. ICES Journal of Marine Science, 13(2):221–235, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/2/221/783966.

Cooper:1933:SRU

[Coo33] L. H. N. Cooper. A system of rational units for reporting nutrient salts in sea water. ICES Journal of Marine Science, 8(3):331–334, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/331/730294.

Corkan:1934:R

[Cor34] R. H. Corkan. Reviews. ICES Journal of Marine Science, 9(3):407– 408, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/407/615797.

Carter:1937:GMC

[CT37] Neal M. Carter and John P. Tully. A graphical method for the conversion of chlorinity per litre to chlorinity per kilogram (Cl per mil) in sea-water analyses. ICES Journal of Marine Science, 12(1):36–39, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
36/626354.

D:1930:Ra

[D.30a] A. D. Reviews. ICES Journal of Marine Science, 5(3):430–432, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/430/648276.

D:1930:Rb

[D.30b] A. D. Reviews. ICES Journal of Marine Science, 5(3):432, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/432/648279.

D:1935:Ra

[D.35a] G. E. R. D. Reviews. ICES Journal of Marine Science, 10(2): 196–197, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/196/672868.

D:1935:Rb

[D.35b] G. E. R. D. Reviews. ICES Journal of Marine Science, 10(3):326– 327, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/326/659243.

D:1937:R

[D.37] A. D. Reviews. ICES Journal of Marine Science, 12(2):204, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
204/753831.

D:1938:Ra

[D.38a] G. E. R. D. Reviews. ICES Journal of Marine Science, 13(3):376– 377, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/376/635446.

D:1938:Rb

[D.38b] G. E. R. D. Reviews. ICES Journal of Marine Science, 13(3):377– 378, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/377/635447.

Dahlgren:1936:FDT

[Dah36] Edwin H. Dahlgren. Further developments in the tagging of the Pacific herring, Clupea pallasii. ICES Journal of Marine Science, 11(2): 229–248, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/229/663218.

Dahl:1938:R

[Dah38] Knut Dahl. Reviews. ICES Journal of Marine Science, 13(2): 264–265, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/264/784114.

Dannevig:1930:DFV

[Dan30a] Alf Dannevig. The death of fish in very cold winters. ICES Journal of Marine Science, 5(2):194–196, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/5/2/194/627798.

Dannevig:1930:R

[Dan30b] Alf Dannevig. Reviews. ICES Journal of Marine Science, 5(2): 255–256, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/255/627867.

Dannevig:1932:ILC

[Dan32a] Alf Dannevig. The influence of light on the cod. ICES Journal of Marine Science, 7(1):53–59, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/7/1/53/947006.

Dannevig:1932:NVC

[Dan32b] Alf Dannevig. Is the number of vertebrae in the cod influenced by light or high temperature during the early stages? ICES Journal of Marine Science, 7(1):60–62, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/7/1/60/947014.

Daniel:1933:JJ

[Dan33a] R. J. Daniel. James Johnstone (1870–1932). ICES Journal of Marine Science, 8(1):3–6, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/1/3/682444.

Dannevig:1933:NVG

[Dan33b] Alf Dannevig. The number of vertebrae in Gadus virens L. from the Norwegian Skagerak Coast. ICES Journal of Marine Science, 8(3):

355–356, December 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/3/355/730303.

deBuen:1934:ERB

[dB34a] Fernando de Buen. Estudios raciales basados en el número de vértebras en la sardina, Sardinia pilchardus (Walb) (Spanish) [Racial studies based on the number of vertebrae in the sardine, Sardinia pilchardus (Walb)]. ICES Journal of Marine Science, 9(3):339–345, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/339/615753.

duBaty:1934:MVS

[dB34b] R. Rallier du Baty. Montagnes et vallées sous-marines du Golfe de Gascogne. (French) [Underwater mountains and valleys of the Bay of Biscay]. ICES Journal of Marine Science, 9(2):181–185, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/181/743129.

Dannevig:1937:SWW

[DD37] Alf Dannevig and Gunnar Dannevig. The season in which “winter” zones in the scales of trout from Southern Norway are formed. ICES Journal of Marine Science, 12(2):192–198, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/2/192/753808.

DeBuen:1931:R

[De 31] F. De Buen. Reviews. ICES Journal of Marine Science, 6(2):309–310, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/309/600505.

Deacon:1937:R

[Dea37] G. E. R. Deacon. Reviews. ICES Journal of Marine Science, 12(2): 201–202, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/201/753816.

Defant:1930:R

[Def30] A. Defant. Reviews. ICES Journal of Marine Science, 5(3):393–396, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/393/648179.

Defant:1938:Ra

[Def38a] A. Defant. Reviews. ICES Journal of Marine Science, 13(1):102–104, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/102/714443.

Defant:1938:Rb

[Def38b] A. Defant. Reviews. ICES Journal of Marine Science, 13(2):243, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
243/783989.

Defant:1938:Rc

[Def38c] A. Defant. Reviews. ICES Journal of Marine Science, 13(2):243– 244, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/243/783993.

Defant:1939:Ra

[Def39a] A. Defant. Reviews. ICES Journal of Marine Science, 14(3):408– 409, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/408/721660.

Defant:1939:Rb

[Def39b] A. Defant. Reviews. ICES Journal of Marine Science, 14(3):409– 410, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/409/721665.

Dannevig:1931:SEC

[DH31] Alf Dannevig and Per Høst. Sources of error in computing l_{1}–l_{2} etc. from scales taken from different parts of the fish. ICES Journal of Marine Science, 6(1):64–93, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/1/64/641723.

deHaan:1936:R

[dH36] J. A. Bierens de Haan. Reviews. ICES Journal of Marine Science, 11(1):112, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/112/640256.

deHaan:1937:Ra

[dH37a] J. A. Bierens de Haan. Reviews. ICES Journal of Marine Science, 12 (3):357, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/357/645428.

deHaan:1937:Rb

[dH37b] J. A. Bierens de Haan. Reviews. ICES Journal of Marine Science, 12(3):357–358, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/12/3/357/645431.

deHaan:1939:R

[dH39] J. A. Bierens de Haan. Reviews. ICES Journal of Marine Science, 14(1):135, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/135/888454.

Dietrich:1937:Rb

[Die37a] G. Dietrich. Reviews. ICES Journal of Marine Science, 12(3):341– 342, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/341/645361.

Dietrich:1937:Ra

[Die37b] Günter Dietrich. Reviews. ICES Journal of Marine Science, 12(1): 80–82, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/80/626384.

Dietrich:1938:Rb

[Die38a] G. Dietrich. Reviews. ICES Journal of Marine Science, 13(2):241, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
241/783982.

Dietrich:1938:Rc

[Die38b] G. Dietrich. Reviews. ICES Journal of Marine Science, 13(3):380– 381, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/380/635457.

Dietrich:1938:Rd

[Die38c] G. Dietrich. Reviews. ICES Journal of Marine Science, 13(3):381– 382, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/381/635459.

Dietrich:1938:Ra

[Die38d] Günter Dietrich. Reviews. ICES Journal of Marine Science, 13(1): 104–105, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/104/714447.

Dietrich:1939:Ra

[Die39a] G. Dietrich. Reviews. ICES Journal of Marine Science, 14(3):410– 411, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/410/721671.

Dietrich:1939:Rb

[Die39b] G. Dietrich. Reviews. ICES Journal of Marine Science, 14(3):411– 412, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/411/721681.

Dixon:1931:AGS

[Dix31a] Boris Dixon. The age and growth of salmon caught in the Polish Baltic. ICES Journal of Marine Science, 6(3):438–448, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/438/724468.

Dixon:1931:AGR

[Dix31b] Boris Dixon. Age and growth rate of the sea-trout (Salmo trutta) of the Rivers Reda and Dunajec. ICES Journal of Marine Science, 6(3): 449–457, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/449/724471.

Dixon:1931:MFS

[Dix31c] Borys Dixon. Morphometrical features of the sea-trout of Polish rivers. ICES Journal of Marine Science, 6(1):94–102, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/6/1/94/641725.

Dixon:1932:MHS

[Dix32] Boris Dixon. The mixture of herrings with sprats in catches with the sprat trawl, and the composition of the sprat stock of the Gulf of Danzig in 1932. ICES Journal of Marine Science, 7(3):386–396, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/386/694230.

Dixon:1934:AGS

[Dix34] Boris Dixon. The age and growth of salmon caught in the Polish Baltic in the years 1931–1933. ICES Journal of Marine Science, 9(1):66–78, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/66/945123.

Deacon:1933:EEM

[DM33] B. Sc. G. E. R. Deacon and D. J. Mathews. An examination of “Eau de Mer normale P 13, 28-VI-1929”. ICES Journal of Marine Science, 8(1):59–63, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/59/682451.

Dodgson:1930:PPS

[Dod30] R. W. Dodgson. The problem of polluted shellfish, and its solution. ICES Journal of Marine Science, 5(2):149–166, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/2/149/627787.

Doodson:1930:COH

[Doo30] A. T. Doodson. Current observations at Horn’s Rev, Varne and Smith’s Knoll in the years 1922 and 1923. ICES Journal of Marine Science, 5(1):22–32, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/1/22/676360.

Dannevig:1933:IVP

[DS33] Alf Dannevig and Erling Sivertsen. On the influence of various physical factors. on cod larvae; experiments at the Flødevig Sea-Fish Hatchery. ICES Journal of Marine Science, 8(1):90–99, April 1933. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/8/1/90/682477.

Devaputra:1932:RSW

[DTU32] D. Devaputra, Thomas G. Thompson, and Clinton L. Utterback. The radioactivity of sea water. ICES Journal of Marine Science, 7(3): 358–366, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/358/694220.

E:1932:R

[E.32] T. E. Reviews. ICES Journal of Marine Science, 7(2):301–302, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/301/708326.

E:1935:R

[E.35] R. E. Reviews. ICES Journal of Marine Science, 10(3):343, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
343/659293.

E:1937:Ra

[E.37a] R. E. Reviews. ICES Journal of Marine Science, 12(2):235, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
235/753967.

E:1937:Rb

[E.37b] R. E. Reviews. ICES Journal of Marine Science, 12(2):235, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
235/753969.

Edser:1934:OAC

[EC34] T. Edser and S. H. Crouchen. Observations on the accuracy of the collection of fishery statistics. ICES Journal of Marine Science, 9 (1):83–86, April 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/1/83/945128.

Eh:1930:R

[Eh30] Eh. Reviews. ICES Journal of Marine Science, 5(3):423–424, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/423/648258.

Ehrenbaum:1936:Ra

[Ehr36a] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 11(3): 373–375, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/373/621026.

Ehrenbaum:1936:Rb

[Ehr36b] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 11(3): 375–376, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/375/621028.

Ehrenbaum:1937:R

[Ehr37] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 12(1): 94–96, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/94/626449.

Ehrenbaum:1938:R

[Ehr38] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 13(2): 262–263, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/262/784100.

Ehrenbaum:1939:R

[Ehr39] E. Ehrenbaum. Reviews. ICES Journal of Marine Science, 14(3): 429–430, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/429/721808.

Ekman:1930:R

[Ekm30] V. Walfrid Ekman. Reviews. ICES Journal of Marine Science, 5(3): 396–398, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/396/648186.

Ekman:1932:ITC

[Ekm32] V. Walfrid Ekman. An improved type of current-meter. ICES Journal of Marine Science, 7(1):3–10, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/7/1/3/946992.

Ekman:1933:Ra

[Ekm33a] V. W. Ekman. Reviews. ICES Journal of Marine Science, 8(1):101– 102, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/101/682376.

Ekman:1933:Rb

[Ekm33b] V. W. Ekman. Reviews. ICES Journal of Marine Science, 8(3): 395–397, December 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/3/395/730333.

Ekman:1934:R

[Ekm34] V. W. Ekman. Reviews. ICES Journal of Marine Science, 9(1):102– 104, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/102/944953.

Elmhirst:1930:R

[Elm30] Richard Elmhirst. Reviews. ICES Journal of Marine Science, 5(1): 107–108, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/107/676260.

Ertel:1934:R

[Ert34] H. Ertel. Reviews. ICES Journal of Marine Science, 9(3):401–402, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/401/615784.

F:1930:Rf

[F.30a] A. F. Reviews. ICES Journal of Marine Science, 5(3):415–417, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/415/648240.

F:1930:Rd

[F.30b] E. F. Reviews. ICES Journal of Marine Science, 5(3):406, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/406/648211.

F:1930:Re

[F.30c] E. F. Reviews. ICES Journal of Marine Science, 5(3):414–415, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/414/648232.

F:1930:Ra

[F.30d] G. P. F. Reviews. ICES Journal of Marine Science, 5(1):105–106, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/105/676252.

F:1930:Rb

[F.30e] G. P. F. Reviews. ICES Journal of Marine Science, 5(1):106–107, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/106/676256.

F:1930:Rc

[F.30f] G. P. F. Reviews. ICES Journal of Marine Science, 5(2):258–260, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/258/627872.

F:1931:Ra

[F.31a] G. P. F. Reviews. ICES Journal of Marine Science, 6(1):116–117, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/116/641613.

F:1931:Rb

[F.31b] G. P. F. Reviews. ICES Journal of Marine Science, 6(1):124–125, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/124/641648.

F:1931:Rc

[F.31c] G. P. F. Reviews. ICES Journal of Marine Science, 6(3):487–488, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/487/724557.

F:1931:Rd

[F.31d] G. P. F. Reviews. ICES Journal of Marine Science, 6(3):492–493, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/492/724595.

F:1932:Ra

[F.32a] G. P. F. Reviews. ICES Journal of Marine Science, 7(1):92–93, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/92/947096.

F:1932:Rb

[F.32b] G. P. F. Reviews. ICES Journal of Marine Science, 7(2):312–313, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/312/708372.

F:1933:Ra

[F.33a] G. P. F. Reviews. ICES Journal of Marine Science, 8(2):266–267, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/266/618408.

F:1933:Rb

[F.33b] G. P. F. Reviews. ICES Journal of Marine Science, 8(2):273–274, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/273/618420.

F:1934:R

[F.34] G. P. F. Reviews. ICES Journal of Marine Science, 9(1):116–117, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/116/944996.

F:1935:Rb

[F.35a] G. P. F. Reviews. ICES Journal of Marine Science, 10(1):116–117, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/116/615442.

F:1935:Rc

[F.35b] G. P. F. Reviews. ICES Journal of Marine Science, 10(2):199–200, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/199/672881.

F:1935:Rd

[F.35c] G. P. F. Reviews. ICES Journal of Marine Science, 10(2):212, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
212/672917.

F:1935:Ra

[F.35d] L. F. Reviews. ICES Journal of Marine Science, 10(1):115–116, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
115/615440.

F:1936:Ra

[F.36a] G. P. F. Reviews. ICES Journal of Marine Science, 11(2):271–272, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/2/271/663246.

F:1936:Rb

[F.36b] G. P. F. Reviews. ICES Journal of Marine Science, 11(2):272–273, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/2/272/663247.

F:1936:Rc

[F.36c] G. P. F. Reviews. ICES Journal of Marine Science, 11(3):372, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/3/
372/621016.

F:1937:Ra

[F.37a] G. P. F. Reviews. ICES Journal of Marine Science, 12(1):83, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
83/626398.

F:1937:Rb

[F.37b] G. P. F. Reviews. ICES Journal of Marine Science, 12(1):83–84, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
83/626405.

F:1937:Rc

[F.37c] L. F. Reviews. ICES Journal of Marine Science, 12(3):366–367, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/3/366/645481.

F:1938:Rd

[F.38a] F. C. F. Reviews. ICES Journal of Marine Science, 13(2):261, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
261/784095.

F:1938:Ra

[F.38b] G. P. F. Reviews. ICES Journal of Marine Science, 13(2):247–248, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/247/784008.

F:1938:Rb

[F.38c] G. P. F. Reviews. ICES Journal of Marine Science, 13(2):248–249, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/248/784014.

F:1938:Rc

[F.38d] G. P. F. Reviews. ICES Journal of Marine Science, 13(2):249–250, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/249/784021.

F:1938:Re

[F.38e] G. P. F. Reviews. ICES Journal of Marine Science, 13(3):394– 395, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/394/635489.

F:1938:Rf

[F.38f] G. P. F. Reviews. ICES Journal of Marine Science, 13(3):395, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
395/635490.

F:1939:R

[F.39] J. H. F. Reviews. ICES Journal of Marine Science, 14(2):288–289, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
288/737187.

Fage:1933:R

[Fag33] L. Fage. Reviews. ICES Journal of Marine Science, 8(1):115–116, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/115/682412.

Farran:1936:MHD

[Far36] G. P. Farran. On the mesh of herring drift-nets in relation to the condition factor of the fish. ICES Journal of Marine Science, 11 (1):43–52, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/43/640285.

Farran:1938:SNO

[Far38] G. P. Farran. On the size and number of the ova of Irish herrings. ICES Journal of Marine Science, 13(1):91–100, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/13/1/91/714638.

Fraser-Brunner:1934:ANC

[FB34] A. Fraser-Brunner. An arrangement of nets for capturing small oceanic fishes. ICES Journal of Marine Science, 9(1):46–48, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/9/1/
46/945117.

Fjeldstad:1934:Ra

[Fje34a] J. E. Fjeldstad. Reviews. ICES Journal of Marine Science, 9(3):402, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/402/615787.

Fjeldstad:1934:Rb

[Fje34b] J. E. Fjeldstad. Reviews. ICES Journal of Marine Science, 9(3):402, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/402/615789.

Fjeldstad:1934:Rc

[Fje34c] J. E. Fjeldstad. Reviews. ICES Journal of Marine Science, 9(3): 403–404, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/403/615791.

Fleming:1939:CDP

[Fle39] Richard H. Fleming. The control of diatom populations by grazing. ICES Journal of Marine Science, 14(2):210–227, August 1939. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/14/2/210/737099.

Ford:1933:NVH

[For33] E. Ford. The “number of vertebrae” in the herring and its variation. ICES Journal of Marine Science, 8(2):211–222, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/8/2/211/618357.

Fraenkel:1934:R

[Fra34] G. Fraenkel. Reviews. ICES Journal of Marine Science, 9(3):408–409, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/408/615800.

Fraser:1937:DCS

[Fra37] James H. Fraser. The distribution of Chaetognatha in Scottish waters during 1936, with notes on the Scottish indicator species. ICES Journal of Marine Science, 12(3):311–320, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/3/311/645337.

Fraser:1939:DCS

[Fra39] James H. Fraser. The distribution of Chaetognatha in Scottish waters in 1937. ICES Journal of Marine Science, 14(1):25–34, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
25/888497.

G:1930:R

[G.30] M. G. Reviews. ICES Journal of Marine Science, 5(2):242, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/242/627843.

G:1931:Ra

[G.31a] A. C. G. Reviews. ICES Journal of Marine Science, 6(1):117–119, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/117/641618.

G:1931:Rd

[G.31b] J. G. Reviews. ICES Journal of Marine Science, 6(2):333, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/333/600664.

G:1931:Re

[G.31c] J. G. Reviews. ICES Journal of Marine Science, 6(2):333–334, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/333/600670.

G:1931:Rb

[G.31d] M. G. Reviews. ICES Journal of Marine Science, 6(2):313–314, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/313/600581.

G:1931:Rc

[G.31e] M. G. Reviews. ICES Journal of Marine Science, 6(2):317–319, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/317/600615.

G:1931:Rf

[G.31f] M. G. Reviews. ICES Journal of Marine Science, 6(3):495–496, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/495/724602.

G:1931:Rg

[G.31g] M. G. Reviews. ICES Journal of Marine Science, 6(3):503–504, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/503/724613.

G:1932:Ra

[G.32a] A. C. G. Reviews. ICES Journal of Marine Science, 7(2):293–295, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/293/708286.

G:1932:Rb

[G.32b] A. C. G. Reviews. ICES Journal of Marine Science, 7(2):297–300, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/297/708311.

G:1932:Rc

[G.32c] A. C. G. Reviews. ICES Journal of Marine Science, 7(3):427–429, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/427/694311.

G:1933:R

[G.33] M. G. Reviews. ICES Journal of Marine Science, 8(2):274–275, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/274/618422.

G:1934:Ra

[G.34a] A. C. G. Reviews. ICES Journal of Marine Science, 9(1):117–119, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/117/945006.

G:1934:Re

[G.34b] A. C. G. Reviews. ICES Journal of Marine Science, 9(3):416–418, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/416/615814.

G:1934:Rb

[G.34c] M. G. Reviews. ICES Journal of Marine Science, 9(1):124–125, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/124/945022.

G:1934:Rc

[G.34d] M. G. Reviews. ICES Journal of Marine Science, 9(1):125–127, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/125/945032.

G:1934:Rd

[G.34e] M. G. Reviews. ICES Journal of Marine Science, 9(1):127, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/9/1/
127/945040.

G:1934:Rf

[G.34f] M. G. Reviews. ICES Journal of Marine Science, 9(3):418–419, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/418/615817.

G:1935:Rc

[G.35a] A. C. G. Reviews. ICES Journal of Marine Science, 10(2):200–201, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/200/672883.

G:1935:Rd

[G.35b] A. C. G. Reviews. ICES Journal of Marine Science, 10(2):201–202, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/201/672886.

G:1935:Re

[G.35c] A. C. G. Reviews. ICES Journal of Marine Science, 10(2):202–203, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/202/672890.

G:1935:Rf

[G.35d] A. C. G. Reviews. ICES Journal of Marine Science, 10(2):203–204, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/203/672897.

G:1935:Rg

[G.35e] A. C. G. Reviews. ICES Journal of Marine Science, 10(2):204–205, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/204/672899.

G:1935:Rk

[G.35f] A. C. G. Reviews. ICES Journal of Marine Science, 10(3):329– 330, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/329/659257.

G:1935:Ra

[G.35g] M. G. Reviews. ICES Journal of Marine Science, 10(1):117–118, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
117/615445.

G:1935:Rh

[G.35h] M. G. Reviews. ICES Journal of Marine Science, 10(2):210, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
210/672914.

G:1935:Ri

[G.35i] M. G. Reviews. ICES Journal of Marine Science, 10(2):216, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
216/672928.

G:1935:Rl

[G.35j] M. G. Reviews. ICES Journal of Marine Science, 10(3):343, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
343/659291.

G:1935:Rb

[G.35k] R. G. Reviews. ICES Journal of Marine Science, 10(1):126, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
126/615470.

G:1935:Rj

[G.35l] S. G. Reviews. ICES Journal of Marine Science, 10(3):328–329, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/3/328/659254.

G:1936:RS

[G.36a] A. C. G. Rowland Southern. (1882–1935). ICES Journal of Marine Science, 11(1):3–6, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/11/1/3/640278.

G:1936:Ra

[G.36b] M. G. Reviews. ICES Journal of Marine Science, 11(1):115–116, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
115/640264.

G:1936:Rb

[G.36c] M. G. Reviews. ICES Journal of Marine Science, 11(2):273, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
273/663248.

G:1936:Rd

[G.36d] M. G. Reviews. ICES Journal of Marine Science, 11(3):378–379, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/3/378/621044.

G:1936:Rc

[G.36e] R. G. Reviews. ICES Journal of Marine Science, 11(3):373, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/3/
373/621024.

G:1937:Rb

[G.37a] M. G. Reviews. ICES Journal of Marine Science, 12(2):225, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
225/753906.

G:1937:Ra

[G.37b] R. G. Reviews. ICES Journal of Marine Science, 12(2):212, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
212/753856.

G:1938:Ra

[G.38a] M. G. Reviews. ICES Journal of Marine Science, 13(1):123–124, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
123/714559.

G:1938:Rb

[G.38b] M. G. Reviews. ICES Journal of Marine Science, 13(2):255–256, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
255/784064.

G:1939:Ra

[G.39a] A. C. G. Reviews. ICES Journal of Marine Science, 14(1):120–122, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/120/888367.

G:1939:Re

[G.39b] A. C. G. Reviews. ICES Journal of Marine Science, 14(3):418– 420, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/418/721770.

G:1939:Rd

[G.39c] E. R. G. Reviews. ICES Journal of Marine Science, 14(2):279–280, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/2/279/737148.

G:1939:Rf

[G.39d] J. L. G. Reviews. ICES Journal of Marine Science, 14(3):425– 426, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/425/721799.

G:1939:Rb

[G.39e] M. G. Reviews. ICES Journal of Marine Science, 14(1):126–128, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
126/888397.

G:1939:Rc

[G.39f] M. G. Reviews. ICES Journal of Marine Science, 14(1):128–130, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
128/888402.

Gardiner:1935:R

[Gar35] J. Stanley Gardiner. Reviews. ICES Journal of Marine Science, 10 (1):111–113, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/111/615431.

Gardiner:1937:PPP

[Gar37] A. C. Gardiner. Phosphate production by planktonic animals. ICES Journal of Marine Science, 12(2):144–146, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/2/144/753746.

Gessner:1933:PNP

[Ges33] Fritz Gessner. Phosphat, Nitrat und Planktongehalt im Arkonabecken: Ein Beitrag zur Produktionsbiologie der Ostsee. (German) [Phosphate, nitrate and plankton content in the Arkona Basin: a contribution to the production biology of the Baltic Sea]. ICES Journal of Marine Science, 8(2):181–194, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/8/2/181/618349.

Gibbons:1937:CPS

[GF37a] S. G. Gibbons and J. H. Fraser. The centrifugal pump and suction hose as a method of collecting plankton samples. ICES Journal of Marine Science, 12(2):155–170, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/2/155/753758.

Gibbons:1937:EHP

[GF37b] S. G. Gibbons and J. H. Fraser. Experiments with the “Hardy” plankton indicator in Scottish waters. I. ICES Journal of Marine Science, 12(1):45–50, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/45/626362.

Gibbons:1933:NTS

[Gib33] S. G. Gibbons. A new type of sampler for use in the quantitative analysis of plankton collections. ICES Journal of Marine Science, 8 (2):195–200, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/195/618351.

Gibbons:1937:R

[Gib37] S. G. Gibbons. Reviews. ICES Journal of Marine Science, 12(3): 351–352, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/351/645410.

Gibbons:1939:HN

[Gib39] S. G. Gibbons. The Hensen net. ICES Journal of Marine Science, 14 (2):242–248, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/242/737116.

Gilson:1931:R

[Gil31] Gilson. Reviews. ICES Journal of Marine Science, 6(3):465–466, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/465/724491.

Gilson:1933:RC

[Gil33] G. Gilson. Revue critique. ICES Journal of Marine Science, 8(2): 250–256, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/250/618368.

Gilson:1934:R

[Gil34] G. Gilson. Reviews. ICES Journal of Marine Science, 9(2):278–280, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/278/743431.

Godsil:1936:TT

[God36] H. C. Godsil. Tuna tagging. ICES Journal of Marine Science, 11 (1):94–97, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/94/640302.

Godsil:1938:TT

[God38] H. C. Godsil. Tuna tags. ICES Journal of Marine Science, 13(2): 217–220, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/217/783960.

Graham:1930:R

[Gra30] Michael Graham. Reviews. ICES Journal of Marine Science, 5(1): 112–113, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/112/676294.

Graham:1931:SPH

[Gra31a] Michael Graham. Some problems in herring behaviour: (the substance af a paper read to the Challenger Society in London on January 29th, 1931). ICES Journal of Marine Science, 6(2):252–265, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/252/600442.

Gran:1931:CHO

[Gra31b] H. H. Gran. Carl Hansen Ostenfeld 1873–1931. ICES Journal of Marine Science, 6(2):171–174, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/2/171/600421.

Gran:1932:PMP

[Gra32] H. H. Gran. Phytoplankton. methods and problems. ICES Journal of Marine Science, 7(3):343–358, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/7/3/343/694217.

Graham:1934:NSC

[Gra34] Michael Graham. The North Sea cod. ICES Journal of Marine Science, 9(2):159–171, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/9/2/159/743116.

Graham:1935:LE

[Gra35a] Michael Graham. Letter to the Editor. ICES Journal of Marine Science, 10(3):318, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/10/3/318/659221.

Graham:1935:MTE

[Gra35b] Michael Graham. Modern theory of exploiting a fishery, and application to North Sea trawling. ICES Journal of Marine Science, 10(3): 264–274, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/264/659215.

Graham:1936:LP

[Gra36] Michael Graham. The “Lowestoft photometer”. ICES Journal of Marine Science, 11(2):205–210, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/11/2/205/663208.

Graham:1938:E

[Gra38a] M. Graham. Errata. ICES Journal of Marine Science, 13(2):150, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
150/783901.

Graham:1938:RFN

[Gra38b] Michael Graham. Rates of fishing and natural mortality from the data of marking experiments. ICES Journal of Marine Science, 13 (1):76–90, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/76/714636.

Gripenberg:1937:CCB

[Gri37a] Stina Gripenberg. The calcium content of Baltic water. ICES Journal of Marine Science, 12(3):293–304, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/3/293/645330.

Gripenberg:1937:SMD

[Gri37b] Stina Gripenberg. A simplified method for the determination of calcium in sea water. ICES Journal of Marine Science, 12(3):284– 292, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/284/645328.

H:1930:R

[H.30] W. C. H. Reviews. ICES Journal of Marine Science, 5(1):110–112, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/110/676290.

H:1931:Rc

[H.31a] C. F. H. Reviews. ICES Journal of Marine Science, 6(2):319–321, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/319/600619.

H:1931:Rd

[H.31b] C. F. H. Reviews. ICES Journal of Marine Science, 6(2):331–332, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/331/600654.

H:1931:Re

[H.31c] E. H. Reviews. ICES Journal of Marine Science, 6(3):478–479, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/478/724531.

H:1931:Ra

[H.31d] H. W. H. Reviews. ICES Journal of Marine Science, 6(1):104, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/104/641575.

H:1931:Rb

[H.31e] H. W. H. Reviews. ICES Journal of Marine Science, 6(2):298–299, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/298/600471.

H:1931:Rf

[H.31f] W. C. H. Reviews. ICES Journal of Marine Science, 6(3):493–495, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/493/724598.

H:1932:Rf

[H.32a] C. F. H. Reviews. ICES Journal of Marine Science, 7(3):429–430, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/429/694316.

H:1932:Rb

[H.32b] E. H. Reviews. ICES Journal of Marine Science, 7(1):93–94, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/7/1/
93/947106.

H:1932:Rc

[H.32c] E. H. Reviews. ICES Journal of Marine Science, 7(2):296–297, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/296/708305.

H:1932:Ra

[H.32d] H. W. H. Reviews. ICES Journal of Marine Science, 7(1):90–91, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/90/947085.

H:1932:Rd

[H.32e] W. C. H. Reviews. ICES Journal of Marine Science, 7(2):305–306, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/305/708338.

H:1932:Re

[H.32f] W. C. H. Reviews. ICES Journal of Marine Science, 7(3):425, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/425/694300.

H:1933:Rb

[H.33a] C. F. H. Reviews. ICES Journal of Marine Science, 8(2):275–276, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/275/618423.

H:1933:Rc

[H.33b] H. W. H. Reviews. ICES Journal of Marine Science, 8(3):400, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/400/730340.

H:1933:Ra

[H.33c] W. C. H. Reviews. ICES Journal of Marine Science, 8(1):114–115, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/114/682411.

H:1934:Rd

[H.34a] A. H. Reviews. ICES Journal of Marine Science, 9(2):271, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/271/743385.

H:1934:WWO

[H.34b] A. H. Wilhelm Walter Otto Mielck. (1878–1933). ICES Journal of Marine Science, 9(1):3–14, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/9/1/3/945109.

H:1934:Ra

[H.34c] C. F. H. Reviews. ICES Journal of Marine Science, 9(1):128–129, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/128/945048.

H:1934:Rb

[H.34d] E. H. Reviews. ICES Journal of Marine Science, 9(2):266–267, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/266/743332.

H:1934:Rc

[H.34e] E. H. Reviews. ICES Journal of Marine Science, 9(2):267–268, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/267/743362.

H:1934:Re

[H.34f] E. H. Reviews. ICES Journal of Marine Science, 9(3):409–411, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/409/615802.

H:1935:Rb

[H.35a] E. H. Reviews. ICES Journal of Marine Science, 10(1):106, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
106/615408.

H:1935:Rc

[H.35b] E. H. Reviews. ICES Journal of Marine Science, 10(1):107, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
107/615412.

H:1935:Rd

[H.35c] E. H. Reviews. ICES Journal of Marine Science, 10(1):107, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
107/615415.

H:1935:Re

[H.35d] E. H. Reviews. ICES Journal of Marine Science, 10(1):107, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
107/615417.

H:1935:Rf

[H.35e] E. H. Reviews. ICES Journal of Marine Science, 10(2):212–213, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
212/672918.

H:1935:Ra

[H35f] H. W. H. Reviews. ICES Journal of Marine Science, 10(1):103–104, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/103/615392.

H:1936:Rd

[H.36a] A. H. Reviews. ICES Journal of Marine Science, 11(1):113, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
113/640258.

H:1936:Ra

[H.36b] E. H. Reviews. ICES Journal of Marine Science, 11(1):102–103, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
102/640225.

H:1936:Rb

[H.36c] E. H. Reviews. ICES Journal of Marine Science, 11(1):103–104, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
103/640227.

H:1936:Rc

[H.36d] E. H. Reviews. ICES Journal of Marine Science, 11(1):105–106, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
105/640231.

H:1937:Ra

[H.37a] E. H. Reviews. ICES Journal of Marine Science, 12(1):84–85, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
84/626410.

H:1937:Rb

[H.37b] E. H. Reviews. ICES Journal of Marine Science, 12(1):85–87, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
85/626415.

H:1937:Re

[H.37c] E. H. Reviews. ICES Journal of Marine Science, 12(2):216, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
216/753875.

H:1937:Rf

[H.37d] E. H. Reviews. ICES Journal of Marine Science, 12(2):217, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
217/753876.

H:1937:Rg

[H.37e] E. H. Reviews. ICES Journal of Marine Science, 12(2):218–219, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
218/753878.

H:1937:Rh

[H.37f] E. H. Reviews. ICES Journal of Marine Science, 12(2):219, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
219/753879.

H:1937:Rd

[H.37g] H. W. H. Reviews. ICES Journal of Marine Science, 12(2):215– 216, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/215/753873.

H:1937:Rc

[H.37h] W. C. H. Reviews. ICES Journal of Marine Science, 12(1):92–93, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
92/626438.

H:1938:Rb

[H.38a] B. H. Reviews. ICES Journal of Marine Science, 13(2):272–273, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
272/784136.

H:1938:Rg

[H.38b] C. F. H. Reviews. ICES Journal of Marine Science, 13(3):397– 398, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/397/635498.

H:1938:Rh

[H.38c] C. F. H. Reviews. ICES Journal of Marine Science, 13(3):409, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
409/635530.

H:1938:Ra

[H.38d] E. H. Reviews. ICES Journal of Marine Science, 13(1):116, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
116/714502.

H:1938:Rc

[H.38e] E. H. Reviews. ICES Journal of Marine Science, 13(3):392–393, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/3/392/635475.

H:1938:Rd

[H.38f] E. H. Reviews. ICES Journal of Marine Science, 13(3):393, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
393/635478.

H:1938:Re

[H.38g] E. H. Reviews. ICES Journal of Marine Science, 13(3):393, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
393/635481.

H:1938:Rf

[H.38h] E. H. Reviews. ICES Journal of Marine Science, 13(3):393, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
393/635482.

H:1939:Rb

[H39a] Chr. H. Reviews. ICES Journal of Marine Science, 14(1):124–126, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/124/888390.

H:1939:Ra

[H.39b] E. H. Reviews. ICES Journal of Marine Science, 14(1):120, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
120/888363.

H:1939:Rc

[H.39c] E. H. Reviews. ICES Journal of Marine Science, 14(2):295–296, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
295/737252.

H:1939:Rd

[H.39d] E. H. Reviews. ICES Journal of Marine Science, 14(2):296, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
296/737255.

Hagmeier:1931:R

[Hag31] A. Hagmeier. Reviews. ICES Journal of Marine Science, 6(3):482– 484, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/482/724545.

Hagmeier:1933:Rc

[Hag33a] Hagmeier. Reviews. ICES Journal of Marine Science, 8(3):402–403, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/402/730381.

Hagmeier:1933:Ra

[Hag33b] A. Hagmeier. Reviews. ICES Journal of Marine Science, 8(2): 268–271, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/268/618412.

Hagmeier:1933:Rb

[Hag33c] A. Hagmeier. Reviews. ICES Journal of Marine Science, 8(2): 271–272, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/271/618414.

Hagmeier:1937:Ra

[Hag37a] A. Hagmeier. Reviews. ICES Journal of Marine Science, 12(2):232, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/232/753933.

Hagmeier:1937:Rb

[Hag37b] A. Hagmeier. Reviews. ICES Journal of Marine Science, 12(2): 232–233, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/232/753934.

Hamaker:1933:ROT

[Ham33] H. C. Hamaker. Results obtained with a thermograph for surface temperatures during the “Snellius” Expedition. ICES Journal of Marine Science, 8(1):64–72, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/1/64/682455.

Hans:1938:R

[Han38] Böker Hans. Reviews. ICES Journal of Marine Science, 13(3): 397, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/397/635495.

Hart:1935:NCM

[Har35a] E. G. Hart. Note concerning a measuring plankton-net. ICES Journal of Marine Science, 10(2):179–184, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/10/2/179/672835.

Hart:1935:SDM

[Har35b] E. G. Hart. Some devices for the manipulation of marine plankton collections on board ship. ICES Journal of Marine Science, 10(2): 173–178, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/173/672832.

Hart:1936:MMB

[Har36] John Lawson Hart. A modified measuring board. ICES Journal of Marine Science, 11(3):363, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/11/3/363/620997.

Hass:1939:UJP

[Has39] Günther Hass. Untersuchungen an Jungflundern (Pleuronectes flesus L) der Samlandküste. (German) [Investigations on young flounder (Pleuronectes flesus L) of the Samland coast]. ICES Journal of Marine Science, 14(2):261–270, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/2/261/737128.

Havinga:1930:GCVa

[Hav30a] B. Havinga. Der Granat (Crangon vulgaris Fabr.) in den holländischen Gewässern. (German) [The garnet (Crangon vulgaris Fabr.) in the Dutch waters]. ICES Journal of Marine Science, 5(1):55–56, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/5/1/
55/676374.

Havinga:1930:GCVb

[Hav30b] B. Havinga. Der Granat (Crangon vulgaris Fabr.) in den holländischen Gewässern. (German) [The garnet (Crangon vulgaris Fabr.) in the Dutch waters]. ICES Journal of Marine Science, 5(1):57–87, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/5/1/
57/676380.

Havinga:1931:R

[Hav31] B. Havinga. Reviews. ICES Journal of Marine Science, 6(2):326–328, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/326/600638.

Havinga:1933:R

[Hav33] B. Havinga. Reviews. ICES Journal of Marine Science, 8(2):284–285, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/284/618444.

Havinga:1934:R

[Hav34] B. Havinga. Reviews. ICES Journal of Marine Science, 9(3):425–426, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/425/615828.

Havinga:1937:Ra

[Hav37a] B. Havinga. Reviews. ICES Journal of Marine Science, 12(2): 233–234, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/233/753963.

Havinga:1937:Rb

[Hav37b] B. Havinga. Reviews. ICES Journal of Marine Science, 12(2): 234–235, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/234/753964.

Havinga:1939:PTS

[Hav39a] B. Havinga. Prediction of the time of setting of oyster spat and a method of control. ICES Journal of Marine Science, 14(3):394– 400, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/394/721642.

Havinga:1939:Ra

[Hav39b] B. Havinga. Reviews. ICES Journal of Marine Science, 14(3):422– 423, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/422/721787.

Havinga:1939:Rb

[Hav39c] B. Havinga. Reviews. ICES Journal of Marine Science, 14(3):430– 431, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/430/721809.

Heldt:1931:R

[Hel31] H. Heldt. Reviews. ICES Journal of Marine Science, 6(2):314–316, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/314/600608.

Heldt:1935:R

[Hel35] H. Heldt. Reviews. ICES Journal of Marine Science, 10(3):342– 343, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/342/659287.

Heldt:1939:R

[Hel39] H. Heldt. Reviews. ICES Journal of Marine Science, 14(1):134–135, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/134/888450.

Hentschel:1931:R

[Hen31] E. Hentschel. Reviews. ICES Journal of Marine Science, 6(1): 119–120, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/119/641622.

Henking:1932:R

[Hen32] H. Henking. Reviews. ICES Journal of Marine Science, 7(1):108–110, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/108/946931.

Hentschel:1938:QSP

[Hen38] Ernst Hentschel. Über quantitative Seihmethoden in der Planktonforschung. (German) [On quantitative screening methods in plankton research]. ICES Journal of Marine Science, 13(3):304–308, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
304/635431.

Henschel:1939:ATG

[Hen39] J. Henschel. Der Atmungsmechanismus der Teleosteer. (German) [The breathing mechanism of the Teleosteer]. ICES Journal of Marine Science, 14(2):249–260, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/2/249/737123.

Heron:1930:R

[Her30] Crawford Heron. Reviews. ICES Journal of Marine Science, 5(3): 421–422, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/421/648252.

Hessle:1931:BSR

[Hes31] Chr. Hessle. Biological statistics regarding the herrings along the Baltic coast of Sweden. ICES Journal of Marine Science, 6(1):21–27, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/21/641708.

Hickling:1931:CVD

[Hic31] C. F. Hickling. A comparison of the Vigneron–Dahl trawling gear with the steam otter trawl, in the hake fishery. ICES Journal of Marine Science, 6(3):421–431, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/3/421/724459.

Hickling:1939:SAD

[Hic39] C. F. Hickling. The selective action of the drift-net on the Cornish pilchard. ICES Journal of Marine Science, 14(1):67–80, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
67/888511.

Hjort:1939:MK

[Hjo39] Johan Hjort. N. M. Knipovich. 1862–1939. ICES Journal of Marine Science, 14(3):335–336, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/3/335/721621.

Hodgson:1933:FES

[Hod33] William C. Hodgson. Further experiments on the selective action of commercial drift nets. ICES Journal of Marine Science, 8(3):344–354, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/344/730301.

Hodgson:1935:LE

[Hod35] W. C. Hodgson. Letters to the Editor. ICES Journal of Marine Science, 10(1):99–101, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/10/1/99/615501.

Huntsman:1938:IPF

[Hun38] A. G. Huntsman. International Passamaquoddy Fishery Investigations. ICES Journal of Marine Science, 13(3):357–369, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
357/635442.

I:1934:Ra

[I.34a] C. I. Reviews. ICES Journal of Marine Science, 9(1):105–106, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/105/944958.

I:1934:Rb

[I.34b] C. I. Reviews. ICES Journal of Marine Science, 9(2):259–260, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/259/743201.

I:1935:R

[I.35] C. I. Reviews. ICES Journal of Marine Science, 10(2):190–191, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
190/672846.

I:1937:R

[I.37] C. O’D. I. Reviews. ICES Journal of Marine Science, 12(2):199– 200, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/199/753812.

I:1938:R

[I.38] C. O’D. I. Reviews. ICES Journal of Marine Science, 13(1):101–102, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/101/714438.

I:1939:R

[I.39] C. O’D. I. Reviews. ICES Journal of Marine Science, 14(2):277– 278, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/277/737141.

Ibanez:1933:NES

[Iba33a] O. Gomez Ibañez. Note on the effect of salts in the determination of phosphate in sea water by Denigès’ method. ICES Journal of Marine Science, 8(3):326–329, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/3/326/730289.

Ibanez:1933:PSW

[Iba33b] O. Gomez Ibañez. Preservation of sea water samples with fluoride for phosphate estimation by Denigès’ method. ICES Journal of Marine Science, 8(3):329–330, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/3/329/730292.

Idelson:1931:FMB

[Ide31] M. Idelson. Fish marking in the Barents Sea. ICES Journal of Marine Science, 6(3):432–433, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/3/432/724462.

Ilyin:1934:OFR

[Ily34] B. Ilyin. Organization of fishery research in the U.S.S.R. ICES Journal of Marine Science, 9(3):392–393, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/9/3/392/615770.

Iselin:1939:R

[Ise39] C. O’D. Iselin. Reviews. ICES Journal of Marine Science, 14 (1):116, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/116/887978.

J:1930:TWF

[J.30] J. T. J. Thomas Wemyss Fulton. (1855–1929). ICES Journal of Marine Science, 5(1):3–6, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/1/3/676368.

J:1931:R

[J.31] J. T. J. Reviews. ICES Journal of Marine Science, 6(3):499–500, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/499/724607.

J:1934:R

[J.34] J. T. J. Reviews. ICES Journal of Marine Science, 9(1):129–130, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/129/945053.

J:1935:R

[J.35] J. T. J. Reviews. ICES Journal of Marine Science, 10(3):335, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
335/659272.

J:1936:R

[J.36] J. T. J. Reviews. ICES Journal of Marine Science, 11(2):269–271, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
269/663245.

J:1937:R

[J.37] J. T. J. Reviews. ICES Journal of Marine Science, 12(1):88–91, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
88/626426.

Jacobsen:1930:R

[Jac30] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 5(2): 231–233, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/231/627818.

Jacobsen:1932:Ra

[Jac32a] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 7(2): 284–286, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/284/708224.

Jacobsen:1932:Rb

[Jac32b] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 7(2): 289–291, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/289/708267.

Jacobsen:1933:Ra

[Jac33a] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 8(2):260, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/260/618384.

Jacobsen:1933:Rb

[Jac33b] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 8(2): 260–262, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/260/618388.

Jacobsen:1934:R

[Jac34] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 9(2): 262–263, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/262/743235.

Jacobsen:1935:Ra

[Jac35a] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 10(2): 193–194, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/193/672861.

Jacobsen:1935:Rb

[Jac35b] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 10(2): 195–196, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/195/672864.

Jacobsen:1936:BR

[Jac36a] J. P. Jacobsen. Book reviews. ICES Journal of Marine Science, 11(3): 364–366, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/364/620999.

Jacobsen:1936:R

[Jac36b] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 11(2): 260–261, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/260/663233.

Jacobsen:1937:R

[Jac37] J. P. Jacobsen. Reviews. ICES Journal of Marine Science, 12(3): 337, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/337/645346.

Janssen:1939:TYS

[Jan39] John F. Janssen, Jr. Two years of sardine tagging in California. ICES Journal of Marine Science, 14(1):48–66, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/1/48/888507.

Jeffreys:1932:R

[Jef32] Harold Jeffreys. Reviews. ICES Journal of Marine Science, 7(3): 417–418, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/417/694277.

Jensen:1934:R

[Jen34] Aage J. C. Jensen. Reviews. ICES Journal of Marine Science, 9(2): 277–278, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/277/743424.

Jensen:1939:FRC

[Jen39a] Aage J. C. Jensen. Fluctuations in the racial characters of the plaice and the dab. ICES Journal of Marine Science, 14(3):370–384, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/3/
370/721636.

Jensen:1939:R

[Jen39b] Aage J. C. Jensen. Reviews. ICES Journal of Marine Science, 14(3): 424–425, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/424/721796.

Jespersen:1930:R

[Jes30] P. Jespersen. Reviews. ICES Journal of Marine Science, 5(3):423, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/423/648257.

Jespersen:1931:Ra

[Jes31a] P. Jespersen. Reviews. ICES Journal of Marine Science, 6(1):130–131, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/130/641669.

Jespersen:1931:Rb

[Jes31b] P. Jespersen. Reviews. ICES Journal of Marine Science, 6(2):316–317, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/316/600613.

Jespersen:1931:Rc

[Jes31c] P. Jespersen. Reviews. ICES Journal of Marine Science, 6(2):322–323, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/322/600624.

Jespersen:1932:Ra

[Jes32a] P. Jespersen. Reviews. ICES Journal of Marine Science, 7(2):310–312, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/310/708364.

Jespersen:1932:Rb

[Jes32b] P. Jespersen. Reviews. ICES Journal of Marine Science, 7(3):418– 419, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/418/694282.

Jespersen:1935:R

[Jes35] P. Jespersen. Reviews. ICES Journal of Marine Science, 10(2): 210–212, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/210/672916.

Jespersen:1937:R

[Jes37] P. Jespersen. Reviews. ICES Journal of Marine Science, 12(3): 354, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/354/645419.

Jesperaen:1939:R

[Jes39] P. Jesperaen. Reviews. ICES Journal of Marine Science, 14(1): 130–131, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/130/888443.

Johnstone:1931:R

[Joh31] Jas. Johnstone. Reviews. ICES Journal of Marine Science, 6(3): 464–465, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/464/724487.

Johnstone:1932:R

[Joh32] Jas. Johnstone. Reviews. ICES Journal of Marine Science, 7(2):284, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/284/708218.

Joubin:1932:R

[Jou32] L. Joubin. Reviews. ICES Journal of Marine Science, 7(1):111–112, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/111/946944.

Juday:1931:R

[Jud31] C. Juday. Reviews. ICES Journal of Marine Science, 6(2):311, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/311/600507.

K:1938:CG

[K.38] S. K. Charles Green 1876–1938. ICES Journal of Marine Science, 13(1):3–5, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/3/714610.

K:1939:R

[K.39] P. K. Reviews. ICES Journal of Marine Science, 14(2):295, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
295/737248.

Kandler:1932:UBW

[Kän32] Rudolf Kändler. Unsicherheiten bei Bestimmung der Wirbelzahl infolge Verwachsungserscheinungen. (German) [Uncertainties in determining the number of vortices due to adhesion phenomena]. ICES Journal of Marine Science, 7(3):373–385, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/7/3/373/694227.

Kandler:1933:Ra

[Kän33a] R. Kändler. Reviews. ICES Journal of Marine Science, 8(3):403–404, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/403/730385.

Kandler:1933:Rb

[Kän33b] R. Kändler. Reviews. ICES Journal of Marine Science, 8(3):405–406, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/405/730390.

Kandler:1934:R

[Kän34] R. Kändler. Reviews. ICES Journal of Marine Science, 9(1):109–114, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/109/944977.

Kiilerich:1932:R

[Kii32] A. Kiilerich. Reviews. ICES Journal of Marine Science, 7(3):409–410, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/409/694247.

Knudsen:1930:R

[Knu30] Martin Knudsen. Reviews. ICES Journal of Marine Science, 5(3): 399–400, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/399/648195.

Knudsen:1939:RVP

[Knu39] Martin Knudsen. S. A. R. Valdemar, Prince de Danemark, 1858–1939. ICES Journal of Marine Science, 14(1):3–4, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/1/3/888500.

Kotthaus:1936:SDB

[Kot36] Adolf Kotthaus. Die Schollenbevölkerung der Deutschen Bucht im Jahre 1935. (German) [The plaice population of the German Bight in 1935]. ICES Journal of Marine Science, 11(1):27–32, April 1936. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/11/1/27/640274.

Kramp:1934:EOR

[Kra34] P. L. Kramp. An exceptional occurrence of Rhizostoma octopus and Chrysaora hysoscella in the Danish waters in 1933. ICES Journal of Marine Science, 9(2):211–221, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/2/211/743147.

Kramp:1939:R

[Kra39] P. L. Kramp. Reviews. ICES Journal of Marine Science, 14(2): 293–295, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/293/737245.

Krey:1939:BCM

[Kre39] Johannes Krey. Die Bestimmung des Chlorophylls in Meerwasser– Schöpfproben. (German) [Determination of chlorophyll in sea water– scoop samples]. ICES Journal of Marine Science, 14(2):201–209, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
201/737091.

Kunne:1931:R

[Kün31] Künne. Reviews. ICES Journal of Marine Science, 6(3):482, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/482/724543.

KUNne:1932:R

[KÜN32] KÜNne. Reviews. ICES Journal of Marine Science, 7(2):295, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/295/708297.

KuUnne:1932:R

[KüU32] KüUnne. Reviews. ICES Journal of Marine Science, 7(1):91–92, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/91/947093.

Kreps:1930:SCP

[KV30] E. Kreps and N. Verjbinskaya. Seasonal changes in the phosphate and nitrate content and in hydrogen ion concentration in the Barents Sea. ICES Journal of Marine Science, 5(3):329–346, December 1930. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/5/3/329/648149.

Kreps:1932:CNS

[KV32] E. Kreps and N. Verjbinskaya. The consumption of nutrient salts in the Barents Sea. ICES Journal of Marine Science, 7(1):25–46, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/25/946980.

L:1930:Rd

[L.30a] H. L. Reviews. ICES Journal of Marine Science, 5(3):410–411, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/410/648219.

L:1930:Ra

[L.30b] J. R. L. Reviews. ICES Journal of Marine Science, 5(2):239–240, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/239/627837.

L:1930:Rc

[L.30c] M. V. L. Reviews. ICES Journal of Marine Science, 5(2):244–245, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/244/627851.

L:1931:Rb

[L.31a] H. L. Reviews. ICES Journal of Marine Science, 6(1):125–127, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/125/641654.

L:1931:Ra

[L.31b] J. R. L. Reviews. ICES Journal of Marine Science, 6(1):104–105, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/104/641578.

L:1931:Rc

[L.31c] M. V. L. Reviews. ICES Journal of Marine Science, 6(1):140–141, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/140/641697.

L:1932:Ra

[L.32a] H. L. Reviews. ICES Journal of Marine Science, 7(1):110, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/7/1/
110/946937.

L:1932:Rc

[L.32b] H. L. Reviews. ICES Journal of Marine Science, 7(2):308–309, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/308/708350.

L:1932:Rb

[L.32c] M. V. L. Reviews. ICES Journal of Marine Science, 7(1):113, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/113/946951.

L:1932:Rd

[L.32d] M. V. L. Reviews. ICES Journal of Marine Science, 7(3):419–420, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/419/694285.

L:1932:Re

[L.32e] M. V. L. Reviews. ICES Journal of Marine Science, 7(3):421, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/421/694288.

L:1933:Rd

[L.33a] J. L. Reviews. ICES Journal of Marine Science, 8(2):283–284, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/283/618443.

L:1933:Ra

[L.33b] J. R. L. Reviews. ICES Journal of Marine Science, 8(1):105–106, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/105/682382.

L:1933:Rb

[L.33c] M. V. L. Reviews. ICES Journal of Marine Science, 8(1):110, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/110/682395.

L:1933:Rc

[L.33d] M. V. L. Reviews. ICES Journal of Marine Science, 8(1):116–117, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/116/682415.

L:1934:Rb

[L.34a] J. R. L. Reviews. ICES Journal of Marine Science, 9(2):262, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/262/743228.

L:1934:Ra

[L.34b] M. V. L. Reviews. ICES Journal of Marine Science, 9(1):115–116, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/115/944985.

L:1934:Rc

[L.34c] M. V. L. Reviews. ICES Journal of Marine Science, 9(3):411–412, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/411/615806.

L:1935:Ra

[L.35a] M. V. L. Reviews. ICES Journal of Marine Science, 10(1):107–108, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/107/615420.

L:1935:Rb

[L.35b] M. V. L. Reviews. ICES Journal of Marine Science, 10(2):198–199, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/198/672878.

L:1936:R

[L.36] M. V. L. Reviews. ICES Journal of Marine Science, 11(1):104–105, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/104/640229.

L:1937:Ra

[L.37a] M. V. L. Reviews. ICES Journal of Marine Science, 12(1):91, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
91/626432.

L:1937:Rb

[L.37b] M. V. L. Reviews. ICES Journal of Marine Science, 12(2):212–213, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/212/753860.

L:1937:Rc

[L.37c] M. V. L. Reviews. ICES Journal of Marine Science, 12(2):213–214, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/213/753866.

L:1937:Rd

[L.37d] M. V. L. Reviews. ICES Journal of Marine Science, 12(2):214, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
214/753870.

L:1937:Rg

[L.37e] M. V. L. Reviews. ICES Journal of Marine Science, 12(3):349– 350, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/349/645406.

L:1937:Re

[L.37f] W. V. L. Reviews. ICES Journal of Marine Science, 12(3):345– 347, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/345/645392.

L:1937:Rf

[L.37g] W. V. L. Reviews. ICES Journal of Marine Science, 12(3):347– 348, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/347/645397.

L:1938:Ra

[L.38a] M. V. L. Reviews. ICES Journal of Marine Science, 13(1):116–117, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/116/714506.

L:1938:Rb

[L.38b] M. V. L. Reviews. ICES Journal of Marine Science, 13(1):117, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
117/714528.

L:1939:R

[L.39] M. V. L. Reviews. ICES Journal of Marine Science, 14(2):289–290, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/2/289/737236.

Lb:1938:R

[Lb.38] Lb. Reviews. ICES Journal of Marine Science, 13(1):120–121, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
120/714550.

Lunde:1930:ICA

[LBC30] Gulbrand Lunde, Jens Böe, and Karl Closs. Iodine content of American marine animals. ICES Journal of Marine Science, 5(2):216–225, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/216/627808.

Lebour:1933:ILM

[Leb33] Marie V. Lebour. The importance of larval Mollusca in the plankton. ICES Journal of Marine Science, 8(3):335–343, December 1933. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/8/3/335/730296.

Lees:1930:PGC

[Lee30] H. Lees. Paper and glass collectors for oyster spat. ICES Journal of Marine Science, 5(3):383–384, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/3/383/648164.

Lewis:1934:PCO

[Lew34] B. A. H. Mabel Lewis. The physical and chemical oceanography of the Irish Sea. a bibliography. ICES Journal of Marine Science, 9(1):87– 92, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/87/945130.

L:1930:Rb

[LH30] M. V. L. and H. W. H. Reviews. ICES Journal of Marine Science, 5 (2):243–244, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/243/627845.

Lissner:1930:VBS

[Lis30] H. Lissner. Ein verbesserter Becher für das Standard-Netz. (German) [An improved mug for the standard network]. ICES Journal of Marine Science, 5(2):192–193, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/2/192/627797.

Lissner:1934:RH

[Lis34] H. Lissner. On races of herrings. ICES Journal of Marine Science, 9(3):346–364, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/9/3/346/615756.

Lisitzin:1938:ZZW

[Lis38] Eugenie Lisitzin. Über den Zusammenhang zwischen Wind und Strom bei dem Feuerschiff Storbrotten im nördlichen Ålandsmeer. (German)

[On the connection between wind and electricity at the lightship Storbrotten in the northern Ålands Sea]. ICES Journal of Marine Science, 13(3):293–303, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/3/293/635428.

Lubbert:1931:PEA

[Lüb31] H. Lübbert. Professor Ehrenbaum’s Ausscheiden aus dem Staatsdienst. (German) [Professor Ehrenbaum’s resignation from the Civil Service]. ICES Journal of Marine Science, 6(3):458–459, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/458/724477.

Lucas:1936:CIR

[Luc36a] C. E. Lucas. On certain inter-relations between phytoplankton and zooplankton under experimental conditions. ICES Journal of Marine Science, 11(3):343–362, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/11/3/343/620994.

Lucas:1936:DVS

[Luc36b] C. E. Lucas. On the diurnal variation of size-groups of trawlcaught herring. ICES Journal of Marine Science, 11(1):53–59, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
53/640290.

Lucas:1938:Ra

[Luc38a] C. E. Lucas. Reviews. ICES Journal of Marine Science, 13(2): 250–251, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/250/784025.

Lucas:1938:Rb

[Luc38b] C. E. Lucas. Reviews. ICES Journal of Marine Science, 13(2): 252–253, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/252/784033.

Lucas:1938:SAI

[Luc38c] C. E. Lucas. Some aspects of integration in plankton communities. ICES Journal of Marine Science, 13(3):309–322, December 1938. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/13/3/309/635433.

Lucas:1939:R

[Luc39] C. E. Lucas. Reviews. ICES Journal of Marine Science, 14(3):420– 421, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/420/721773.

M:1930:R

[M.30] W. J. M. M. Reviews. ICES Journal of Marine Science, 5(2):261–262, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/261/627878.

M:1931:WCM

[M.31a] A. M. William Carmichael M’Intosh. (1838–1931). ICES Journal of Marine Science, 6(2):165–169, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/2/165/600420.

M:1931:Rd

[M.31b] D. J. M. Reviews. ICES Journal of Marine Science, 6(1):109–110, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/109/641597.

M:1931:Re

[M.31c] D. J. M. Reviews. ICES Journal of Marine Science, 6(1):110–111, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/110/641600.

M:1931:Rf

[M.31d] D. J. M. Reviews. ICES Journal of Marine Science, 6(2):290–291, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/290/600459.

M:1931:Ra

[M.31e] H. A. M. Reviews. ICES Journal of Marine Science, 6(1):105–106, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/105/641583.

M:1931:Rb

[M.31f] H. A. M. Reviews. ICES Journal of Marine Science, 6(1):106–108, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/106/641587.

M:1931:Rc

[M.31g] H. A. M. Reviews. ICES Journal of Marine Science, 6(1):108–109, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/108/641591.

M:1931:DTJ

[M.31h] W. L. M. David Thomas Jones. (1866–1931). ICES Journal of Marine Science, 6(2):161–164, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/2/161/600419.

M:1932:Rb

[M.32a] D. J. M. Reviews. ICES Journal of Marine Science, 7(3):410–413, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/410/694252.

M:1932:Rc

[M.32b] D. J. M. Reviews. ICES Journal of Marine Science, 7(3):413–414, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/413/694257.

M:1932:Ra

[M.32c] H. A. M. Reviews. ICES Journal of Marine Science, 7(1):87–89, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/87/947076.

M:1932:Rd

[M.32d] N. A. M. Reviews. ICES Journal of Marine Science, 7(3):426–427, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/426/694307.

M:1933:Ra

[M.33a] A. M. Reviews. ICES Journal of Marine Science, 8(1):111–112, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/111/682401.

M:1933:Rb

[M.33b] A. M. Reviews. ICES Journal of Marine Science, 8(3):401–402, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/401/730378.

M:1934:Rb

[M.34a] D. J. M. Reviews. ICES Journal of Marine Science, 9(2):255, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/255/743175.

M:1934:Rc

[M.34b] D. J. M. Reviews. ICES Journal of Marine Science, 9(2):256, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/256/743182.

M:1934:Ra

[M.34c] S. M. M. Reviews. ICES Journal of Marine Science, 9(1):114–115, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/114/944980.

M:1935:Rc

[M.35a] A. M. Reviews. ICES Journal of Marine Science, 10(3):333, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
333/659268.

M:1935:Ra

[M.35b] H. B. M. Reviews. ICES Journal of Marine Science, 10(3):331– 332, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/331/659263.

M:1935:Rb

[M.35c] H. B. M. Reviews. ICES Journal of Marine Science, 10(3):332– 333, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/332/659266.

M:1937:Ra

[M.37a] H. B. M. Reviews. ICES Journal of Marine Science, 12(2):228– 229, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/228/753924.

M:1937:Rb

[M.37b] H. B. M. Reviews. ICES Journal of Marine Science, 12(3):364, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/3/364/645465.

M:1937:Rc

[M.37c] H. B. M. Reviews. ICES Journal of Marine Science, 12(3):364– 365, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/364/645470.

M:1937:Rd

[M.37d] H. B. M. Reviews. ICES Journal of Marine Science, 12(3):365– 366, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/365/645476.

M:1938:Ra

[M.38a] S. M. M. Reviews. ICES Journal of Marine Science, 13(1):118, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
118/714541.

M:1938:Rb

[M.38b] S. M. M. Reviews. ICES Journal of Marine Science, 13(1):118–120, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/118/714544.

M:1938:Rc

[M.38c] S. M. M. Reviews. ICES Journal of Marine Science, 13(3):394, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/3/394/635485.

M:1939:Ra

[M.39a] C. H. M. Reviews. ICES Journal of Marine Science, 14(1):122–123, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/122/888378.

M:1939:Rb

[M.39b] J. W. S. M. Reviews. ICES Journal of Marine Science, 14(2): 292–293, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/292/737241.

M:1939:Rc

[M.39c] W. J. M. M. Reviews. ICES Journal of Marine Science, 14(2): 307–308, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/307/737356.

M:1939:Rd

[M.39d] W. J. M. M. Reviews. ICES Journal of Marine Science, 14(2): 308–309, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/308/737360.

Mannsfeld:1937:SPP

[Man37] Wilh. Mannsfeld. Über die Scholle (Pleuronectes platessa) an der lettländischen Ostseeküste. (German) [On plaice (Pleuronectes platessa) on the Latvian Baltic coast]. ICES Journal of Marine Science, 12(2):182–191, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/2/182/753802.

Marmer:1932:CT

[Mar32] H. A. Marmer. On the classification of tides. ICES Journal of Marine Science, 7(1):11–24, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/1/11/946934.

Matthews:1930:Ra

[Mat30a] Donald J. Matthews. Reviews. ICES Journal of Marine Science, 5 (1):101–102, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/101/676236.

Matthews:1930:Rb

[Mat30b] Donald J. Matthews. Reviews. ICES Journal of Marine Science, 5 (1):102–103, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/102/676242.

Maurice:1939:JOB

[Mau39] H. G. Maurice. John Oliver Borley. 1872–1938. ICES Journal of Marine Science, 14(1):5–6, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/1/5/888508.

Mitchell:1936:EST

[MBR36] Philip H. Mitchell, Kurt Buch, and Norris W. Rakestraw. The effect of salinity and temperature upon the dissociation of cresol red and phenol red in sea water. ICES Journal of Marine Science, 11(2): 183–189, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/183/663202.

McEwen:1933:R

[McE33] Geo. F. McEwen. Reviews. ICES Journal of Marine Science, 8(1): 106–108, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/106/682386.

Meek:1930:R

[Mee30] Alexander Meek. Reviews. ICES Journal of Marine Science, 5(2): 270–271, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/270/627907.

Menzies:1932:SMS

[Men32] W. J. M. Menzies. The spawning mark on sea trout scales. ICES Journal of Marine Science, 7(1):81–84, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/7/1/81/947032.

Menzies:1934:MF

[Men34] W. J. M. Menzies. Migration of fishes. ICES Journal of Marine Science, 9(1):94–96, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/9/1/94/945134.

Menzies:1936:RER

[Men36] W. J. M. Menzies. The run of elvers in the River Bann, Northern Ireland. ICES Journal of Marine Science, 11(2):249–259, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
249/663225.

Montgomery:1939:R

[MI39] R. B. Montgomery and C. O’D. Iselin. Reviews. ICES Journal of Marine Science, 14(1):114–116, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/1/114/887975.

Mielck:1933:R

[Mie33] W. Mielck. Reviews. ICES Journal of Marine Science, 8(1):117–119, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/117/682417.

M:1931:Rg

[MO31] S. M. M. and A. P. O. Reviews. ICES Journal of Marine Science, 6(2):304, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/304/600490.

Mohr:1930:BAW

[Moh30] Erna W. Mohr. Bibliographie der Alters- und Wachstums-Bestimmung bei Fischen: II. Nachträge und Fortsetzung. (German) [Bibliography of age and growth determination in fish: II. Supplements and continuation]. ICES Journal of Marine Science, 5(1):88– 100, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/88/676392.

Mohr:1934:BAW

[Moh34] Erna W. Mohr. Bibliographie der Alters- und Wachstums-Bestimmung bei Fischen. III. Nachträge und Fortsetzung. (German)

[Bibliography of age and growth determination in fish: III. Supplements and continuation]. ICES Journal of Marine Science, 9(3): 377–391, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/377/615763.

Molander:1935:Ra

[Mol35a] Molander. Reviews. ICES Journal of Marine Science, 10(1):119–120, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/119/615450.

Molander:1935:Rb

[Mol35b] Molander. Reviews. ICES Journal of Marine Science, 10(3):333– 334, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/333/659270.

Mortensen:1931:R

[Mor31] Th. Mortensen. Reviews. ICES Journal of Marine Science, 6(2): 301–303, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/301/600487.

Mosby:1937:Ra

[Mos37a] Håkon Mosby. Reviews. ICES Journal of Marine Science, 12(2): 202–203, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/202/753823.

Mosby:1937:Rb

[Mos37b] Håkon Mosby. Reviews. ICES Journal of Marine Science, 12(2): 203–204, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/203/753827.

Mosby:1938:R

[Mos38] Håkon Mosby. Reviews. ICES Journal of Marine Science, 13(2): 242, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/242/783984.

Mosby:1939:R

[Mos39] Håkon Mosby. Reviews. ICES Journal of Marine Science, 14(2): 278–279, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/278/737144.

Mottram:1931:SMS

[Mot31] M. B. Ldn. J. C. Mottram. The spawning mark on sea trout scales. ICES Journal of Marine Science, 6(3):434–437, December 1931. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/6/3/434/724463.

Moult:1939:LAR

[Mou39] Frank H. Moult. Laboratory apparatus for the rearing of young salmon from freshly fertilized eggs. ICES Journal of Marine Science, 14(2): 271–273, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/271/737133.

Mitchell:1935:DCC

[MT35] Philip H. Mitchell and Ivon R. Taylor. The dissociation constant of cresol red in sea water. ICES Journal of Marine Science, 10(2): 169–172, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/169/672830.

Murray:1932:TTT

[Mur32] F. Z. S. D. K. Wolfe Murray. Tunny (Thunnus thynnus L.) in the North Sea. ICES Journal of Marine Science, 7(2):251–254, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/251/708205.

N:1930:Ra

[N.30a] J. R. N. Reviews. ICES Journal of Marine Science, 5(1):122–123, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/122/676338.

N:1930:Rb

[N.30b] J. R. N. Reviews. ICES Journal of Marine Science, 5(2):260–261, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/260/627875.

N:1931:R

[N.31] J. R. N. Reviews. ICES Journal of Marine Science, 6(1):131, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/6/1/
131/641672.

N:1934:R

[N.34] J. R. N. Reviews. ICES Journal of Marine Science, 9(3):423–424, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/423/615824.

N:1938:Rb

[N.38a] J. R. N. Reviews. ICES Journal of Marine Science, 13(2):259, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
259/784077.

N:1938:Ra

[N.38b] L. N. Reviews. ICES Journal of Marine Science, 13(2):246–247, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
246/784002.

N:1938:KS

[N.38c] W. N. Knud Schøning 1860–1938. ICES Journal of Marine Science, 13 (2):151–154, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/151/783907.

Nall:1932:SMS

[Nal32] G. Herbert Nall. The spawning mark on sea trout scales: a reply to Dr. Mottram’s criticisms. ICES Journal of Marine Science, 7(1):76– 80, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/76/947024.

Nesbit:1934:CMP

[Nes34] Robert A. Nesbit. A convenient method for preparing celluloid impressions of fish scales. ICES Journal of Marine Science, 9(3):373–376, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/373/615760.

Nesbit:1937:R

[Nes37] R. A. Nesbit. Reviews. ICES Journal of Marine Science, 12(3):359– 360, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/359/645438.

Nielsen:1933:QUM

[Nie33] E. Steemann Nielsen. Über quantitative Untersuchung von marinem Plankton mit Utermöhls umgekehrtem Mikroskop. (German) [On quantitative investigation of marine plankton with Utermöl’s inverted microscope]. ICES Journal of Marine Science, 8(2):201–210, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/201/618356.

Nielsen:1935:MEQ

[Nie35] E. Steemann Nielsen. Eine Methode zur exakten quantitativen Bestimmung von Zooplankton. Mit allgemeinen Bemerkungen über quantitative Planktonarbeiten. (German) [A method for the exact quantitative determination of zooplankton. With general remarks on quantitative plankton work]. ICES Journal of Marine Science, 10(3):302– 314, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/302/659218.

Nielsen:1937:RBQ

[Nie37a] E. Steemann Nielsen. On the relation between the quantities of phytoplankton and zooplankton in the sea. ICES Journal of Marine Science, 12(2):147–154, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/2/147/753753.

Nielsen:1937:R

[Nie37b] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 12 (2):211, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/211/753851.

Nielsen:1938:Ra

[Nie38a] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 13 (2):253, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/253/784039.

Nielsen:1938:Rb

[Nie38b] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 13 (2):254, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/254/784043.

Nielsen:1938:Rc

[Nie38c] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 13 (2):254, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/254/784053.

Nielsen:1938:ANQ

[Nie38d] E. Steemann Nielsen. Über die Anwendung von Netzfängen bei quantitativen Phytoplanktonuntersuchungen. (German) [On the use of nets in quantitative phytoplankton surveys]. ICES Journal of Marine Science, 13(2):197–205, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/2/197/783946.

Nielsen:1939:LE

[Nie39] E. Steemann Nielsen. Letter to the Editor. ICES Journal of Marine Science, 14(1):106–107, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/1/106/887970.

Nolte:1937:ZKP

[Nol37] W. Nolte. Die Zucht und Konservierung der Polychaetenlarven. (German) [The breeding and conservation of polychaete larvae]. ICES Journal of Marine Science, 12(1):51–76, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/1/51/626367.

Norman:1936:R

[Nor36] J. R. Norman. Reviews. ICES Journal of Marine Science, 11(3): 376, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/376/621033.

Norman:1939:Ra

[Nor39a] J. R. Norman. Reviews. ICES Journal of Marine Science, 14(1): 135–136, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/135/888459.

Norman:1939:Rb

[Nor39b] J. R. Norman. Reviews. ICES Journal of Marine Science, 14(3): 427, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/427/721802.

Norman:1939:Rc

[Nor39c] J. R. Norman. Reviews. ICES Journal of Marine Science, 14(3): 427–428, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/427/721804.

O:1930:R

[O.30] A. P. O. Reviews. ICES Journal of Marine Science, 5(3):401, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/401/648199.

O:1931:Rb

[O.31a] A. P. O. Reviews. ICES Journal of Marine Science, 6(2):332–333, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/332/600658.

O:1931:Rc

[O.31b] A. P. O. Reviews. ICES Journal of Marine Science, 6(3):476, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/476/724511.

O:1931:Rd

[O.31c] A. P. O. Reviews. ICES Journal of Marine Science, 6(3):480–481, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/480/724536.

O:1931:Ra

[O.31d] J. H. O. Reviews. ICES Journal of Marine Science, 6(1):131–132, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/131/641676.

O:1938:R

[O.38] F. D. O. Reviews. ICES Journal of Marine Science, 13(3):407– 409, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/407/635528.

O:1939:R

[O.39] F. D. O. Reviews. ICES Journal of Marine Science, 14(1):138–139, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/138/888469.

Ommanney:1935:R

[Omm35] F. D. Ommanney. Reviews. ICES Journal of Marine Science, 10(2): 219–220, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/219/672937.

Orton:1930:PCS

[Ort30] J. H. Orton. The production of clean shellfish. ICES Journal of Marine Science, 5(3):385–392, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/5/3/385/648168.

Orton:1934:R

[Ort34] J. H. Orton. Reviews. ICES Journal of Marine Science, 9(2):273–274, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/273/743410.

Orton:1938:R

[Ort38] J. H. Orton. Reviews. ICES Journal of Marine Science, 13(2): 263–264, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/263/784105.

Orton:1939:R

[Ort39] J. H. Orton. Reviews. ICES Journal of Marine Science, 14(1):143, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/143/888491.

Ostenfeld:1930:R

[Ost30] C. H. Ostenfeld. Reviews. ICES Journal of Marine Science, 5(3): 404–405, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/404/648209.

Otterstrom:1931:TAW

[Ott31] C. V. Otterstrøm. Die Turbinen und die abwärts wandernden jungen Lachse und Forellen. (German) [Turbines and young salmon and trout migrating downwards]. ICES Journal of Marine Science, 6(2): 273–286, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/273/600448.

Otterstrom:1932:TAW

[Ott32] C. V. Otterstrøm. Die Turbinen und die abwärts wandernden jungen Lachse und Forellen: 2. Teil. (German) [Turbines and young salmon and trout migrating downwards. Part 2]. ICES Journal of Marine Science, 7(1):63–75, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/7/1/63/947018.

Otterstrom:1933:RLS

[Ott33] C. V. Otterstrøm. Reife Lachse (Salmo salar) in Teichen. (German) [Ripe salmon (Salmo salar) in ponds]. ICES Journal of Marine Science, 8(1):83–89, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/8/1/83/682468.

Ottestad:1934:LE

[Ott34] Per Ottestad. Letter to the Editor. ICES Journal of Marine Science, 9 (2):249–253, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/249/743167.

Otterstrom:1936:TAW

[Ott36] C. V. Otterstrøm. Die Turbinen und die abwärts wandernden jungen Lachse und Forellen (sowie Aale) 3. Teil. (German) [Turbines and young salmon and trout (and eels) migrating downwards. Part 3]. ICES Journal of Marine Science, 11(1):72–93, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/11/1/72/640300.

Ottestad:1937:R

[Ott37] Per Ottestad. Reviews. ICES Journal of Marine Science, 12(3):350– 351, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/350/645408.

Ottestad:1938:R

[Ott38] Per Ottestad. Reviews. ICES Journal of Marine Science, 13(3):404– 405, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/404/635519.

P:1934:Ra

[P.34a] E. P. Reviews. ICES Journal of Marine Science, 9(1):106–107, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/106/944963.

P:1934:Rb

[P.34b] E. P. Reviews. ICES Journal of Marine Science, 9(1):107, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/9/1/
107/944966.

P:1934:Rc

[P.34c] E. P. Reviews. ICES Journal of Marine Science, 9(1):107–108, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/107/944969.

P:1935:Ra

[P.35a] J. P. Reviews. ICES Journal of Marine Science, 10(2):191–192, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
191/672850.

P:1935:Rb

[P.35b] J. P. Reviews. ICES Journal of Marine Science, 10(3):324–325, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/3/324/659235.

P:1936:R

[P.36] H. H. P. Reviews. ICES Journal of Marine Science, 11(3):371, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/3/
371/621014.

P:1937:Rb

[P.37a] F. T. K. P. Reviews. ICES Journal of Marine Science, 12(1):98, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
98/626459.

P:1937:Rc

[P.37b] F. T. K. P. Reviews. ICES Journal of Marine Science, 12(1):98, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
98/626464.

P:1937:Ra

[P.37c] J. P. Reviews. ICES Journal of Marine Science, 12(1):82–83, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
82/626390.

P:1938:Rd

[P.38a] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(1): 113–114, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/113/714493.

P:1938:Re

[P.38b] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(1):127, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
127/714582.

P:1938:Rf

[P.38c] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(2): 268–269, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/268/784126.

P:1938:Rg

[P.38d] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(3): 402, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/402/635512.

P:1938:Rh

[P.38e] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(3):402– 403, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/402/635515.

P:1938:Ri

[P.38f] F. T. K. P. Reviews. ICES Journal of Marine Science, 13(3):403– 404, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/403/635516.

P:1938:Ra

[P.38g] H. H. P. Reviews. ICES Journal of Marine Science, 13(1):111, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
111/714484.

P:1938:Rb

[P.38h] H. H. P. Reviews. ICES Journal of Marine Science, 13(1):111–113, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/111/714488.

P:1938:Rc

[P.38i] H. H. P. Reviews. ICES Journal of Marine Science, 13(1):113, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
113/714491.

P:1939:Ra

[P.39a] F. T. K. P. Reviews. ICES Journal of Marine Science, 14(2): 305–307, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/305/737353.

P:1939:Rb

[P.39b] F. T. K. P. Reviews. ICES Journal of Marine Science, 14(2):311, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
311/737364.

P:1939:Rd

[P.39c] F. T. K. P. Reviews. ICES Journal of Marine Science, 14(3):426– 427, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/426/721800.

P:1939:Re

[P.39d] F. T. K. P. Reviews. ICES Journal of Marine Science, 14(3):428– 429, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/428/721806.

P:1939:Rc

[P.39e] H. H. P. Reviews. ICES Journal of Marine Science, 14(3):415– 416, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/415/721731.

Palmen:1931:Ra

[Pal31a] E. Palmén. Reviews. ICES Journal of Marine Science, 6(2):299, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/299/600476.

Palmen:1931:Rb

[Pal31b] E. Palmén. Reviews. ICES Journal of Marine Science, 6(2):299–300, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/299/600479.

Palmen:1931:BTT

[Pal31c] E. Palmén. Zur Bestimmung des Triftstromes aus Terminbeobachtungen. (German) [Determination of the drift current from deadline observations]. ICES Journal of Marine Science, 6(3):387–401, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/387/724418.

Palmen:1932:R

[Pal32] E. Palmén. Reviews. ICES Journal of Marine Science, 7(2):291–292, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/291/708273.

Parr:1936:PRB

[Par36a] A. E. Parr. On the probable relationship between vertical stability and lateral mixing processes. ICES Journal of Marine Science, 11(3): 308–313, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/308/620988.

Parr:1936:RBD

[Par36b] A. E. Parr. On the relationship between dynamic topography and direction of current under the influence of external (climatic) factors. ICES Journal of Marine Science, 11(3):299–307, December 1936. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/11/3/299/620984.

Parr:1937:LE

[Par37] A. E. Parr. Letter to the Editor. ICES Journal of Marine Science, 12 (1):77–78, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/77/626373.

Parr:1939:AER

[Par39] A. E. Parr. Analysis of evidence relating to eddy diffusion processes at mid-depths in the Eastern Caribbean. ICES Journal of Marine Science, 14(3):347–356, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/3/347/721630.

Paulsen:1932:JN

[Pau32] Ove Paulsen. J. N. Nielsen. (1877–1932). ICES Journal of Marine Science, 7(3):343, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/3/343/694214.

Pentelow:1932:R

[Pen32] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 7(1): 111, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/111/946942.

Pentelow:1933:R

[Pen33] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 8 (1):119–120, April 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/1/119/682421.

Pentelow:1934:Ra

[Pen34a] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 9 (1):130–131, April 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/1/130/945056.

Pentelow:1934:Rb

[Pen34b] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 9 (2):286, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/286/743473.

Pentelow:1935:R

[Pen35] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 10(2): 205–206, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/205/672902.

Pentelow:1937:Ra

[Pen37a] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 12 (1):98–100, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/98/626469.

Pentelow:1937:Rb

[Pen37b] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 12(3): 367–368, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/367/645486.

Pentelow:1938:R

[Pen38] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 13 (1):126–127, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/126/714577.

Pettersson:1930:GE

[Pet30] Otto Pettersson. Gustaf Ekmax 1852–1930. ICES Journal of Marine Science, 5(3):287–290, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/3/287/648145.

Pettersson:1935:SDT

[Pet35] Hans Pettersson. Submarine daylight and the transparency of sea water. ICES Journal of Marine Science, 10(1):48–65, April 1935. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/10/1/48/615492.

Poulsen:1931:R

[Pou31] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 6(2): 308–309, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/308/600501.

Poulsen:1932:Ra

[Pou32a] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 7(3): 422–423, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/422/694293.

Poulsen:1932:Rb

[Pou32b] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 7(3): 424–425, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/424/694295.

Poulsen:1938:R

[Pou38] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 13(1): 121–122, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/121/714554.

Pratje:1934:R

[Pra34] Pratje. Reviews. ICES Journal of Marine Science, 9(2):271–273, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/271/743391.

Proudman:1931:R

[Pro31] J. Proudman. Reviews. ICES Journal of Marine Science, 6(2): 300–301, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/300/600484.

Proudman:1932:R

[Pro32] J. Proudman. Reviews. ICES Journal of Marine Science, 7(1):87, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/87/947059.

P-T:1933:R

[PT33] T. E. P-T. Reviews. ICES Journal of Marine Science, 8(3):412–414, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/412/730431.

P-T:1936:R

[PT36] T. E. P.-T. Reviews. ICES Journal of Marine Science, 11(3):381– 382, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/381/621057.

P-T:1938:R

[PT38] T. E. P.-T. Reviews. ICES Journal of Marine Science, 13(2): 265–266, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/265/784116.

P-T:1939:R

[PT39] T. E. P.-T. Reviews. ICES Journal of Marine Science, 14(2): 309–311, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/309/737362.

Phifer:1935:SMO

[PU35] Lyman D. Phifer and C. L. Utterback. Some meteorological observations. ICES Journal of Marine Science, 10(3):315–317, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
315/659220.

R:1930:R

[R.30] H. C. R. Reviews. ICES Journal of Marine Science, 5(2):245–247, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/245/627854.

R:1931:Ra

[R.31a] F. S. R. Reviews. ICES Journal of Marine Science, 6(1):120, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/6/1/
120/641625.

R:1931:Rb

[R.31b] F. S. R. Reviews. ICES Journal of Marine Science, 6(3):481–482, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/481/724539.

R:1931:Rc

[R.31c] F. S. R. Reviews. ICES Journal of Marine Science, 6(3):484–486, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/484/724552.

R:1932:Ra

[R.32a] J. A. R. Reviews. ICES Journal of Marine Science, 7(1):94–95, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/94/947125.

R:1932:Rb

[R.32b] J. A. R. Reviews. ICES Journal of Marine Science, 7(1):95–96, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/95/947127.

R:1932:Rc

[R.32c] J. A. R. Reviews. ICES Journal of Marine Science, 7(1):96–97, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/96/947130.

R:1932:Rd

[R.32d] J. A. R. Reviews. ICES Journal of Marine Science, 7(2):309–310, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/309/708356.

R:1933:JS

[R.33a] C. T. R. Johannes Schmidt (1877–1933). ICES Journal of Marine Science, 8(2):145–152, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/2/145/618338.

R:1933:Ra

[R.33b] F. S. R. Reviews. ICES Journal of Marine Science, 8(1):111, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/8/1/
111/682397.

R:1933:Rb

[R.33c] J. A. R. Reviews. ICES Journal of Marine Science, 8(1):124–125, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/124/682441.

R:1934:Rb

[R.34a] E. S. R. Reviews. ICES Journal of Marine Science, 9(1):132, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/9/1/
132/945063.

R:1934:Ra

[R.34b] J. R. Reviews. ICES Journal of Marine Science, 9(1):131–132, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/131/945060.

R:1935:Ra

[R.35a] D. S. R. Reviews. ICES Journal of Marine Science, 10(1):118–119, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/118/615449.

R:1935:Rb

[R.35b] F. S. R. Reviews. ICES Journal of Marine Science, 10(3):324, December 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/3/
324/659228.

R:1935:Rc

[R.35c] H. C. R. Reviews. ICES Journal of Marine Science, 10(3):337– 338, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/337/659278.

R:1936:Ra

[R.36a] A. R. Reviews. ICES Journal of Marine Science, 11(1):109–110, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
109/640245.

R:1936:Rb

[R.36b] D. S. R. Reviews. ICES Journal of Marine Science, 11(1):110, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
110/640250.

R:1936:Rc

[R.36c] D. S. R. Reviews. ICES Journal of Marine Science, 11(1):110–111, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/110/640252.

R:1936:Rd

[R.36d] F. S. R. Reviews. ICES Journal of Marine Science, 11(2):266–269, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/2/266/663244.

R:1936:Rf

[R.36e] F. S. R. Reviews. ICES Journal of Marine Science, 11(3):372– 373, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/372/621020.

R:1936:Re

[R.36f] J. A. R. Reviews. ICES Journal of Marine Science, 11(2):274–275, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/2/274/663250.

R:1937:Ra

[R.37a] D. S. R. Reviews. ICES Journal of Marine Science, 12(1):93–94, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
93/626443.

R:1937:Rf

[R.37b] D. S. R. Reviews. ICES Journal of Marine Science, 12(3):358– 359, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/358/645435.

R:1937:Rb

[R.37c] E. S. R. Reviews. ICES Journal of Marine Science, 12(1):102, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
102/626326.

R:1937:Rc

[R.37d] H. C. R. Reviews. ICES Journal of Marine Science, 12(1):103–104, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/1/103/626332.

R:1937:Rd

[R.37e] H. C. R. Reviews. ICES Journal of Marine Science, 12(1):104, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
104/626337.

R:1937:Re

[R.37f] H. C. R. Reviews. ICES Journal of Marine Science, 12(1):105, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
105/626340.

R:1937:Rg

[R.37g] H. C. R. Reviews. ICES Journal of Marine Science, 12(3):360– 361, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/360/645448.

R:1937:Rh

[R.37h] H. C. R. Reviews. ICES Journal of Marine Science, 12(3):362, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/3/
362/645453.

R:1937:Ri

[R.37i] H. C. R. Reviews. ICES Journal of Marine Science, 12(3):362– 363, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/362/645458.

R:1938:Ra

[R.38a] H. R. Reviews. ICES Journal of Marine Science, 13(2):267–268, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
267/784121.

R:1938:Rb

[R.38b] H. C. R. Reviews. ICES Journal of Marine Science, 13(2):269–272, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/269/784133.

R:1938:Rc

[R.38c] H. C. R. Reviews. ICES Journal of Marine Science, 13(3):389– 390, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/389/635469.

R:1939:Ra

[R39a] F. S. R. Reviews. ICES Journal of Marine Science, 14(1):122, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
122/888372.

R:1939:Rb

[R.39b] H. C. R. Reviews. ICES Journal of Marine Science, 14(2):312, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
312/737366.

R:1939:Rc

[R.39c] H. C. R. Reviews. ICES Journal of Marine Science, 14(2):312–313, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/2/312/737370.

R:1939:Rd

[R.39d] H. C. R. Reviews. ICES Journal of Marine Science, 14(3):421– 422, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/421/721778.

R:1939:Re

[R.39e] H. C. R. Reviews. ICES Journal of Marine Science, 14(3):422, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/3/
422/721782.

Rasmussen:1939:Ra

[RA39a] H. Baggesgaard Rasmussen and V. Sten Andersen. Reviews. ICES Journal of Marine Science, 14(2):280–282, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/2/280/737150.

Rasmussen:1939:Rb

[RA39b] H. Baggesgaard Rasmussen and V. Sten Andersen. Reviews. ICES Journal of Marine Science, 14(2):282–285, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/2/282/737153.

Rasmussen:1939:Rc

[RA39c] H. Baggesgaard Rasmussen and V. Sten Andersen. Reviews. ICES Journal of Marine Science, 14(2):285–287, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/14/2/285/737157.

Rae:1939:MEL

[Rae39] Bennet B. Rae. Marking experiments on lemon soles at Faroe 1923– 1936, with a note on Icelandic markings 1925. ICES Journal of Marine Science, 14(1):35–47, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/1/35/888505.

Raitt:1934:PAS

[Rai34] D. S. Raitt. A preliminary account of the sandeels of Scottish waters. ICES Journal of Marine Science, 9(3):365–372, December 1934. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/9/3/365/615759.

Raitt:1935:UAS

[Rai35] D. S. Raitt. On the use of alizarin as a stain for the skeletal structures of fishes. ICES Journal of Marine Science, 10(1):75–80, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
75/615496.

Raitt:1936:SRF

[Rai36] D. S. Raitt. Stock replenishment and fishing intensity in the haddock of the North Sea. ICES Journal of Marine Science, 11(2): 211–218, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/211/663212.

Ramalho:1930:R

[Ram30] A. Ramalho. Reviews. ICES Journal of Marine Science, 5(2):262–264, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/262/627880.

Ramalho:1937:Ra

[Ram37a] A. Ramalho. Reviews. ICES Journal of Marine Science, 12(2): 205–206, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/205/753833.

Ramalho:1937:Rb

[Ram37b] A. Ramalho. Reviews. ICES Journal of Marine Science, 12(2): 222–223, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/222/753891.

Ramalho:1937:Rc

[Ram37c] A. Ramalho. Reviews. ICES Journal of Marine Science, 12(2):236, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/236/753973.

Russell:1932:SBF

[RB32] E. S. Russell and H. O. Bull. A selected bibliography of fish behaviour. ICES Journal of Marine Science, 7(2):255–276, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/7/2/255/708208.

Rounsefell:1933:TEP

[RD33] George A. Rounsefell and Edwin H. Dahlgren. Tagging experiments on the Pacific herring, Clupea pallasii. ICES Journal of Marine Science, 8(3):371–384, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/8/3/371/730312.

Redeke:1934:OTP

[Red34] H. C. Redeke. On the occurrence of two pelagic copepods, Acartia bifilosa and Acartia tonsa, in the brackish waters of the Netherlands. ICES Journal of Marine Science, 9(1):39–45, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/9/1/39/945116.

Redeke:1935:SRP

[Red35] H. C. Redeke. Some recent publications on the fauna of brackish waters in the Netherlands. ICES Journal of Marine Science, 10(3):

319–323, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/319/659222.

Redeke:1939:ECZ

[Red39] H. C. Redeke. The effect of the closure of the zuiderzee on fish and fisheries. ICES Journal of Marine Science, 14(3):337–346, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/3/
337/721627.

Reibisch:1931:KBG

[Rei31] J. Reibisch. Karl Brandt, gestorben am 7. januar 1931. ICES Journal of Marine Science, 6(2):157–159, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/6/2/157/600417.

Riel:1932:R

[Rie32] V. Riel. Reviews. ICES Journal of Marine Science, 7(2):292–293, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/292/708278.

Ritchie:1936:R

[Rit36] James Ritchie. Reviews. ICES Journal of Marine Science, 11(3): 384–385, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/384/621073.

Robertson:1933:NOA

[Rob33] J. Armitage Robertson. Notes on the optical appearance of winter rings in some clupeoid scales. ICES Journal of Marine Science, 8(3): 357–370, December 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/3/357/730309.

Robertson:1936:OLP

[Rob36a] J. Armitage Robertson. The occurrence of Lee’s phenomenon in the sprat, and the size-relation between fish and scale. ICES Journal of Marine Science, 11(2):219–228, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/11/2/219/663215.

Robertson:1936:PAS

[Rob36b] J. Armitage Robertson. A preliminary account of the sprat and sprat fishery in England. ICES Journal of Marine Science, 11(1): 60–71, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/60/640294.

Robertson:1939:R

[Rob39] J. A. Robertson. Reviews. ICES Journal of Marine Science, 14(2): 304–305, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/304/737351.

Rollefsen:1932:SCE

[Rol32] Gunnar Rollefsen. The susceptibility of cod eggs to external influences. ICES Journal of Marine Science, 7(3):367–373, December 1932. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/7/3/367/694223.

Rosenberg:1938:R

[Ros38] M. Rosenberg. Reviews. ICES Journal of Marine Science, 13(3): 400–402, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/400/635509.

Roule:1934:MF

[Rou34] Louis Roule. Migration of fishes. ICES Journal of Marine Science, 9(1):96–97, April 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/1/96/945137.

Ross:1939:IFC

[RU39] F. W. Ross and C. L. Utterback. Intensity fluctuations in components of solar radiation with atmospheric conditions. ICES Journal of Marine Science, 14(2):193–200, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/14/2/193/737082.

Runnstrom:1933:RNF

[Run33] Sven Runnstrøm. Über die Rassenverhältnisse bei dem norwegischen Frühjahrshering mit besonderer Berücksichtigung der Konstanz der Rassenmerkmale. (German) [On the racial relations of the Norwegian spring herring with special consideration for the constancy of racial characteristics]. ICES Journal of Marine Science, 8(2):235–249, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/235/618366.

Runnstrom:1937:RNH

[Run37a] Sven Runnstrøm. A review of the Norwegian herring investigations in recent years. ICES Journal of Marine Science, 12(2):123–143, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
123/753740.

Runnstrom:1937:Ra

[Run37b] Sven Runnstrøm. Reviews. ICES Journal of Marine Science, 12(2): 229–231, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/229/753928.

Runnstrom:1937:Rb

[Run37c] Sven Runnstrøm. Reviews. ICES Journal of Marine Science, 12(2): 231–232, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/231/753930.

Runnstrom:1938:R

[Run38] Sven Runnström. Reviews. ICES Journal of Marine Science, 13(3): 400, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/400/635507.

Runnstrom:1939:R

[Run39] Sven Runnström. Reviews. ICES Journal of Marine Science, 14 (1):131–132, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/131/888445.

Russell:1931:STC

[Rus31] E. S. Russell. Some theoretical considerations on the “overfishing” problem. ICES Journal of Marine Science, 6(1):3–20, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/6/1/3/641714.

Russell:1934:PS

[Rus34] F. S. Russell. Plymouth sagitta. ICES Journal of Marine Science, 9(1):93, April 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/1/93/945131.

Russell:1939:R

[Rus39a] E. S. Russell. Reviews. ICES Journal of Marine Science, 14(3):423– 424, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/423/721792.

Russell:1939:HBC

[Rus39b] F. S. Russell. Hydrographical and biological conditions in the North Sea as indicated by plankton organisms. ICES Journal of Marine Science, 14(2):171–192, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/2/171/737070.

Ruud:1930:NPS

[Ruu30] Johan T. Ruud. Nitrates and phosphates in the Southern Seas. ICES Journal of Marine Science, 5(3):347–360, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/3/347/648151.

Ruud:1931:Ra

[Ruu31a] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 6(1): 136–140, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/136/641688.

Ruud:1931:Rb

[Ruu31b] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 6(2): 328–330, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/328/600644.

Ruud:1933:R

[Ruu33] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 8(1):102– 104, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/102/682379.

Ruud:1935:Ra

[Ruu35a] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 10(1): 108–110, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/108/615425.

Ruud:1935:Rb

[Ruu35b] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 10(2): 220–222, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/220/672939.

Ruud:1938:R

[Ruu38] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 13(3): 405–407, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/405/635523.

Ruud:1939:R

[Ruu39] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 14(1): 136–138, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/136/888464.

Robinson:1934:FAA

[RW34a] Rex J. Robinson and Henry E. Wirth. Free ammonia, albuminoid nitrogen, and organic nitrogen in the waters of the Pacific Ocean off the coasts of Washington and Vancouver Island. ICES Journal of Marine Science, 9(2):187–195, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/2/187/743137.

Robinson:1934:RFA

[RW34b] Rex J. Robinson and Henry E. Wirth. Report on the free ammonia, albuminoid nitrogen and organic nitrogen in the waters of the Puget Sound Area, during the summers of 1931 and 1932. ICES Journal of Marine Science, 9(1):15–27, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/1/15/945092.

S:1930:Ri

[S.30a] E. J. S. Reviews. ICES Journal of Marine Science, 5(3):430, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/430/648273.

S:1930:Rb

[S.30b] H. P. S. Reviews. ICES Journal of Marine Science, 5(1):123–124, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/123/676343.

S:1930:Rg

[S.30c] H. P. S. Reviews. ICES Journal of Marine Science, 5(2):267–270, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/267/627896.

S:1930:Rc

[S.30d] M. S. Reviews. ICES Journal of Marine Science, 5(2):242, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/242/627842.

S:1930:Rd

[S.30e] M. S. Reviews. ICES Journal of Marine Science, 5(2):245, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/245/627852.

S:1930:Re

[S.30f] M. S. Reviews. ICES Journal of Marine Science, 5(2):252, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/252/627862.

S:1930:Ra

[S.30g] O. S. Reviews. ICES Journal of Marine Science, 5(1):121–122, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/121/676331.

S:1930:Rh

[S.30h] O. S. Reviews. ICES Journal of Marine Science, 5(3):417–419, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/417/648243.

S:1930:Rf

[S.30i] R. E. S. Reviews. ICES Journal of Marine Science, 5(2):252–255, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/252/627866.

S:1931:Rb

[S.31a] H. P. S. Reviews. ICES Journal of Marine Science, 6(2):323–325, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/323/600626.

S:1931:Rc

[S.31b] H. P. S. Reviews. ICES Journal of Marine Science, 6(2):325–326, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/325/600632.

S:1931:Rf

[S.31c] H. P. S. Reviews. ICES Journal of Marine Science, 6(3):504–505, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/504/724616.

S:1931:Rd

[S.31d] H. U. S. Reviews. ICES Journal of Marine Science, 6(3):466–467, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/466/724496.

S:1931:Re

[S.31e] H. U. S. Reviews. ICES Journal of Marine Science, 6(3):467–468, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/467/724500.

S:1931:Ra

[S.31f] R. S. Reviews. ICES Journal of Marine Science, 6(1):134–136, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/134/641684.

S:1932:Rh

[S.32a] G. A. S. Reviews. ICES Journal of Marine Science, 7(2):300–301, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/300/708318.

S:1932:Ri

[S.32b] G. A. S. Reviews. ICES Journal of Marine Science, 7(2):315–317, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/315/708382.

S:1932:Ra

[S.32c] G. M. S. Reviews. ICES Journal of Marine Science, 7(1):91, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/7/1/
91/947089.

S:1932:Re

[S.32d] H. P. S. Reviews. ICES Journal of Marine Science, 7(1):114–115, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/114/946967.

S:1932:Rk

[S.32e] H. P. S. Reviews. ICES Journal of Marine Science, 7(2):318–319, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/318/708394.

S:1932:Rl

[S.32f] H. P. S. Reviews. ICES Journal of Marine Science, 7(2):319–320, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/319/708402.

S:1932:Rn

[S.32g] H. P. S. Reviews. ICES Journal of Marine Science, 7(3):425–426, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/425/694303.

S:1932:Rf

[S.32h] H. U. S. Reviews. ICES Journal of Marine Science, 7(2):288–289, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/288/708246.

S:1932:Rg

[S.32i] H. U. S. Reviews. ICES Journal of Marine Science, 7(2):289, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/289/708260.

S:1932:Rm

[S.32j] H. U. S. Reviews. ICES Journal of Marine Science, 7(3):415, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/415/694265.

S:1932:Rb

[S.32k] O. S. Reviews. ICES Journal of Marine Science, 7(1):103–104, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/103/946901.

S:1932:Rc

[S.32l] O. S. Reviews. ICES Journal of Marine Science, 7(1):104–105, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/104/946907.

S:1932:Rd

[S.32m] R. S. Reviews. ICES Journal of Marine Science, 7(1):106–108, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/106/946926.

S:1932:Rj

[S.32n] R. S. Reviews. ICES Journal of Marine Science, 7(2):317–318, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/317/708389.

S:1933:Rc

[S.33a] G. A. S. Reviews. ICES Journal of Marine Science, 8(2):276–277, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/276/618428.

S:1933:Rb

[S.33b] H. U. S. Reviews. ICES Journal of Marine Science, 8(2):262–263, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/262/618392.

S:1933:Rg

[S.33c] H. U. S. Reviews. ICES Journal of Marine Science, 8(3):401, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/401/730349.

S:1933:Rd

[S.33d] O. S. Reviews. ICES Journal of Marine Science, 8(2):278–279, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/278/618432.

S:1933:Re

[S.33e] O. S. Reviews. ICES Journal of Marine Science, 8(2):279–280, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/279/618437.

S:1933:Rf

[S.33f] O. S. Reviews. ICES Journal of Marine Science, 8(2):281–282, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/281/618439.

S:1933:Ra

[S.33g] R. E. S. Reviews. ICES Journal of Marine Science, 8(1):121, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/8/1/
121/682432.

S:1934:Rd

[S.34a] G. M. S. Reviews. ICES Journal of Marine Science, 9(3):429–431, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/429/615831.

S:1934:Rc

[S.34b] H. P. S. Reviews. ICES Journal of Marine Science, 9(2):286–287, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/286/743485.

S:1934:Ra

[S.34c] H. U. S. Reviews. ICES Journal of Marine Science, 9(2):261, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/261/743219.

S:1934:Rb

[S.34d] R. S. Reviews. ICES Journal of Marine Science, 9(2):283–286, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/283/743454.

S:1935:Rd

[S.35a] A. C. S. Reviews. ICES Journal of Marine Science, 10(2):206–207, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/206/672905.

S:1935:Re

[S.35b] A. C. S. Reviews. ICES Journal of Marine Science, 10(2):215–216, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/215/672926.

S:1935:Ra

[S.35c] G. A. S. Reviews. ICES Journal of Marine Science, 10(1):120–121, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/120/615454.

S:1935:Rb

[S.35d] O. S. Reviews. ICES Journal of Marine Science, 10(1):121–122, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
121/615456.

S:1935:Rc

[S.35e] O. S. Reviews. ICES Journal of Marine Science, 10(1):122–123, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
122/615459.

S:1936:Rc

[S.36a] G. A. S. Reviews. ICES Journal of Marine Science, 11(1):116, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
116/640267.

S:1936:Ra

[S.36b] H. U. S. Reviews. ICES Journal of Marine Science, 11(1):98, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
98/640304.

S:1936:Rb

[S.36c] H. U. S. Reviews. ICES Journal of Marine Science, 11(1):98–99, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/1/
98/640305.

S:1936:Rd

[S.36d] H. U. S. Reviews. ICES Journal of Marine Science, 11(2):260, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
260/663230.

S:1936:Re

[S.36e] O. S. Reviews. ICES Journal of Marine Science, 11(2):265–266, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
265/663240.

S:1937:Rd

[S.37a] A. C. S. Reviews. ICES Journal of Marine Science, 12(2):227–228, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/227/753921.

S:1937:Rh

[S.37b] A. C. S. Reviews. ICES Journal of Marine Science, 12(3):363, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/3/
363/645461.

S:1937:Ri

[S.37c] A. C. S. Reviews. ICES Journal of Marine Science, 12(3):364, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/3/
364/645463.

S:1937:Re

[S.37d] H. P. S. Reviews. ICES Journal of Marine Science, 12(2):233, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
233/753960.

S:1937:Ra

[S.37e] H. U. S. Reviews. ICES Journal of Marine Science, 12(1):79, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
79/626378.

S:1937:Rf

[S.37f] H. U. S. Reviews. ICES Journal of Marine Science, 12(3):339– 340, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/339/645352.

S:1937:Rb

[S.37g] O. S. Reviews. ICES Journal of Marine Science, 12(2):219, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
219/753881.

S:1937:Rc

[S.37h] O. S. Reviews. ICES Journal of Marine Science, 12(2):220, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
220/753884.

S:1937:KA

[S.37i] O. S. S. K. Asserson. (1882–1937). ICES Journal of Marine Science, 12(3):259–262, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/3/259/645323.

S:1937:Rg

[S.37j] S. E. S. Reviews. ICES Journal of Marine Science, 12(3):340– 341, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/340/645357.

S:1938:Rd

[S.38a] G. A. S. Reviews. ICES Journal of Marine Science, 13(2):260, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
260/784084.

S:1938:Re

[S.38b] H. P. S. Reviews. ICES Journal of Marine Science, 13(2):264, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
264/784110.

S:1938:Rc

[S.38c] H. U. S. Reviews. ICES Journal of Marine Science, 13(2):244–245, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/244/783995.

S:1938:Rf

[S.38d] H. U. S. Reviews. ICES Journal of Marine Science, 13(3):379– 380, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/379/635453.

S:1938:Rb

[S.38e] R. E. S. Reviews. ICES Journal of Marine Science, 13(1):122–123, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/122/714557.

S:1938:Ra

[S.38f] S. E. S. Reviews. ICES Journal of Marine Science, 13(1):105–106, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/105/714451.

S:1939:Rb

[S.39a] H. P. S. Reviews. ICES Journal of Marine Science, 14(1):142–143, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/142/888486.

S:1939:Ra

[S.39b] O. S. Reviews. ICES Journal of Marine Science, 14(1):123–124, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
123/888384.

Schnakenbeck:1930:R

[Sch30a] W. Schnakenbeck. Reviews. ICES Journal of Marine Science, 5(3): 407–408, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/407/648213.

Schulz:1930:HB

[Sch30b] Bruno Schulz. Die hydrographie des barentsmeeres. ICES Journal of Marine Science, 5(3):291–319, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/3/291/648146.

Schulz:1930:Ra

[Sch30c] Bruno Schulz. Reviews. ICES Journal of Marine Science, 5(2): 235–236, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/235/627825.

Schulz:1930:Rb

[Sch30d] Bruno Schulz. Reviews. ICES Journal of Marine Science, 5(3):392– 393, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/392/648174.

Schnakenbeck:1931:Ra

[Sch31a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 6(1): 127–128, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/127/641658.

Schnakenbeck:1931:Rb

[Sch31b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 6(1):129, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/129/641663.

Schnakenbeck:1931:RFG

[Sch31c] W. Schnakenbeck. Zum Rassenproblem bei den Fischen. (German) [On the race problem of fish]. ICES Journal of Marine Science, 6 (1):28–40, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/28/641712.

Schott:1931:R

[Sch31d] G. Schott. Reviews. ICES Journal of Marine Science, 6(3):468–471, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/468/724503.

Schnakenbeck:1932:R

[Sch32a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 7(2): 306–308, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/306/708344.

Schubert:1932:Rb

[Sch32b] O. V. Schubert. Reviews. ICES Journal of Marine Science, 7(3): 414–415, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/414/694261.

Schubert:1932:Ra

[Sch32c] V. Schubert. Reviews. ICES Journal of Marine Science, 7(2):288, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
2/288/708236.

Schulz:1932:Ra

[Sch32d] Bruno Schulz. Reviews. ICES Journal of Marine Science, 7(3):415– 417, December 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/3/415/694271.

Schulz:1932:Rb

[Sch32e] Bruno Schulz. Reviews. ICES Journal of Marine Science, 7(3):417, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
3/417/694275.

Schubert:1933:R

[Sch33a] O. V. Schubert. Reviews. ICES Journal of Marine Science, 8(1):104– 105, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/104/682380.

Schumacher:1933:R

[Sch33b] A. Schumacher. Reviews. ICES Journal of Marine Science, 8(3): 397–400, December 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/3/397/730336.

Schnakenbeck:1934:Ra

[Sch34a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 9(2): 274–277, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/274/743417.

Schnakenbeck:1934:Rb

[Sch34b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 9(3):426– 429, December 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/3/426/615829.

Schulz:1934:R

[Sch34c] Bruno Schulz. Reviews. ICES Journal of Marine Science, 9(1):101, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/101/944948.

Schumacher:1934:R

[Sch34d] A. Schumacher. Reviews. ICES Journal of Marine Science, 9(2): 257–259, August 1934. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/9/2/257/743191.

Schnakenbeck:1935:Ra

[Sch35a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 10(3): 339–340, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/339/659281.

Schnakenbeck:1935:Rb

[Sch35b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 10(3): 340–341, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/340/659284.

Schott:1935:R

[Sch35c] G. Schott. Reviews. ICES Journal of Marine Science, 10(3):327– 328, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/327/659252.

Schnakenbeck:1936:Ra

[Sch36a] Schnakenbeck. Reviews. ICES Journal of Marine Science, 11(1): 106–108, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/106/640235.

Schnakenbeck:1936:Rb

[Sch36b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 11(3): 379–381, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/379/621048.

Schulz:1936:BR

[Sch36c] Bruno Schulz. Book reviews. ICES Journal of Marine Science, 11(3): 366–367, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/366/621003.

Schumacher:1936:R

[Sch36d] A. Schumacher. Reviews. ICES Journal of Marine Science, 11(1): 99–100, April 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/1/99/640306.

Scherhag:1937:EAG

[Sch37a] R. Scherhag. Die Erwärmung der Arktis. (German) [Warming of the Arctic]. ICES Journal of Marine Science, 12(3):263–276, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/3/
263/645324.

Schnakenbeck:1937:R

[Sch37b] Schnakenbeck. Reviews. ICES Journal of Marine Science, 12(3): 354–356, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/354/645422.

Sch:1938:R

[Sch38a] G. Sch. Reviews. ICES Journal of Marine Science, 13(2):239–240, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
239/783977.

Schach:1938:AEM

[Sch38b] Helmut Schach. Ein Apparat zur Entnahme von Meerwasserproben aus der Tiefe für bakteriologische Untersuchungen. (German) [An apparatus for taking seawater samples from the depths for bacteriological investigations]. ICES Journal of Marine Science, 13(3):349– 356, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/349/635440.

Schulz:1938:R

[Sch38c] Bruno Schulz. Reviews. ICES Journal of Marine Science, 13(2): 237–239, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/237/783971.

Schach:1939:R

[Sch39a] Helmut Schach. Reviews. ICES Journal of Marine Science, 14(2): 287–288, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/287/737159.

Schott:1939:R

[Sch39b] G. Schott. Reviews. ICES Journal of Marine Science, 14(3):401– 403, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/401/721645.

Seiwell:1931:OPC

[Sei31] H. R. Seiwell. Observations on the phosphate-content and hydrogenion concentration of the North Sea, the southern entrance to the Norwegian Sea, and the water south of Iceland. ICES Journal of Marine Science, 6(2):213–231, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/2/213/600428.

Seiwell:1935:AOP

[Sei35] H. R. Seiwell. The annual organic production and nutrient phosphorus requirement in the Tropical Western North Atlantic. ICES Journal of Marine Science, 10(1):20–32, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/10/1/20/615478.

Seiwell:1937:OPL

[Sei37] H. R. Seiwell. The oxygen-poor layer in the Western North Atlantic. ICES Journal of Marine Science, 12(3):277–283, December 1937. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/12/3/277/645326.

Seiwell:1938:EDM

[Sei38] H. R. Seiwell. Eddy diffusion at mid-depths in the Caribbean Sea region. ICES Journal of Marine Science, 13(2):155–162, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
155/783916.

Seiwell:1939:DTV

[Sei39] H. R. Seiwell. Daily temperature variations in the Western North Atlantic. ICES Journal of Marine Science, 14(3):357–369, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/3/
357/721634.

Savage:1934:LIE

[SH34] R. E. Savage and W. C. Hodgson. Lunar influence on the East Anglian herring fishery. ICES Journal of Marine Science, 9(2):223–239, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/223/743151.

Sherwood:1931:OIN

[She31] B. A. H. P. Sherwood. The oyster industry in North America: a record of a brief tour of some of the centres on the Atlantic and Pacific coasts, and of a summer in Canada. ICES Journal of Marine Science, 6(3): 361–386, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/361/724400.

Shelford:1933:Ra

[She33a] V. E. Shelford. Reviews. ICES Journal of Marine Science, 8(1):108, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/108/682388.

Shelford:1933:Rb

[She33b] V. E. Shelford. Reviews. ICES Journal of Marine Science, 8(1):109, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/109/682390.

Shelford:1933:Rc

[She33c] V. E. Shelford. Reviews. ICES Journal of Marine Science, 8(1):109, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/109/682392.

Siedlecki:1930:R

[Sie30] M. Siedlecki. Reviews. ICES Journal of Marine Science, 5(2):264–266, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/264/627883.

Sklower:1939:UIO

[Skl39] Alfred Sklower. Untersuchungen über die inkretorischen Organe der Fische: II. Die Metamorphose der Plattfische und die Bedeutung der Schilddrüse für diesen Vorgang. (German) [Investigations into the incretory organs of fish: II. The metamorphosis of flatfish and the importance of the thyroid for this process]. ICES Journal of Marine Science, 14(1):81–85, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/81/888518.

Savage:1935:NAL

[SML35] R. E. Savage, A. More, and J. R. Lumby. A note on amoebae living in a sea water sample. ICES Journal of Marine Science, 10(2): 185–186, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/185/672839.

Southern:1930:R

[Sou30] R. Southern. Reviews. ICES Journal of Marine Science, 5(3):426–428, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/426/648263.

Sparck:1934:R

[Spä34] R. Spärck. Reviews. ICES Journal of Marine Science, 9(3):424–425, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/424/615825.

Sparck:1935:IQI

[Spä35] R. Spärck. On the importance of quantitative investigation of the bottom fauna in marine biology. ICES Journal of Marine Science, 10(1):3–19, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/3/615481.

Sparck:1936:R

[Spä36] R. Spärck. Reviews. ICES Journal of Marine Science, 11(3):381, December 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/3/381/621051.

Speerschneider:1938:R

[Spe38] C. I. Speerschneider. Reviews. ICES Journal of Marine Science, 13(3): 388–389, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/388/635468.

Soot-Ryen:1932:CNC

[SR32] T. Soot-Ryen. Changes in the nitrite content through the year in North-Norwegian fjords. ICES Journal of Marine Science, 7(2): 246–250, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/246/708203.

Schmidt-Ries:1939:POV

[SR39] H. Schmidt-Ries. Das Plankton der Ostsee vor der ostpreussischen Küste. (German) [Plankton of the Baltic Sea off the East Prussian coast]. ICES Journal of Marine Science, 14(1):7–24, April 1939. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/14/1/7/888514.

Stamp:1936:GEN

[Sta36] L. Dudley Stamp. The geographical evolution of the North Sea Basin. ICES Journal of Marine Science, 11(2):135–163, August 1936. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/11/2/135/663183.

Stenij:1930:R

[Ste30] S. E. Stenij. Reviews. ICES Journal of Marine Science, 5(2):240–242, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/240/627840.

Stephensen:1931:R

[Ste31] K. Stephensen. Reviews. ICES Journal of Marine Science, 6(1): 121–122, March 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/1/121/641630.

Steuer:1937:R

[Ste37] Ad. Steuer. Reviews. ICES Journal of Marine Science, 12(3):348– 349, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/348/645402.

Storrow:1939:ER

[Sto39] B. Storrow. Essay review. ICES Journal of Marine Science, 14(1): 108–113, April 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/1/108/887972.

Sund:1931:CSD

[Sun31a] Oscar Sund. Colorimetry at sea, with a description of a new colorimeter. ICES Journal of Marine Science, 6(2):241–245, August 1931. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/6/2/241/600435.

Sund:1931:NSS

[Sun31b] Oscar Sund. A new speed surface sampler. ICES Journal of Marine Science, 6(3):419–420, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/3/419/724457.

Sund:1931:R

[Sun31c] Oscar Sund. Reviews. ICES Journal of Marine Science, 6(1):111–112, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/111/641605.

Sund:1935:R

[Sun35] Oscar Sund. Reviews. ICES Journal of Marine Science, 10(1): 123–125, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/123/615463.

Sund:1936:R

[Sun36] Oscar Sund. Reviews. ICES Journal of Marine Science, 11(1):108, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/1/108/640240.

Sund:1939:R

[Sun39] O. Sund. Reviews. ICES Journal of Marine Science, 14(1):114, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
114/887974.

Sverdrup:1931:R

[Sve31] H. U. Sverdrup. Reviews. ICES Journal of Marine Science, 6(2): 289–290, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/289/600456.

Sven:1938:R

[Sve38a] Runnström Sven. Reviews. ICES Journal of Marine Science, 13(3): 398–399, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/398/635503.

Sverdrup:1938:EOM

[Sve38b] H. U. Sverdrup. On the explanation of the oxygen minima and maxima in the oceans. ICES Journal of Marine Science, 13(2):163–172, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
163/783919.

T:1930:Ra

[T.30a] H. T. Reviews. ICES Journal of Marine Science, 5(1):113–115, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/113/676311.

T:1930:Rb

[T.30b] H. T. Reviews. ICES Journal of Marine Science, 5(1):115–117, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/115/676318.

T:1930:Rc

[T.30c] H. T. Reviews. ICES Journal of Marine Science, 5(1):117–121, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/117/676325.

T:1930:Rf

[T.30d] H. T. Reviews. ICES Journal of Marine Science, 5(3):412–414, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/412/648227.

T:1930:Rd

[T.30e] J. B. T. Reviews. ICES Journal of Marine Science, 5(2):233–235, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/233/627823.

T:1930:Re

[T.30f] J. J. T. Reviews. ICES Journal of Marine Science, 5(2):249–252, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/249/627860.

T:1931:Rb

[T.31a] A. K. T. Reviews. ICES Journal of Marine Science, 6(2):305, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/305/600492.

T:1931:Re

[T.31b] H. T. Reviews. ICES Journal of Marine Science, 6(3):490–492, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/490/724591.

T:1931:Ra

[T.31c] J. B. T. Reviews. ICES Journal of Marine Science, 6(1):112–115, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/112/641608.

T:1931:Rc

[T.31d] J. J. T. Reviews. ICES Journal of Marine Science, 6(2):311–313, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/311/600511.

T:1931:Rd

[T.31e] J. J. T. Reviews. ICES Journal of Marine Science, 6(2):321–322, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/321/600620.

T:1931:Rf

[T.31f] J. J. T. Reviews. ICES Journal of Marine Science, 6(3):500–503, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/500/724609.

T:1933:Rc

[T.33a] H. T. Reviews. ICES Journal of Marine Science, 8(3):409–411, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/409/730413.

T:1933:Ra

[T.33b] J. J. T. Reviews. ICES Journal of Marine Science, 8(1):100–101, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/100/682374.

T:1933:Rb

[T.33c] J. J. T. Reviews. ICES Journal of Marine Science, 8(1):123–124, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/123/682437.

T:1933:Rd

[T.33d] J. J. T. Reviews. ICES Journal of Marine Science, 8(3):411–412, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/411/730425.

T:1934:R

[T.34] J. B. T. Reviews. ICES Journal of Marine Science, 9(3):399–400, December 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
3/399/615782.

T:1935:Ra

[T.35a] J. J. T. Reviews. ICES Journal of Marine Science, 10(2):207–208, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
207/672910.

T:1935:Rb

[T.35b] J. J. T. Reviews. ICES Journal of Marine Science, 10(2):208–209, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
208/672912.

T:1937:R

[T.37] Å. V. T. Reviews. ICES Journal of Marine Science, 12(2):220–222, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
12/2/220/753887.

T:1938:Rb

[T.38a] J. B. T. Reviews. ICES Journal of Marine Science, 13(3):384– 386, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/384/635462.

T:1938:Ra

[T.38b] J. J. T. Reviews. ICES Journal of Marine Science, 13(2):257–259, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
257/784071.

T:1939:Ra

[T.39a] J. J. T. Reviews. ICES Journal of Marine Science, 14(1):132–134, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/132/888448.

T:1939:Rb

[T.39b] J. J. T. Reviews. ICES Journal of Marine Science, 14(2):274–277, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
274/737137.

T:1939:Rc

[T.39c] J. J. T. Reviews. ICES Journal of Marine Science, 14(2):296–298, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
296/737261.

T:1939:Rd

[T.39d] J. J. T. Reviews. ICES Journal of Marine Science, 14(2):298–300, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
298/737309.

Tait:1933:R

[Tai33] J. B. Tait. Reviews. ICES Journal of Marine Science, 8(3):400–401, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/400/730345.

Tait:1936:SFH

[Tai36] John B. Tait. Salient features of the hydrography of the Northern North Sea and the Faroe–Shetland Channel in 1935. ICES Journal of Marine Science, 11(2):164–168, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/11/2/164/663189.

Taaning:1931:R

[Tån31] Å. Vedel Tåning. Reviews. ICES Journal of Marine Science, 6(3): 497–499, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/497/724605.

Taaning:1932:R

[Tån32] Å. Vedel Tåning. Reviews. ICES Journal of Marine Science, 7(2): 302–305, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/302/708331.

Taaning:1933:MMC

[Tån33] Å. Vedel Tåning. A method for marking cod: Capt. Hansen’s marking needle. ICES Journal of Marine Science, 8(3):385–387, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/385/730318.

Taaning:1937:SFM

[Tån37] Å. Vedel Tåning. Some features in the migration of cod. ICES Journal of Marine Science, 12(1):3–35, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/12/1/3/626348.

Taaning:1938:MCS

[Tån38a] Å. Vedel Tåning. A method for cutting sections of otoliths of cod and other fish. ICES Journal of Marine Science, 13(2):213–216, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
213/783955.

Taaning:1938:MSH

[Tån38b] Å. Vedel Tåning. Migrations of small halibut marked in Faroese waters. ICES Journal of Marine Science, 13(3):370–375, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
370/635443.

Taaning:1939:R

[Tån39] Å. Vedel Tåning. Reviews. ICES Journal of Marine Science, 14(2): 302–304, August 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/2/302/737347.

Thompson:1935:DIS

[TB35a] Thomas G. Thompson and Raymond W. Bremner. The determination of iron in sea water. ICES Journal of Marine Science, 10 (1):33–38, April 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/1/33/615484.

Thompson:1935:OIW

[TB35b] Thomas G. Thompson and Raymond W. Bremner. The occurrence of iron in the waters of the North-East Pacific Ocean. ICES Journal of Marine Science, 10(1):39–47, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/10/1/39/615487.

Tborade:1933:R

[Tbo33] H. Tborade. Reviews. ICES Journal of Marine Science, 8(2):263, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/263/618394.

Tully:1938:SPC

[TC38] John P. Tully and Neal M. Carter. A simplified procedure for colorimetric analyses of sea water. ICES Journal of Marine Science, 13 (1):58–66, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/58/714619.

Tester:1938:VMV

[Tes38] Albert L. Tester. Variation in the mean vertebral count of herring (Clupea pallasii) with water temperature. ICES Journal of Marine Science, 13(1):71–75, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/1/71/714633.

Thamdrup:1938:VVB

[Tha38a] H. M. Thamdrup. Der van Veen-Bodengreifer. Vergleichsversuche über die Leistungsfähigkeit des van Veen- und des Petersen-Bodengreifers. (German) [The van Veen grapple. Comparative tests on the performance of the van Veen and Petersen floor trawl]. ICES Journal of Marine Science, 13(2):206–212, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/13/2/206/783951.

Thamdrup:1938:R

[Tha38b] H. M. Thamdrup. Reviews. ICES Journal of Marine Science, 13(3): 390–392, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/390/635472.

Thompson:1930:FNO

[Tho30a] D’Arcy W. Thompson. Fridtjof Nansen. 10. October 1861–13. May 1930. ICES Journal of Marine Science, 5(2):141–148, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/141/627782.

Thompson:1930:Ra

[Tho30b] D’Arcy W. Thompson. Reviews. ICES Journal of Marine Science, 5 (2):229–230, August 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/2/229/627815.

Thompson:1930:Rb

[Tho30c] D’Arcy Wentworth Thompson. Reviews. ICES Journal of Marine Science, 5(3):419–421, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/3/419/648248.

Thorade:1930:R

[Tho30d] H. Thorade. Reviews. ICES Journal of Marine Science, 5(3):398–399, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/398/648189.

Thorade:1932:R

[Tho32] H. Thorade. Reviews. ICES Journal of Marine Science, 7(1):85–86, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/85/947038.

Thomsen:1933:CDI

[Tho33] Helge Thomsen. The circulation in the depths of the Indian Ocean. ICES Journal of Marine Science, 8(1):73–79, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/8/1/73/682461.

Thorade:1934:Ra

[Tho34a] H. Thorade. Reviews. ICES Journal of Marine Science, 9(1):108–109, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/108/944974.

Thorade:1934:Rb

[Tho34b] H. Thorade. Reviews. ICES Journal of Marine Science, 9(2):260–261, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/260/743210.

Thorade:1936:R

[Tho36] H. Thorade. Reviews. ICES Journal of Marine Science, 11(2):266, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
11/2/266/663242.

Thorade:1937:Ra

[Tho37a] H. Thorade. Reviews. ICES Journal of Marine Science, 12(3):342– 343, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/342/645368.

Thorade:1937:Rb

[Tho37b] H. Thorade. Reviews. ICES Journal of Marine Science, 12(3): 343, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/343/645373.

Thorade:1937:Rc

[Tho37c] H. Thorade. Reviews. ICES Journal of Marine Science, 12(3):343– 344, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/343/645379.

Thomsen:1938:R

[Tho38a] Helge Thomsen. Reviews. ICES Journal of Marine Science, 13(1): 107–108, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/107/714470.

Thorade:1938:Ra

[Tho38b] H. Thorade. Reviews. ICES Journal of Marine Science, 13(1): 106–107, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/106/714456.

Thorade:1938:Rb

[Tho38c] H. Thorade. Reviews. ICES Journal of Marine Science, 13(1):107, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/1/107/714459.

Thomsen:1939:NSV

[Tho39a] Helge Thomsen. Note on a secular variation of the current at the lightvessels “Horns-Rev” and “Vyl”. ICES Journal of Marine Science, 14(3):400, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/400/721643.

Thomsen:1939:R

[Tho39b] Helge Thomsen. Reviews. ICES Journal of Marine Science, 14(3): 412–414, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/412/721687.

Thorade:1939:Ra

[Tho39c] H. Thorade. Reviews. ICES Journal of Marine Science, 14(3):416– 417, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/416/721753.

Thorade:1939:Rb

[Tho39d] H. Thorade. Reviews. ICES Journal of Marine Science, 14(3): 417, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/417/721758.

Thorade:1939:Rc

[Tho39e] H. Thorade. Reviews. ICES Journal of Marine Science, 14(3): 417, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/417/721762.

Thompson:1931:SCR

[TJW31] Thomas G. Thompson, William R. Johnston, and Henry E. Wirth. The sulfate-chlorinity ratio in ocean waters. ICES Journal of Marine Science, 6(2):246–251, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/6/2/246/600439.

Thursby-Pelham:1930:CB

[TP30a] Thursby-Pelham. Current bibliography. ICES Journal of Marine Science, 5(1):127–137, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/1/127/676352.

Thursby-Pelham:1930:R

[TP30b] D. E. Thursby-Pelham. Reviews. ICES Journal of Marine Science, 5 (1):108–110, April 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/1/108/676262.

T-P:1931:R

[TP31] D. E. T.-P. Reviews. ICES Journal of Marine Science, 6(3):486–487, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/486/724553.

T-P:1932:R

[TP32a] D. E. T.-P. Reviews. ICES Journal of Marine Science, 7(1):99–100, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/99/947146.

Thubsby-Pelham:1932:CB

[TP32b] Thubsby-Pelham. Current bibliography. ICES Journal of Marine Science, 7(2):322–335, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/2/322/708410.

Thursby-Pelham:1932:CB

[TP32c] Thursby-Pelham. Current bibliography. ICES Journal of Marine Science, 7(3):431–443, December 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/3/431/694322.

T-P:1933:Ra

[TP33a] D. E. T-P. Reviews. ICES Journal of Marine Science, 8(3):406–407, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/406/730398.

T-P:1933:Rb

[TP33b] D. E. T-P. Reviews. ICES Journal of Marine Science, 8(3):407–408, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/407/730402.

Thursby-Pelham:1936:DLP

[TP36] D. E. Thursby-Pelham. The decline in the Lowestoft plaice fisheries. ICES Journal of Marine Science, 11(1):33–42, April 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/11/1/33/640280.

T-P:1937:R

[TP37] D. E. T.-P. Reviews. ICES Journal of Marine Science, 12(2): 223–224, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/223/753896.

Thursby-Pelham:1938:CBa

[TP38a] Thursby-Pelham. Current bibliography. ICES Journal of Marine Science, 13(2):274–290, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/2/274/784140.

Thursby-Pelham:1938:CBb

[TP38b] Thursby-Pelham. Current bibliography. ICES Journal of Marine Science, 13(3):412–433, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/13/3/412/635533.

Thursby-Pelham:1938:IVX

[TP38c] Thursby-Pelham. Index to volume XIII. ICES Journal of Marine Science, 13(3):434–439, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/13/3/434/635536.

Thursby-Pelham:1939:CB

[TP39] Thursby-Pelham. Current bibliography. ICES Journal of Marine Science, 14(2):316–332, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/14/2/316/737381.

Trewavas:1938:Ra

[Tre38a] E. Trewavas. Reviews. ICES Journal of Marine Science, 13(1): 124–125, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/124/714562.

Trewavas:1938:Rb

[Tre38b] E. Trewavas. Reviews. ICES Journal of Marine Science, 13(1): 125–126, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/125/714569.

Thomas:1934:ECS

[TTU34] Bertram D. Thomas, Thomas G. Thompson, and Clinton L. Utterback. The electrical conductivity of sea water. ICES Journal of Marine Science, 9(1):28–34, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/1/28/945094.

Tully:1937:GMC

[Tul37] John P. Tully. A graphical method for calculating the corrections on deep-sea reversing thermometers. ICES Journal of Marine Science, 12 (1):40–44, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/40/626359.

Tully:1938:ESW

[Tul38] John P. Tully. The expression of sea-water analyses as milligramatoms per litre. ICES Journal of Marine Science, 13(1):67–70, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
67/714626.

Tutin:1936:Ra

[Tut36a] T. G. Tutin. Reviews. ICES Journal of Marine Science, 11(3):382– 383, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/382/621062.

Tutin:1936:Rb

[Tut36b] T. G. Tutin. Reviews. ICES Journal of Marine Science, 11(3):383– 384, December 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/3/383/621068.

Thompson:1931:SGS

[TW31] Thomas G. Thompson and Henry E. Wirth. The specific gravity of sea water at zero degrees in relation to the chlorinity. ICES Journal of Marine Science, 6(2):232–240, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/6/2/232/600430.

U:1938:R

[U.38] C. L. U. Reviews. ICES Journal of Marine Science, 13(3):382– 384, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/382/635460.

U:1939:R

[U.39] C. L. U. Reviews. ICES Journal of Marine Science, 14(1):119–120, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
14/1/119/888356.

Utterback:1934:ADN

[UJ34] C. L. Utterback and Wilhelm Jorgensen. Absorption of daylight in the North Pacific Ocean. ICES Journal of Marine Science, 9(2):197–209, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/197/743142.

Utterback:1937:VCS

[UM37] C. L. Utterback and Eugene K. Miller. Variations in components of submarine daylight for 1935 and 1936. ICES Journal of Marine Science, 12(3):305–310, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/12/3/305/645334.

Utterback:1933:MSI

[Utt33] C. L. Utterback. Measurement of submarine illumination. ICES Journal of Marine Science, 8(3):388–389, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/8/3/388/730319.

Utterback:1934:RCT

[UTT34] C. L. Utterback, Thomas G. Thompson, and Bertram D. Thomas. Refractivity-chlorinity-temperature relationships of ocean waters. ICES Journal of Marine Science, 9(1):35–38, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/9/1/35/945113.

uV:1937:R

[uV37] u. V. Reviews. ICES Journal of Marine Science, 12(2):208–210, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
208/753848.

VanCleve:1937:EPN

[Van37a] Richard Van Cleve. An electrical plankton-net closing device. ICES Journal of Marine Science, 12(2):171–173, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/12/2/171/753765.

VanCleve:1937:FGW

[Van37b] Richard Van Cleve. A fifteen-gallon water-bottle. ICES Journal of Marine Science, 12(2):174–177, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/12/2/174/753771.

van:1938:R

[van38] E. van. Reviews. ICES Journal of Marine Science, 13(3):378– 379, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/378/635449.

van:1939:Ra

[van39a] E. van. Reviews. ICES Journal of Marine Science, 14(3):403– 405, December 1939. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/14/3/403/721649.

van:1939:Rb

[van39b] E. van. Reviews. ICES Journal of Marine Science, 14(3):414, December 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/3/
414/721723.

vonBrandt:1936:KOE

[vB36] Andres von Brandt. Zum Kalkgehalt der Ostsee: Ergebnisse der ozeanographischen Untersuchungen des Fischerei-Instituts der Universität Königsberg Pr. im Jahre 1935–36 in der mittleren Ostsee. I. (German) [On the lime content of the Baltic Sea: Results of the oceanographic investigations by the Fisheries Institute of the University of Koenigsberg Prussia in 1935–36 in the central Baltic Sea. I]. ICES Journal of Marine Science, 11(3):314–342, December 1936. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/11/3/314/620991.

vBuddenbrock:1937:Ra

[vB37a] v. Buddenbrock. Reviews. ICES Journal of Marine Science, 12 (2):226, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/226/753910.

vBuddenbrock:1937:Rb

[vB37b] v. Buddenbrock. Reviews. ICES Journal of Marine Science, 12(2): 226–227, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/226/753914.

vBuddenbrock:1937:Rc

[vB37c] W. v. Buddenbrock. Reviews. ICES Journal of Marine Science, 12 (2):236, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/2/236/753972.

vonBrand:1938:QDN

[vB38] Theodor von Brand. Quantitative determination of the nitrogen in the particulate matter of the sea. ICES Journal of Marine Science, 13(2): 187–196, August 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/2/187/783937.

VanOosten:1934:MMD

[VDJ34] John Van Oosten, H. J. Deason, and Frank W. Jobes. A microprojection machine designed for the study of fish scales. ICES Journal of Marine Science, 9(2):241–248, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/9/2/241/743160.

vanEverdingen:1930:SE

[vE30] D. Sc. E. van Everdingen. The Snellius Expedition. ICES Journal of Marine Science, 5(3):320–328, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/5/3/320/648148.

vanEverdingen:1936:R

[vE36] E. van Everdingen. Reviews. ICES Journal of Marine Science, 11(2): 263–265, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/263/663237.

vanEverdingen:1937:R

[vE37] E. van Everdingen. Reviews. ICES Journal of Marine Science, 12 (3):338, December 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/3/338/645349.

Verjbinskaya:1932:ONC

[Ver32] N. Verjbinskaya. Observations on the nitrite changes in the Barents Sea. ICES Journal of Marine Science, 7(1):47–52, April 1932. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/7/1/47/947000.

vanRiel:1932:SE

[vR32] P. M. van Riel. The Snellius Expedition. ICES Journal of Marine Science, 7(2):212–217, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/7/2/212/708169.

vanRiel:1935:R

[vR35] P. M. van Riel. Reviews. ICES Journal of Marine Science, 10 (2):193, August 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/2/193/672856.

vanRiel:1936:VTS

[vR36] P. M. van Riel. Variations in temperature and salinity in the Northern North Sea. ICES Journal of Marine Science, 11(2):169–182, August 1936. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/11/2/
169/663195.

VonBuch:1934:STT

[VV34] Kurt Von Buch and Charley Von Gustafsson. Der Salz- und Temperaturfehler von Thymolblau. (German) [Salt and temperature error of thymol blue]. ICES Journal of Marine Science, 9(2):173–180, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/173/743123.

vanVeen:1937:LE

[vV37] J. van Veen. Letter to the Editor. ICES Journal of Marine Science, 12(3):335–336, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/12/3/335/645343.

vanVeen:1938:WMS

[vV38] Joh. van Veen. Water movements in the Straits of Dover. ICES Journal of Marine Science, 13(1):7–36, April 1938. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/13/1/7/714629.

W:1930:Rb

[W.30a] F. S. W. Reviews. ICES Journal of Marine Science, 5(3):428–429, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/428/648265.

W:1930:Rc

[W.30b] F. S. W. Reviews. ICES Journal of Marine Science, 5(3):429–430, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/429/648269.

W:1930:Ra

[W.30c] H. M. W. Reviews. ICES Journal of Marine Science, 5(1):124–126, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
1/124/676347.

W:1931:R

[W.31] J. F. G. W. Reviews. ICES Journal of Marine Science, 6(1):115–116, March 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
1/115/641611.

W:1932:Ra

[W.32a] F. S. W. Reviews. ICES Journal of Marine Science, 7(1):97–98, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/97/947133.

W:1932:Rb

[W.32b] F. S. W. Reviews. ICES Journal of Marine Science, 7(1):98, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/7/1/
98/947136.

W:1932:Rc

[W.32c] F. S. W. Reviews. ICES Journal of Marine Science, 7(1):98, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/7/1/
98/947141.

W:1932:Rd

[W.32d] H. M. W. Reviews. ICES Journal of Marine Science, 7(1):114, April 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/7/
1/114/946963.

W:1933:Ra

[W.33a] F. S. W. Reviews. ICES Journal of Marine Science, 8(1):120–121, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
1/120/682429.

W:1933:Rb

[W.33b] H. W. Reviews. ICES Journal of Marine Science, 8(2):264, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/264/618398.

W:1933:Rc

[W.33c] H. W. Reviews. ICES Journal of Marine Science, 8(2):264–265, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/264/618400.

W:1933:Rd

[W.33d] H. W. Reviews. ICES Journal of Marine Science, 8(2):265–266, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/265/618405.

W:1933:Rf

[W.33e] H. W. Reviews. ICES Journal of Marine Science, 8(3):392, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/392/730324.

W:1933:Rg

[W.33f] H. W. Reviews. ICES Journal of Marine Science, 8(3):392–394, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/392/730326.

W:1933:Rh

[W.33g] H. W. Reviews. ICES Journal of Marine Science, 8(3):394, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/394/730328.

W:1933:Re

[W.33h] R. S. W. Reviews. ICES Journal of Marine Science, 8(2):282–283, August 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
2/282/618441.

W:1934:Ra

[W.34a] H. W. Reviews. ICES Journal of Marine Science, 9(1):100–101, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/100/944925.

W:1934:Rc

[W.34b] H. W. Reviews. ICES Journal of Marine Science, 9(2):265–266, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/265/743286.

W:1934:Rd

[W.34c] H. W. Reviews. ICES Journal of Marine Science, 9(2):266, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/266/743294.

W:1934:Re

[W.34d] R. S. W. Reviews. ICES Journal of Marine Science, 9(2):268–269, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/268/743370.

W:1934:Rf

[W.34e] R. S. W. Reviews. ICES Journal of Marine Science, 9(2):269–270, August 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
2/269/743377.

W:1934:Rb

[W.34f] S. A. W. Reviews. ICES Journal of Marine Science, 9(1):133, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/133/945067.

W:1935:Ra

[W.35a] H. W. Reviews. ICES Journal of Marine Science, 10(1):102, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
102/615377.

W:1935:Rb

[W.35b] H. W. Reviews. ICES Journal of Marine Science, 10(1):102–103, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/1/
102/615383.

W:1935:Re

[W.35c] P. B. W. Reviews. ICES Journal of Marine Science, 10(2):218, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/10/2/
218/672934.

W:1935:Rc

[W.35d] R. S. W. Reviews. ICES Journal of Marine Science, 10(1):110–111, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/110/615429.

W:1935:Rd

[W.35e] R. S. W. Reviews. ICES Journal of Marine Science, 10(1):125–126, April 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/1/125/615467.

W:1935:Rf

[W.35f] R. S. W. Reviews. ICES Journal of Marine Science, 10(3):330– 331, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/330/659260.

W:1936:R

[W.36] H. J. B. W. Reviews. ICES Journal of Marine Science, 11(2): 261–262, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/261/663235.

W:1937:Rc

[W.37a] F. S. W. Reviews. ICES Journal of Marine Science, 12(3):360, December 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/3/
360/645444.

W:1937:Ra

[W.37b] H. W. Reviews. ICES Journal of Marine Science, 12(1):87–88, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/1/
87/626423.

W:1937:Rb

[W.37c] H. W. Reviews. ICES Journal of Marine Science, 12(2):224–225, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/12/2/
224/753900.

W:1938:Ra

[W.38a] F. S. W. Reviews. ICES Journal of Marine Science, 13(1):130, April 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/1/
130/714595.

W:1938:Rb

[W.38b] H. W. Reviews. ICES Journal of Marine Science, 13(2):245–246, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/2/
245/783999.

W:1938:Rc

[W.38c] R. S. W. Reviews. ICES Journal of Marine Science, 13(2):254–255, August 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
13/2/254/784060.

W:1939:Ra

[W.39a] H. W. Reviews. ICES Journal of Marine Science, 14(1):117, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
117/887981.

W:1939:Rb

[W.39b] H. W. Reviews. ICES Journal of Marine Science, 14(1):117, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
117/887984.

W:1939:Rc

[W.39c] H. W. Reviews. ICES Journal of Marine Science, 14(1):117–118, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
117/887988.

W:1939:Rd

[W.39d] H. W. Reviews. ICES Journal of Marine Science, 14(1):118, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
118/888341.

W:1939:Re

[W.39e] H. W. Reviews. ICES Journal of Marine Science, 14(1):118, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
118/888347.

W:1939:Rf

[W.39f] H. W. Reviews. ICES Journal of Marine Science, 14(1):119, April 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/1/
119/888352.

W:1939:Rg

[W.39g] H. W. Reviews. ICES Journal of Marine Science, 14(2):301–302, August 1939. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/14/2/
301/737343.

Watkin:1930:R

[Wat30a] E. Emrys Watkin. Reviews. ICES Journal of Marine Science, 5(3): 411–412, December 1930. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/5/3/411/648224.

Wattenberg:1930:R

[Wat30b] H. Wattenberg. Reviews. ICES Journal of Marine Science, 5(3):400, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
3/400/648197.

Wattenberg:1931:R

[Wat31] H. Wattenberg. Reviews. ICES Journal of Marine Science, 6(2): 287–289, August 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/2/287/600452.

Wattenberg:1932:R

[Wat32] H. Wattenberg. Reviews. ICES Journal of Marine Science, 7(2): 286–288, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/286/708230.

Webb:1930:ARV

[Web30] H. Marshall Webb. An apparatus for recording the valve movements and the extrusion of dejecta of oysters. ICES Journal of Marine Science, 5(3):361–382, December 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/5/3/361/648155.

White:1937:R

[Whi37] P. Bruce White. Reviews. ICES Journal of Marine Science, 12(1): 100–101, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/100/626324.

White:1938:R

[Whi38] P. Bruce White. Reviews. ICES Journal of Marine Science, 13(3): 410–411, December 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/3/410/635532.

Williams:1931:R

[Wil31] E. H. Williams. Reviews. ICES Journal of Marine Science, 6(3): 477–478, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/477/724528.

Wimpenny:1937:NFH

[Wim37] R. S. Wimpenny. A new form of Hensen net bucket. ICES Journal of Marine Science, 12(2):178–181, August 1937. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/12/2/178/753798.

Wimpenny:1938:DVF

[Wim38a] R. S. Wimpenny. Diurnal variation in the feeding and breeding of zooplankton related to the numerical balance of the zoo-phytoplankton community. ICES Journal of Marine Science, 13(3):323–337, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/13/3/
323/635435.

Wimpenny:1938:RME

[Wim38b] R. S. Wimpenny. A routine method for the estimation of fat in plankton and its application to herring tissues. ICES Journal of Marine Science, 13(3):338–348, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/13/3/338/635437.

Wiseman:1938:R

[Wis38] John D. H. Wiseman. Reviews. ICES Journal of Marine Science, 13(3):386–388, December 1938. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/13/3/386/635465.

Witting:1932:CMU

[Wit32] Rolf Witting. A current-meter, its use and some results. ICES Journal of Marine Science, 7(2):218–232, August 1932. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/7/2/218/708173.

Wood:1931:R

[Woo31] Henry Wood. Reviews. ICES Journal of Marine Science, 6(3):496– 497, December 1931. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/6/3/496/724604.

Wortbington:1933:R

[Wor33] E. B. Wortbington. Reviews. ICES Journal of Marine Science, 8(2): 267–268, August 1933. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/8/2/267/618410.

Wright:1938:R

[Wri38] F. S. Wright. Reviews. ICES Journal of Marine Science, 13(1): 130–131, April 1938. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/13/1/130/714599.

Wust:1930:MST

[Wüs30] Georg Wüst. Meridionale Schichtung und Tiefenzirkulation in den Westhälften der drei Ozeane. (German) [Meridional stratification and deep circulation in the western halves of the three oceans]. ICES Journal of Marine Science, 5(1):7–21, April 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/5/1/7/676385.

Wust:1931:R

[Wüs31] Georg Wüst. Reviews. ICES Journal of Marine Science, 6(2):294–298, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
2/294/600468.

Wust:1933:R

[Wüs33] G. Wüst. Reviews. ICES Journal of Marine Science, 8(3):390–392, December 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/8/
3/390/730321.

Wust:1936:R

[Wüs36] G. Wüst. Reviews. ICES Journal of Marine Science, 11(2):262– 263, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/262/663236.

Winsor:1936:SVU

[WW36] Charles P. Winsor and Lionel A. Walford. Sampling variations in the use of plankton nets. ICES Journal of Marine Science, 11(2):

190–204, August 1936. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/11/2/190/663205.

Yonge:1930:R

[Yon30] C. M. Yonge. Reviews. ICES Journal of Marine Science, 5(2):271, August 1930. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/5/
2/271/627908.

Yonge:1931:DPM

[Yon31a] C. M. Yonge. Digestive processes in marine invertebrates and fishes. ICES Journal of Marine Science, 6(2):175–212, August 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:

//academic.oup.com/icesjms/article/6/2/175/600423.

Yonge:1931:MBL

[Yon31b] C. M. Yonge. The Marine Biological Laboratory at Low Isles, N. Queensland. ICES Journal of Marine Science, 6(3):459–462, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/459/724483.

Yonge:1932:R

[Yon32] C. M. Yonge. Reviews. ICES Journal of Marine Science, 7(2): 313–315, August 1932. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/7/2/313/708379.

Yonge:1934:R

[Yon34] C. M. Yonge. Reviews. ICES Journal of Marine Science, 9(1):134– 135, April 1934. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/9/
1/134/945071.

Yonge:1935:R

[Yon35] C. M. Yonge. Reviews. ICES Journal of Marine Science, 10(2):217, August 1935. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
10/2/217/672931.

Yonge:1937:R

[Yon37] C. M. Yonge. Reviews. ICES Journal of Marine Science, 12(1): 102–103, August 1937. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/12/1/102/626329.

Zenkewitch:1931:ABW

[Zen31] L. Zenkewitch. On the aeration of the bottom waters through vertical circulation. ICES Journal of Marine Science, 6(3):402–418, December 1931. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/6/
3/402/724452.

Zarins:1935:UZM

[ZO35] E. Zarins and J. Ozolins. Untersuchungen über die Zusammensetzung des Meerwassers im Rigaschen Meerbusen und an der lettländischen Küste des Baltischen Meeres. (German) [Investigations into the composition of sea water in the Gulf of Riga and on the Latvian coast of the Baltic Sea]. ICES Journal of Marine Science, 10(3):275– 301, December 1935. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/10/3/275/659216.